[Search for users] [Overall Top Noters] [List of all Conferences] [Download this site]

Conference asdg::cmp

Title:CMP
Notice:Welcome to the CMP notesfile
Moderator:FABSIX::B_TOWERLLE
Created:Tue Oct 31 1995
Last Modified:Fri Jun 06 1997
Last Successful Update:Fri Jun 06 1997
Number of topics:94
Total number of notes:6381

5.0. "CMP.A1 EE/PE PASSDOWNS" by SUBPAC::BJUBINVILLE (NFD Fire - Rescue) Sat Nov 11 1995 00:04

    
         This note will be used for both Equipment and Process
       Engineering for the passdown for:
    
                    CMP.A1 - Strasbaugh 6DS-SP Polisher
T.RTitleUserPersonal
Name
DateLines
5.95down forceFABSIX::B_FINNTue Nov 07 1995 19:3713
	

	Ops requested that we check down force on CMP.A1.. Found it to be way 
out due to E-P transducers being out of cal.  Calibrated transducers first then 
calibrated down force for both spindles. Both came in within two passes. Both 
are well under 1%

	System is in MQC with new pad.


				A shift


5.96High Table amps, wafers slidding outFABSIX::R_GEEWed Nov 08 1995 07:0230
	* After the system had been idle for sometime breakin wafers
	  were run. On the second set of two a wafer slid out from under
	  the left carrier. The system was EMO'ed before it crashed.

	      System status at the time:

		- 24 wafers on the pad (2 suba IV, 1 IC1000)
		- 112 wafers on the carriers. (Exposure .0065 & .0070)checked.
		- Table amps at the time spiked to 53 amps.
		- Checked slurry flow, it was good.  
		- Down force was 313 lbs
		- Slurry flow set at 417 ml/min
		- 20 sec into a 180 sec polish.
		- Back pressure 14.8
		- Believe the wafers to be oxide monitors.

	   Tried to run a second time following the break-in procedure.
		- 2 wafers at 200 down force  = spike of 20 amps
	   	- 2 wafers at 250 down force  = spike of 26 amps
	
		- 2 wafers at 313 down force  = spike of 47 amps 

	   The wafers on the right carrier slid out , the system was
	   EMO'ed before it could crash. The system is now unschedule
	   down.
	   					B shift


	   

5.97down forceSUBPAC::LANDRYWed Nov 08 1995 19:2113
    Came in this morning to the system spittin' wafers from the left
    spindle.  Tried running the system to repeat the feat, and yes it 
    did, left carrier again.  EMO'd in time to save the pad/carriers.
    
    Upon examination, found several small slices of film had come loose
    from the carrier.  Swapped the carriers and the wafer slipped out again
    on the same carrier (now on the right).  Changed the carrier but the
    problem remained.  Checked the E-P regulator pressures and they were
    off on both spindles.  Brought them in and reperformed the down force
    calibration.  Completed,  and AGAIN its under 1%. Currently changing
    the pad.
    
    				jim,doc,kyle,john_p. Brendan
5.98Table stall-outsSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 09 1995 08:228
    
    
       	After countless table stalls and many wafer slide outs,
      we up'ed the slurry flow on pump #1 in Step #2 from 417 ml/min
      to 500 ml/min.  The pad break-in and MQC's were perfect.  No
      stall outs and no wafer slide outs.  
    
                                             D shift
5.99A1 INSTALLATION AND FIT-UPSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:13220
    Monday Installation Update...........
    
    There wasn't much activity today.  I got with the contractors (KSI & 
    NEVCO) and clarified a couple of questions they had concerning plumb-
    ing (DI & GCW) and slurry supply and return.  Also, since the systems
    had moved a little due to the pedistal installation, they wanted to
    reverify with me where they could and could not run the hookups.
    
    They are in the process of running pipe under the floor at this time.  
    Not much to see! There was no other activity with this system today.
    
    					Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.2                            CMP.A1                             2 of 356
SUBPAC::LANDRY                                        8 lines   7-JUN-1994 15:53
                                -< 6/7 Update >-
--------------------------------------------------------------------------------
    	Today, KSI hooked up the two GCW drains to the system.  The also
    hooked up the GCW drain on the vacuum cabinet resevoir.  Both of these 
    tasks required substantial installation work under the floor (running
    headers and sticks).  Don't be alarmed as you look through the raised 
    floor and notice some water, they had to cut into an existing line that
    contained some water.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.3                            CMP.A1                             3 of 356
SUBPAC::LANDRY                                       18 lines   9-JUN-1994 15:40
                                -< 6/9 Update >-
--------------------------------------------------------------------------------
    Today, KSI has the main system and vacuum cabinet drain lines under 
    pressure testing.  Alls well so far!
    
    KSI had a question concerning the slurry supply and return line hookup.
    They were under the assumption that the incoming and outgoing lines
    needed to be double contained plumbing.  After confering with George
    Hynes and Tom Richards, it was determined that the lines would NOT have
    to be double containment.  This will help this portion of the install
    go alot quicker.  The double contained wall would have housed a supply
    and a return line within each of the two pump hookups.  They would have 
    had to fabricate a manifold between the slurry valves and the valve
    manifold box to separate the two prior to entering/exiting the VMB's in
    order to maintain double containment integrity.
    
    Also, the electricians are beginning to address what is needed on the
    system.  Some conduit has been run (minus the wires).
    
    						Doc 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.4                            CMP.A1                             4 of 356
SUBPAC::LANDRY                                       18 lines  16-JUN-1994 15:49
                                -< 6/16 Update >-
--------------------------------------------------------------------------------
    Installation status:
    
    The electrical hookup is complete.
    
    The GCW drains have been leak/pressure checked to satisfaction.  The
    lines were then connected up to the main system.
    
    The TCU installation has begun.  There is a rack mounted that houses
    the incoming/outgoing process cooling lines.  They still need to be
    tied into the TCU.  Also, need the hookup from the TCU to the main
    system to be performed.  The electrical hookup is complete.
    
    The vacuum cabinet has the GCW drain hooked up.  The electrical is also
    complete.  Waiting for the CDA hookup to be performed.
    
    No slurry hook ups have been performed as of yet.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.5                            CMP.A1                             5 of 356
SUBPAC::LANDRY                                       15 lines  24-JUN-1994 14:05
                         -< 6/24 Installation Update >-
--------------------------------------------------------------------------------
    KSI personnel ran and hooked up the slurry lines on the system today.
    The lines are a single PVA hook up (NO secondary containment plumbing
    is required).  The lines are supported by a unistrut rack.
    
    KSI also is making good progress with the GCW drains.  The main system
    hookup is COMPLETE!  The vacuum cabinet drains are COMPLETE as well!
    Plumbing work for the TCU is on-going.
    
    The system still needs the following hook ups:
    
    	CDA, UHP DI, Cooling water in/out, vacuum cabinet tied into the
    	main system, vacuum cabinet CDA, ALL TCU plumbing requirements.
    
    						Doc
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.6                            CMP.A1                             6 of 356
SUBPAC::TOOKER                                       37 lines   8-JUL-1994 10:59
                         -< Hook-Up Status for 7/8/94 >-
--------------------------------------------------------------------------------
                CMP.A1 Tool Hook-Up Status For 7/8/94



COMPLETED ITEMS:
----------------

    *  Bulkhead 

    *  DI Water connection except for pressure gauge.

    *  All GCW drain connections

    *  Eelectrical connections

    *  NESLAB TCU soft water, process cooling water, and drain
       connections

    *  Slurry line from valve manifold box to system


OPEN ITEMS:
-----------

    *  DI Water pressure gauge

    *  DI water return line final connection

    *  CDA final connection to main system and vacuum cabinet

    *  N2 final connection

    *  All exhaust work

    *  HEPA filter connection

    *  Slurry valve manifold box valve work

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.11                           CMP.A1                            11 of 356
ASDG::POIRIER                                        64 lines   4-AUG-1994 20:51
                            -< CMP.A1 Mech Accept >-
--------------------------------------------------------------------------------
    
    Power, vacuum, CDA and D.I. water have all been supplied to the tool.
    The D.I. water pressure is still too low. Its currently at 25 p.s.i.
    up from 15 p.s.i. yesterday but still not at the required 40 p.s.i.
    
    There is currently no slurry available. We're hoping the slurry system
    will be operational by next Tuesday for the planned 500 wafer run.
    If not, then we will perform the 500 wafer run with water only. The 
    limitation of doing this is "touching down" on the polishing pad
    momentarily instead of oscillating on the pad.
    
    
    Here is a highlight of some of the work completed thus far on CMP.A1
    
    * The table, spindles and chucks have been indicated and adjusted to the
      proper specifications.
                  
    * The shuttle has been leveled and adjusted in both the forward and
      rear positions. 
    
    * The shuttle speed has been adjusted.
    
    * The shuttle hard stop positions have been adjusted.
    
    * The load stations and cleaning stations have been set to the proper 
      positions.
    
    * Calibration values have been set for Table Min Pressure, Spindle Down
      Soft and Spindle Down Hard.
    
    
    Plans for tomorrow:
    
    * Adjust spindle down force.
    
    * Adjust conditioner down force.
    
    * Calibrate slurry PH meters.
    
    * Calibrate robot positions.
    
    * Calibrate Elevator positions.
    
    * Install elevator tubs.
    
    * Cycle machine.
    
    
    Problems:
    
    * The load cell that was shipped with the machine doesn't work. The rep
      went to Conn. tonight to get another one.
    
    * Some of the poly welds cracked during shippment.
    
    * Rust was discovered in the weldment area.
    
    * The Thompson bar way bearings that support the robot are noisy.
    
    * The lead screw bearing on SEND elevator #2 is binding.
                                            
                                          
                                                      
    * 
5.2SPINDLE AIR CYLINDER REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:1533
    
    * The left spindle air cylinder was replaced today. The reason was that 
      the spindle would not retract the last 1/4 " of its travel. The
      cylinder replacement required us to recalibrate the spindle down force
      as well.
    
    * The conditioning arm assembly was reinstalled after the replacement 
      of the mounting shaft. This shaft suffered broken a weld last
      Thursday. All adjustments and locations were set today.
    
    * We opened the valve in the slurry valve box and slurry is now flowing
      in the tool. The slurry flow rate was also calibrated today.
    
    * The send and recieve elevators were recalibrated. This was done
      because the wafers were dragging out of slots # 23, 24 & 25.
      Since we are mounting the cassettes H-BAR up, I'm not sure if were
      referencing from a surface on the cassette that is held to tight 
      tolerences. Also, since the problem seems to come and go, I'm not
      confident that the cassettes are uniform to each other. I'll take a 
      look at the Empak print when I get a chance.
    
     Plans for tomorrow:
     -------------------
     1) Change polishing pad.
    
     2) Install latest rev software.
    
     3) Start setting up CMP.A2.
    
     F.Y.I.:
     ------
     A new phone has been installed in the CMP service bay, the # is 5804.
                                                          
5.3PCW LEAK AT TABLE ROTARY UNIONSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:1822
ASDG::POIRIER                                        19 lines   6-SEP-1994 23:38
                                 -< PCW leak >-
--------------------------------------------------------------------------------
    
    * A large puddle of PCW was discovered on the floor under CMP.A1 at
      4:30 this afternoon. Tim, Brian and myself, flashlights in hand,
      found the cause to be a loose hose clamp on the hose leading to the 
      rotory union. I tightened the clamp and the leak stopped. While I was 
      under the tool, I tightened the rest of the hose clamps on hoses that
      were flowing PCW. Most of them were fairly loose. 
    
      Brian brought out some spill pillows and we cleaned-up the spill on 
      the floor. I also cleaned-up the inside of the base. Any water in
      the base will be from another source. 
    
    * I also noticed a slight PCW leak from the fitting on the bottom of the
      rotory union. I left the PCW valve off until we repair the fitting.
    
    
    * The system was left in WET IDLE mode.  
    
                                                                     J.P.
5.4CONDITIONER EXCESSIVE PLAYSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:1817
ASDG::POIRIER                                        14 lines  25-AUG-1994 23:43
                          -< Conditioner Performance >-
--------------------------------------------------------------------------------
    
    The cylinder assembly that moves the conditioning arm had quite
    a lot of play. I removed the assembly, disassembled it and tightened
    the lock nut on the end of the lead screw. It was reassembled and a
    test run was performed. The conditioning disk seems to have less
    oscillation the it did this afternoon. I'm concerned that yesterdays 
    crash may have degraded the performance of this assembly although 
    the Strasbaugh rep claims that the larger conditioning disk is 
    responsible. Its easy enough to prove, lets swap disks and see what
    happens. 
    
    Please note that the down force was reduced to 50 and the polish time 
    to 2 seconds for my test run.                       
                          
5.5CONDITIONER ADJUSTMENTSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:1932
                          -< Conditioning arm issues >-
--------------------------------------------------------------------------------
    
    * The new base stiffening bracket was tested today under process
      conditions,(20 lbs of down force). The conditioning arm osculation
      was greatly reduced and the base did not flex. Any remaining movement
      in the arm is due to the design of the conditioning arm mechanisim.
    
    
    * The strasbaugh rep also readjusted the rotating sleave in the
      conditioning arm mechanisim. This sleave did not have enough
      stock milled out of it and thus provides minimal clearance for the 
      socket head cap screws that slide inside the sleave. When the screw
      heads contact the sides of the sleave, a hesitation is induced into
      the vertical motion of the conditioning arm. Larry is working this
      issue with Strasbaugh.
    
     
    * The Strasbaugh rep drilled and tapped 10-32 holes into 2 carriers and
      installed the new check valves. He ran a test run with only the 
      center ten carrier ports open and was able to lift off the pad
      without loosing a wafer. It appears that the new check valves will
      solve the lift off problems we've been experiencing with the carrier
      ports open in the center ten configuration.
    
    * The Strasbaugh rep brought an issue to my attention tonight. One
      the carrier backing plates was damaged by what appear to be hammer
      marks. The carriers are machined to tight specifications and are
      very expensive. Please handle them with extreme care.
    
    
                                                                      John.
5.6TCU FLOW AND ADJUSTMENTSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2022
SUBPAC::LANDRY                                       19 lines   9-SEP-1994 17:14
                      -< The TCU issue has been whipped! >-
--------------------------------------------------------------------------------
    Chiller Update:
    
    *  The plumbing swap was performed yesterday.
    
    *  Glycol was procured and poured into the chiller.
    
    *  Eric Ronnebaum worked on the rotary joint and it's solenoid for
       liquid flow.  Found the connectors on the wires for the solenoid
       to be a problem.  They came from the factory crimped over the 
       wire insulation.  The valve was working intermittenly.  This is
       attributed to the  vibration in the area making/breaking contact.
    
    *  Check for leaks after the work was performed.  All looks fine!
    
    *  The chiller has been shut down for the weekend and the valves that
       pertain to this plumbing loop have been shut off (both at the system
       and at the TCU).
    
                                                       Doc
5.7CONDITIONER DOWNFORCESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2111
ASDG::POIRIER                                         8 lines  13-OCT-1994 23:37
                    -< Inconsistant conditioner down force >-
--------------------------------------------------------------------------------
    
    * Eric noticed that the conditioner downforce is not constant. He
      checked the amplifier and found noise on the line. He's going to 
      check the shielded cable for proper grounding. He is also going to 
      look at the routing of the cable for a possible source of the
      interference.
    
                                                                      John.
5.8CONDITIONER GROUNDING NOISESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2211
ASDG::POIRIER                                         8 lines  18-OCT-1994 00:15
                      -< Conditioner arm control cable. >-
--------------------------------------------------------------------------------
    
    * The Strasbaugh rep re-routed the shielded cable for the conditioner
      arm. He put a scope on the line and the noise was not present.
      It appears that the cable was routed too close to the servo motors.
    
    
                                                                       
                                                                      John. 
5.9LOW VACUUM ON CARRIERSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2312
ASDG::POIRIER                                        19 lines  26-OCT-1994 00:10
                               -< Vacuum errors >-
--------------------------------------------------------------------------------
    
    * The system went down for low vacuum on the carriers. We unloaded the
      carriers, removed the spindle chucks and measured the vacuum at the
      bottom of the spindles. The vacuum was good on both spindles. We then
      replaced the spiindle chucks and took a look at the o-rings. The
      inner most ring was compressed below the surface of the chuck. I 
      swapped the inner most ring on both spindles with the rings from A2
      and the vacuum error was corrected.
    
5.10DI WATER LINES VIBRATINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2419
ASDG::POIRIER                                        16 lines   7-NOV-1994 23:41
                            -< 2nd shift passdown >-
--------------------------------------------------------------------------------
    
    * The D.I. water valves are still chattering on both A1 & A2. Eric
      wants to check the routing of all the tubing going into and out of
      the valve manifold tomorrow. 
    
    * Eric called Strasbaugh and ordered 2 new valves that will be used
      to control D.I. water flow to the conditioner park station dishes.
      The valves will control flow to the conditioning disk rinse lines
      that Jim and I are adding to the park stations on A1 & A2..
    
    * Eric is planning to add a D.I. water valve in line with the
      conditioner park station water supply line. This should eliminate
      the siphoning problem we've been experiencing during WET IDLE mode.
    
                                                                    John.
      
5.11CONDITIONER ARM WOULDN'T GO UPSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2416
ASDG::POIRIER                                        13 lines   9-NOV-1994 23:40
                 -< conditioner assembly micro-switch problem >-
--------------------------------------------------------------------------------
    
    * The system went down for a conditioning arm problem around 7:00 p.m.
      The arm would move to the UP position and lock up during the
      initialization sequence. The problem was with the conditioner
      micro-switch. I readjusted the switch and the problem went away.
      I'm not sure why the adjustment changed, maybe the switch arm
      loosened up and moved a little bit. 
    
    * The system was wiped down and the slurry lines were flushed at the
      end of the shift. 
      
                                                                    John.
    
5.12ELEVATOR WOULDN'T MOVESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2516
ASDG::POIRIER                                        13 lines  18-NOV-1994 23:17
                             -< Elevator problems >-
--------------------------------------------------------------------------------
    
    * Brian and I removed the elevator assembly on SEND ELEVATOR #1. A
      partially disconnected electrical connector was discovered under the
      elevator servo encoder. The connector was repositioned and the
      problem disappeared. 
    
      FYI: I noticed some corrosion starting to form on the terminal block
           connections on the elevator assembly. Its probably caused from
           the nasty environment of D.I. water and slurry that was all 
           over everything. I wish Strasbaugh used connectors designed
           for this type of environment. 
    
                                                                    John.
5.13LOW SLURRY FLOWSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2625
SUBPAC::LANDRY                                       22 lines  10-DEC-1994 16:07
                -< The housing was jammed up and jelly tight! >-
--------------------------------------------------------------------------------
    
    	There was a problem today with the slurry flow.  The day started
    off with a low flow situation and progressively got worse until it
    stopped flowing altogether.  Led by Jim Tooker, the EE grouped jumped
    into action!  It was observed that the pressure gauge on the slurry
    loop was reading nil (0).  Ashland was contacted and they investigated
    on their end.  All was fine at the distribution unit.  Back to EE's
    court.  We noticed that the Speedfam flow was adequate.  Hhhmmm, must
    be an air leak or flow restriction on the system alone.  Sure enough,
    the filter was discovered to be "caked" with a jelly-ous substance, 
    which turned out to be slurry transforming into cement.  Both filters
    were removed from the system and the flow was checked.  Voila, nice
    flow.
    
    The state of the filter housings is as such:  Both are empty, due to we
    currently have no replacement for the 10 micron filter at this time.
    Therefore, the smaller micron filter would have clogged real quick so
    it had to be removed as well.  Bobby  will put the system in a state
    of warning until this situation is rectified.
    
    						EE 
    
5.14ELEVATOR PULLEYS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2715
ASDG::POIRIER                                        12 lines  13-DEC-1994 00:03
                            -< 2ND SHIFT PASSDOWN >-
--------------------------------------------------------------------------------
    
    * WE RECALIBRATED THE SLURRY FLOW RATES ON BOTH SLURRY PUMPS.
    
    * WE REMOVED BOTH (SEND) ELEVATORS ASSEMBLIES TONIGHT.THE
      PLASTIC PULLEYS WERE REPLACED WITH THE NEW ALUMIMUM STYLE
      PULLEYS THAT STRASBAUGH SENT US A FEW WEEKS AGO.
    
      THE ELEVATORS WERE DISASSEMBLED, CLEANED AND RELUBRICATED.
      THEY WERE REINSTALLED INTO THE TOOL AND RELEVELED. A TEST RUN
      WAS COMPLETED WITH NO HANDLING ERRORS. 
    
                                                               JOHN & BOB.
5.15ROBOT BINDING UPSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2713
PATE::GEE                                            10 lines  11-DEC-1994 00:31
                                   -< robot >-
--------------------------------------------------------------------------------
    * Removed robot, robot slide rails, and ball screw. Disassembled
      and cleaned the slide rail assembly. Replaced the pillow block
      bushings with the new ball bearing style. Reassembled and reinstalled
      entire assemble. Ran a test run with no handling problems.
    
      The Pillow block bushings were caked up with dryed Apieson grease.
    
    John and Bob
       
     
5.16ROBOT SENSORS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:278
ASDG::POIRIER                                         5 lines  15-DEC-1994 00:06
                            -< 2nd shift passdown >-
--------------------------------------------------------------------------------
    
    * The Strasbaugh rep replaced the robot (lower arm) home sensor.
      He also replaced the robot servo encoder connectors.
    
                                                              John & Bob.
5.17ROBOT COMMUNICATION PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2810
SUBPAC::LANDRY                                        7 lines  14-DEC-1994 21:13
                         -< The robot is coming along >-
--------------------------------------------------------------------------------
    Today, Eric Ronnebaum arrived to address the infamous robot perils.
    He had special connectors/pins shipped in to try and fix the less than
    acceptable configuration that exists.  The robot is working fine now
    and Eric is finishing up tweaking calibration adjustments on the robot
    to elevator workings.
    
    						Doc
5.18CONDITIONER ACTUATOR ALL LOOSESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:2919
ASDG::POIRIER                                        16 lines  16-DEC-1994 00:15
                            -< 2nd shift passdown. >-
--------------------------------------------------------------------------------
    
    * Disassembled conditioning arm actuator and tightened rack assembly.
      The backlash in the actuator has been increasing over the last few
      weeks. Strasbaugh has agreed to send us their new anti-backlash
      actuators for both tools.
    
    * We replaced the right spindle rotory union. It had developed a leak.
      Strasbaugh is now using an upgraded version rotory union, they are
      sending us 4 new ones next week.
    
    * We realigned both left and right spindle hubs tonight. The right hub
      was out of alignment by .010" and the left by .030". They both are
      now less than .001".
    
    
                                                            John & Bob.
5.19BRUSHES WEREN'T SPINNINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:5610
SUBPAC::BJUBINVILLE "Quit crying and play hockey!"    7 lines  20-DEC-1994 08:15
                            -< Passdown for 12/20 >-
--------------------------------------------------------------------------------
    
       The brushes weren't spinning on the wafer scrub station.
    I found the mounting piece on the brush to be mounted
    upside down.  I flipped it around and the brushes fuctioned
    properly.
    
                                    BJ
5.20ROBOT AND SHUTTLE TRAVEL PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:5611
SUBPAC::LANDRY                                       13 lines  20-DEC-1994 17:06
                    -< slimy work, but it had to be done! >-
--------------------------------------------------------------------------------
    Lots of fun here today..........
    
    *  Started off the shift with the robot not moving it's fuul sweep. 
       Cleaned and lubricated the rails.  Working fine, for now.
    
    *  A half hour later, the shuttle movement was unacceptable.  So, I
       cleaned and lubricated those rails as well.
    
5.21PARTICLESSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:5710
ASDG::POIRIER                                        12 lines  21-DEC-1994 00:12
                         -< Particle problem resolved >-
--------------------------------------------------------------------------------
    
    * Particle problem turned out to be contaminated load stations. This
      condition was brought on by yesterday's wafer breakage. The
      contamination was carried over from the pad to the load stations
      by the carriers. The load stations were cleaned by first shift and
      the MQC's were in spec. (600 particles)
    
5.22SLURRY FLOW PROBLEMSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:5717
SUBPAC::LANDRY                                       14 lines   2-JAN-1995 15:30
              -< much like my post holiday head, it was clogged! >-
--------------------------------------------------------------------------------
    
    *  Slurry flow problem:
    
    	-  Found filter to be clogged.  This was the original filter
    	   that was removed from the system prior to Christmas shut
           down.  The filter had it's ports capped off, hence retaining
           moisture in the filter core.  We attempted to re-install it
           but it proved to be an unsuccessfull endeavor.  The last in-
           house filter was installed and flow performance was verified.
    
    	-  Performed a slurry flow calibration check after the new filter
           was installed.  NO adjustments were necessary for either pump.
    
    							Doc
5.23ROBOT ARMS, POSITIONING PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:5815
SUBPAC::LANDRY                                       12 lines   5-JAN-1995 16:59
                   -< it ain't broke, but it is dislocated! >-
--------------------------------------------------------------------------------
    The robot home position went out of kilter somehow.  It happened as
    the arm was heading in to take a wafer from the load cassette.  The
    top arm was noticeably off by an good inch or two.  WE mechanically
    moved the top arm to line up over the top arm, then referenced it.  
    It is picking up and depositing down the wafer as should.
    
    We're not totally sure how the arm could have shifted as the top arm
    retaining collar appeared to be tight enough to hold it in place.  
    Also, once the arm is "hit", it loses it's torque.  We should probably
    keep an close eye on this.
    
    						1st shifters
5.24CONDITIONER ARM ACTUATOR BACKLASH SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 02:588
SUBPAC::LANDRY                                        6 lines  10-JAN-1995 17:36
--------------------------------------------------------------------------------
    The condition had some excessive backlash this morning.  Brian, Jim
    and myself removed the conditioner motor and tightened up the "spacely"
    sprocket.  Backlash has been reduce, but will be re-done for better
    performance once time is alotted.
    
    					1st shift & BJ
5.25LOAD JET SPRAY NOT COMING ONSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:0122
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             19 lines  18-JAN-1995 09:12
                        -< 3RD SHIFT PASSDOWN FOR 1/18 >-
--------------------------------------------------------------------------------
    
    
        The right load spray jet was not coming on during the
    
    polish cycle.  The problem was experienced before with the
    
    service rep and he had concluded it was just a sticky solenoid. 
    
    The problem was traced to a wire that came out of its socket 
    
    inside of the solenoid head.  It has also been noted that the
    
    solenoid is suppose to be mounted to its panel but an unseen
    
    aggressor thought smacking it would free up the valve.  It's
    
    not going anywhere but I'll look into replacing the valve tonight.
    
                                                       BJ
5.26ELEVATOR MOTOR PULLEYS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:0116
ASDG::POIRIER                                        14 lines  20-JAN-1995 23:35
            -< Weekly P.M. - Elevator Pulleys - Park Station Bowl >-
--------------------------------------------------------------------------------
    
    
    * The plastic servo motor pulleys on both Recieve Elevators were
      replaced with the new aluminum type that Strasbaugh sent us.
      This completes the change over of all eight elevators on both
      Strasbaugh polishers.
    
    * A new D.I. line was tapped into the incomming D.I. supply line to 
      provide water to the conditioner park station bowl during wet idle mode. 
    
    
                                                     John, Bob & Jim M.
                                                                       
5.27SPINDLE ALWAYS DETECTING VACUUMSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:0223
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             20 lines  25-JAN-1995 09:24
                   -< Problems with the left spindle vacuum >-
--------------------------------------------------------------------------------
    
        The left spindle was always detecting vacuum.  A few
    
    checks were done narrowing it down to the vacuum switch.
    
    After speaking with John P. on the phone, it was confirmed. 
    
    We don't stock this item (we will now) so it had to be pulled 
    
    off of CMP.A2 which is down.  The switch was installed on A1
    
    and tested out ok.  Doc is getting the ball rolling on getting
    
    a new switch in here ASAP.  A SID sheet will also be in the works.
    
    We can't get anyone of the phone from Strasbaugh right now so we'll
    
    have to wait until 10am or so.
    
                                                     CMP'er
5.28RIGHT SEND ELEVATOR NOT HOMINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:0420
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             17 lines  26-JAN-1995 09:26
                         -< vacuum and board problems >-
--------------------------------------------------------------------------------
    
    
         The right send elevator wasn't homing or going where it
    
    was suppose to go.  We traced it back to the computer, board #6.
    
    A2 was having all kinds of left spindle vacuum problems so a 
    
    decision was made to take Board #5 out of A2 (all of the elevator
    
    driver cards are the same so it didn't matter which one we swapped 
    
    out).  We also swapped back the vacuum sensor and brushes to get a
    
    full machine running.  It is now in MQC.
    
                                                        Bob, Doc and BJ
5.29BOOT-UP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:0624
SUBPAC::LANDRY                                       32 lines  26-JAN-1995 16:54
                    -< Should have spec'd a Digital PC!!! >-
--------------------------------------------------------------------------------
    Today........                                                     
    
    Bj and I had done some component swapping (driver board, brushes and a
    pressure transducer) to bring this system back up.  Both systems (a1 &
    a2) were down for multiple reasons.  There was only one set of brushes
    so they had to go on the system we brought up.  No driver boards were
    in stock, so again, the set of good driver boards were installed.  Same
    thing with the pressure transducer (status of these component orders
    can be found at the end of this note).  After configuring this system,
    we tested out and all was fine.  We powered down to button up the
    covers, then re-powered up the system and initialized.  All looked
    good, we thought.  The system started coming up with multiple erronious
    error messages.  Took the cover off the computer and rebooted.  Every-
    thing was back to normal.  ^$*$#% IBM computers!  We are strongly 
    beginning to think this beast is causing most of our recent problems.
    Trouble last week with it affecting the conditioner arm and polish
    table!  We have to keep an eye on the cover putting tension on the 
    boards/connectors.  The system ran without problem from 10:00 AM to
    the present, except for minor MQC happenings, which a change of
    carriers should fix.
    
5.30ROBOT TRAVEL VIBRATIONSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1415
SUBPAC::LANDRY                                       12 lines  31-JAN-1995 15:15
                    -< Sha, sha, sha, sha, sha, shakin'!   >-
--------------------------------------------------------------------------------
    Unacceptable robot travel......not only was it not making it's
    destination, it was shaken' profoundly as it approached the elevators.
    Upon eagle eye examination, it was discovered that the bellows that
    protects the rails from moisture, was the culprit.  A couple of bellows
    "fins" were lodged under the robot housing.  I believe it was actually
    contacting the bearings that "ride the rail".  Gently yanked the fins
    out, repositioned the bellows and tested travel.  Hhhmmmm, not to bad!
    But, a tad bit of relubrication was added and it is now "much better
    than new".  
    
    						Doc
    
5.31NO VACUUM AT SPINDLE ERRORSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1514
SUBPAC::LANDRY                                       11 lines  27-JAN-1995 17:03
                    -< A vacuum problem, how surprizing!!! >-
--------------------------------------------------------------------------------
    
    	Well, the system had a vacuum problem mid-morning.  The right
    spindle had the problem.  No, it wasn't a pressure switch this time.
    No, it wasn't o-rings.  Hmmm, let me dig into the past and use the
    ancient art of ATS (analyticle trouble shooting).  Yes, that world
    renown trouble shooting technique came in good use.  Well, without
    getting to deep, the problem turned out to be a leak in the filter
    housing.  The leak was repaired and manufacturing commenced!
    
    						Doc
     
5.32CONDITIONING ARM NOT FUNCTIONINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1629
ASDG::POIRIER                                        26 lines   2-FEB-1995 23:39
           -< Conditioning Arm not functioning/Low Spindle vacuuum >-
--------------------------------------------------------------------------------
    
    * Problem: Conditioning arm not functioning during Manual or Auto
               Polish cycles.
    
    
        Cause: The "conditioning cycles" parameter in the Pad/Condition
               menu was set to zero.
    
        
        Acton: Entered a 1 for the conditioning cycles parameter and
               ran a few pad conditioning cycles. System performed normally.
    
    ---------------------------------------------------------------------------
    
      Problem: Right Spindle vacuum too low.
    
        Cause: The vacuum tubing inside the right Carrier was disconnected
               at the check valve.
    
       Action: The Carrier was removed from the Right Spindle and
               disassembled. The vacuum tubing was reconnected and the
               Carrier was reassembled and mounted on the Right Spindle.
               The Right Spindle now maintains proper vacuum.
    
    
                                                            John & Bob.
5.33COMPUTER BUGSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1618
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             15 lines   2-FEB-1995 10:00
                             -< Passdown for 2/2 >-
--------------------------------------------------------------------------------
    
          The robot wasn't going to where it was suppose to
    when it was told to go to the load cups.  It never errored
    out nor was it binding up.  After referencing the axis',
    it started to run ok.  Later in the morning, the left spindle 
    didn't to turn when told.  The machine was once again powered 
    down.  It looks like the computer is starting to play games 
    with us.  Please keep an eye on it.
    
         We lost slurry tonight due to the Solvent Exhaust dump on 2nd.
    It shutdown the slurry pumps including the power to drive our
    warning lights.  
    
                                                        BJ
     
5.34ROBOT NOT PICKING UP WAFERSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1719
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             16 lines  13-FEB-1995 06:35
                  -< ROBOT not picking up from LOAD or SEND >-
--------------------------------------------------------------------------------
    
    
          The robot tried once to pick up the wafer from the
    left load cup and stopped.  The machine never went into
    alarm.  It just sat there,  Upon discovery, it wasn't
    coming down enough to get the wafer.  All of the robot
    positions were referenced but the problem existed.  It
    was also noted that the robot wasn't picking up wafers
    properly at the SEND station.  The Cal. variables for 
    the robot Z-height at SEND and LOAD were different
    from the latest verion of the variable listings updated
    on 2/9.  After re-aligning the pick-up, the old numbers
    needed to be changed for better pick up.  No problems
    since.
    
                                              BJ
5.35SCRAPPING NOISE COMING FROM BRIDGESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1738
SUBPAC::LANDRY                                       35 lines  10-FEB-1995 16:20
                -< Report from the Noise Reduction Task Force >-
--------------------------------------------------------------------------------
    *  The infamous noise coming from the bridge movement has been solved!
    
    	I was able to get the system this morning for my WEEKLY PM.  After
    that was accomplished, manufacturing was able to let me keep it for a
    short time to check into the movement noise.  I, like the shifts before
    me, checked out the bridge lead screw and guide rails.  All looked
    fine.  The bridge movement didn't have any slop to it, making me think
    the bearings, lead screw and rails were tight and pretty much in ok
    shape.  Manufacturing need the system back at this time, so I gave it
    back.  After all, the system was actually running fine and wafers were
    looking good.  Just had that strange noise.
    
    	In the afternoon, manufacturing was able to give me the system
    again.  So, I started checking again.  I decided it was time to check
    out the motor/pulley/belt.  Alas, wheres the keyway!!  It came loose
    from the motor shaft/pulley.  But how in the world could the bridge be
    moving with the keyway missing?  Well, seems the pulley is a little
    different than your standard pulley.  It has an insert with three
    retaining bolts.  The bolts are to assist in locking the whole pulley
    assembly in place.  So, when the keyway came loose, the pulley slipped
    but the bolts were still tight enough to wedge the pulley onto the
    shaft (a little cocked though).  The pulley slid in up against the
    bracket (starting to slice some nice grooves into it).  Thats the noise
    we were hearing.  It's right under the bearing that was beginning to
    look like the culprit, hence the confusion.  
    
    	Anyway, repositioned the pulley with the keyway inplace and bolts
    firmly doing the secondary locking.  Adjusted the belt tension. 
    Powered up the system and home all axis'.  The bridge responded as
    should.  The noise has returned to it's normal sound.  Manually ran
    the bridge back and forth to asure proper operating performance.  All 
    are happy in Mudville!
    
    							Doc
    
5.36FAIRCHILD EP REGULATORS CALIBRATEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:1914
ASDG::POIRIER                                        15 lines  16-FEB-1995 23:55
             -< Back pressure calibration/Fairchild EP valve cal >-
--------------------------------------------------------------------------------
    
    * The spindle FAIRCHILD electro-pneumatic valves were calibrated
      tonight. The were out of cal by a very small margin. We are of
      the opinion that they are not contributing to the iratic uniformity
      problems we've been experiencing.
    
      We also completed the back pressure calibration procedure. Both
      spindle pressures were within spec.
    
    
                                                        John & Bob.
5.37END EFFECTOR VACUUM PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2013
SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!"             10 lines  14-FEB-1995 02:06
                     -< END EFFECTOR AND VACUUM PROBLEMS >-
--------------------------------------------------------------------------------
    
         Between the end effector and the robot stopping and
    not erroring out is a mystery.  The vacuum sensor was
    checked on the robot and it was ok.  The air was a little
    off so I adjusted it to in spec.  The load cups sensors
    were working but needed adjustment.  After using the fomula,
    my new set point was -3 psi different from the previous setting.
    So far... we've run about 10 wafer and everything ran smooth
    without any problems.
                                               BJ
5.38DOWNFORCE READOUTS IN QUESTION, DOWNFORCE CHECKEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2122
SUBPAC::LANDRY                                       19 lines  23-FEB-1995 16:35
                      -< May the Force Cal be with you! >-
--------------------------------------------------------------------------------
    A "concern" was voiced today centering around the Force readout on the
    screen when the spindles were in the up position.  Both of the spindles
    read differently (as did the other systems two spindles, all four read
    something different).  Well, I hadn't the knowledge of why there were 
    that way and no one else was quite sure as to what the reading range
    should be.  The Manual had no information around this question.  So, 
    the spindle down force was performed to assure that the spindles were
    in fact reading right.  Other than the normal tweak here and there, all
    was fine with the spindle and it's respective load cell.  
    
    After performing the above procedure, the "idle" force read outs had
    changed but were still not in sink with each other.  The readings just
    tracked with the amount of adjustment made during calibration.  Eric B
    did all the cals with me.  He wants to run some MQC's to see what
    results the adjustments made.  Time will tell.
    
    		Doc, with early BJ assistance and late Eric assistance  
    
    
5.39LOW SLURRY PHSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2211
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"        8 lines   9-MAR-1995 06:57
                              -< LOW SLURRY2 Ph >-
--------------------------------------------------------------------------------
    
    
            "Low SLURRY2 Ph"  was coming on during the polish
    cycle.  The problem was traced to the OPTO-22 converter 
    that drives the SLURRY2 meter.  After it was reseated, 
    everything started working ok.
    
                                        BJ and Doc
5.40TABLE STALLED OUTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2220
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"       17 lines  17-MAR-1995 05:46
                    -< Table stopping during polish cycle >-
--------------------------------------------------------------------------------
    
         Entered the fab around 8pm to find A1 down with table
    problems.  During the polish cycle, the amps on the table
    draw up to 40-45amps.  The table and left spindle would come
    to a complete stop.  The servo controller was checked and
    it was ok.  First, just the right spindle was run and the table
    drew only around 10amps.  When just the left spindle was run,
    the table drew 15amps.  The carriers were swapped to verify that
    they were ok.  The slurry pumps were calibrated and found them to
    be out a small margin.  The conditioner disk was changed to try 
    and eliminate that.  After everything was done, both spindles were
    once again tried together and the table drew 30 amps at some point 
    and the table did not stop.  Jesus noticed that the results of two
    pevious wafers that the left spindle took of 3000� more than the 
    right spindle.  Until Strasbaugh can be reached, I had Gwen run 
    only the right spindle.
                                               3rd
5.41TABLE STALL OUTS, OVER CURRENTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2228
ASDG::POIRIER                                        25 lines  21-MAR-1995 20:02
          -< Table over current error / Servo amp loop gain adjusted >-
--------------------------------------------------------------------------------
    
    * PROBLEM: Still experiencing table over current errors.
    
    
    * SUSPECTED
      CAUSE :   Strasbaugh performed some calculations today and determined
                that the present settings on the Table Servo Amp are apt
                cause over current errors. They recommended that we check
                the CURRENT LIMIT pot and increase the LOOP GAIN pot from
                50% to 75%. 
    
    
     * ACTION: We increased the down force to 390 lbs in an attempt to
               cause the Table to over current. It worked, the Servo Amp
               errored every 3rd Auto Cycle on average. We checked the
               CURRENT LIMIT pot to ensure it was set to 100, it was.
               We then increased the LOOP GAIN pot from 50% to 75%.
               Strasbaugh cautioned us to not increase the loop gain
               more than 75% due to oscilation problems that could   
               cause damage to the motor and or gear box. Bobby is
               going to run the machine at the new settings and check
               for the over current problem to reoccur.
    
    
                                                          John & Bob.
5.42TABLE STALLED OUTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2319
PATE::GEE                                            16 lines  22-MAR-1995 00:03
                           -< Loop Gain Adjustment >-
--------------------------------------------------------------------------------
       * After the Loop Gain was adjusted from 50% to 75%, per the 
         Strasbaugh rep.  I began running oxide monitors at a down
         pressure of 390 for 60 seconds, (the same as earlier in the night).
         After the third run the Table amps. started to rise 5 to 10 amps
         with each run. On the seventh run the amps read over 55 and a wafer
         spit out of the left carrier, already smashed. The EMO was hit and
         the right carrier was never touched. The system was cleaned, pad
         removed and carrier disassembled.

         I think that by raising the Loop Gain we just changed the table
         amp trip point.
         

       * Also, replaced the 1/16" Robot End effector vacuum line.  

					Bob
5.43TABLE STALL OUTS, OVER CURRENTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2316
ASDG::POIRIER                                        37 lines  22-MAR-1995 19:16
          -< Table Servo Amp Adjustment results/Robot Vacuum Sensor >-
--------------------------------------------------------------------------------
    
    * Table "over current" problem:
    
      We ran over 50 wafers thru the system today with no "over current"
      errors. Since the "over current" problem was occuring intermittently,
      we cannot be absolutly sure that the present Servo Amp (loop gain) 
      setting will be satisfactory in the long term. However, compared
      to the prevous week's performance, the problem has ceased to exist.
      We are reasonably confident that this problem has been solved and
      will continue to monitor the Table Servo Amp performance regularly.
    
    
    
5.44END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2440
ASDG::POIRIER                                        37 lines  27-MAR-1995 19:52
                    -< Robot End Effector Vacuum Problems >-
--------------------------------------------------------------------------------
    
    
  1) PROBLEM: Robot End Effector Vacuum failing intermittently.
    
    
    
     POSSIBLE
        CAUSE: Found the robot end effector vacuum sensor loose. 
               It may have been rotating around inside the End
               Effector. This condition may have blocked off the 
               ports to the sensor causing the system to detect
               low vacuum and fail intermittently.. 
    
    
       ACTION: Removed and cleaned the sensor. Reinstalled it and
               bonded it in place with Loctite. Performed the "Robot
               Pneumatic Wafer Detection Vacuum Sensor Adjustment"
               procedure on page 36 in the maintenance manual. 
    
    --------------------------------------------------------------------------
    
    2) PROBLEM: No vacuum present at the Robot End Effector / Low vacuum alarms
                on both Spindles.
    
    
         CAUSE: Vacuum line open to atmoshere. Solinoid #43
                (End Effector Vacuum) hose popped off its
                barbed fitting.
    
        ACTION: Reattached vacuum hose to solinoid #43. Vacuum returned
                to its normal level of (28 in/Hg). It was at (14.5 in/Hg) 
                with the hose off. Bob is going to perform the Robot
                Vacuum Sensor Adjustment again on 2nd shift. We made
                the adjustments this afternoon before we realized the
                system vacuum was low.  
    
                                                            John & Bob.
5.45END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2514
SUBPAC::BJUBINVILLE "Brooks, Tritt and Diffie"       11 lines  30-MAR-1995 06:14
                   -< fitting stripped, sensor intermitting >-
--------------------------------------------------------------------------------
    
       End effector failed hard at the beginning of 3rd.  Went right 
    to solenoid #44 and found the white teflon elbow stripped in the
    1/4 fitting, replaced both pieces.  After removing the front cover
    of the robot, the green LED was flashing.  If there was a wafer on
    the robot, the light should stay off.  With no wafer, the light should
    stay on solid.  The machine kept stopping thinking there was a wafer
    on the end effector.  After the sensor was recalibrated, the machine
    ran great all morning.....
    
                                                     BJ
5.46ROLLERS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2627
PATE::GEE                                            14 lines  30-MAR-1995 23:24
                -< Robot End Effector air leak, Clean station >-
--------------------------------------------------------------------------------
        Replaced the clean station rollers and shoulder screws with
        a new design that arrived from Strasbaugh today. These rotate
        much better.

        I didn't get a chance to seal the air leak between the nozzle
        and the robot end effector arm tonight. Production is running
        it and it's running good.
        
        I placed a WARNING in RTC for this problem. We will need an hour
        or so for the RTV to dry plus time to recalibrate the sensor
        and verify the fix. I'll pass it on to 3rd shift.

                                                  Bob
        				
                              -< roller replaced >-
--------------------------------------------------------------------------------
    
        Replaced the new style rollers with the old style.  The
     new style caused the retaining ring to scrape up everywhere. 
     the worse ring is on public display in the chase for your veiwing
     pleasure.  John will be contacting Strasbaugh regarding this
     problem.  Longer shoulder screws would be the best solution...
    
                                                   BJ
5.47VACUUM LEAK AT END EFFECTORSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2810
SUBPAC::LANDRY                                        7 lines  31-MAR-1995 16:53
                              -< Seal that baby! >-
--------------------------------------------------------------------------------
    Currently, the system is down for a robot vacuum malfunction.  There
    is a leak on the arm that is being repaired with RTV.  Bobby has put
    the RTV on and it is currently drying.  Once dried, the vacuum sensor
    will have to be re-calibrated.  
    
    						Doc
    
5.48AIR LEAK AT END EFFECTORSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2927
PATE::GEE                                            13 lines  31-MAR-1995 23:08
                        -< Robot End Effector Air leak >-
--------------------------------------------------------------------------------
	Repaired the air leak between the nozzle and the robot end
        effector, by sealing it with RTV. Recalibrated the end effector
	wafer sensor. Ran 16 wafers with no polish, no vacuum
        problems.
 	
	The system was brought up and the warning taken off in RTC.

	It was put back down in RTC because it needs carriers and a 
        new pad. The system is powered down, the clean station brushes 
        have been remove and are submerged in water next to the sink.
                                                 
							 Bob
        				
    
SUBPAC::LANDRY                                        7 lines  31-MAR-1995 16:53
                              -< Seal that baby! >-
--------------------------------------------------------------------------------
    Currently, the system is down for a robot vacuum malfunction.  There
    is a leak on the arm that is being repaired with RTV.  Bobby has put
    the RTV on and it is currently drying.  Once dried, the vacuum sensor
    will have to be re-calibrated.  
    
    						Doc
    
5.49TABLE STALLSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:2925
SUBPAC::LANDRY                                       23 lines   5-APR-1995 17:16
--------------------------------------------------------------------------------
    Late this afternoon.......
    
    	The polish table rotation slowed down to a slow crawl.  The set
    point was 23 with an actual of 4.  The program was halted to allow the 
    carriersd to rise from the table.  The table continued at a 4 RPM
    speed.  We then "Quit Out" of the screen and the table continued to
    spin at the 4 RPM speed.  Totally ludicrous!  We had to EMO the system
    to get the table to cease.
    
    	After a reboot, they tried to run a couple of wafers to see what
    was up.  The AMPs hit about 50 (I was told) and a wafer spit out from
    under the carrier.  Aqwuick EMO saved a big silicon mess.
    
    	The computer was checked for PC board placement, pinched cables,
    loose connectors, etc.  Nothing unusual was found.  Resat the boards
    and carefull re-inserted the cover.  Rebooted the system wi8th no
    problem.  All components homed.  Ran a condition cycle to check the 
    table speed/amps.  All was normal.  
    
    	Manufacturing will run a few wafers to test the system.
    
    							Doc
    
5.50END EFFECTOR PROBLEMS - TABLE STALL-OUTSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3622
ASDG::POIRIER                                        19 lines   5-APR-1995 20:59
                 -< Robot End Effector Vacuum/Table Servo Amp >-
--------------------------------------------------------------------------------
    
    * Robot End Effector Vacuum:
    
      Two vacuum errors occured on 3rd shift last night. We cleaned the air
      nozzle on the Robot End Effector this morning and haven't had any
      problems.
    
    
    
    * Table RPM failure:
    
      The table RPM slowed down to an unacceptable level around 5:00 P.M.
      today. The tool errored for "table RPM too low". We placed a call
      to Strasbaugh and are waiting for a reply. It's likely that they will
      send a field service rep in next week to exchange the Table Servo
      Amp. The last time it failed was 1 week ago. 2nd shift has resumed
      running the tool since the problem is intermittent in nature.
    
                                                            John & Doc.
5.51TABLE STALL-OUTS, SPINDLE RPM'S TOO LOWSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3758
ASDG::POIRIER                                        12 lines  10-APR-1995 19:16
              -< Table RPM Problems/Troubleshootong Activities. >-
--------------------------------------------------------------------------------
    
    * DC Servo Module #'s 2 & 4 on the DCX Motion Control Board were
      swapped this afternoon. 2nd & 3rd shifts are going to watch for
      "RIGHT SPINDLE RPM TOO LOW" errors during test runs on both
      shifts. If the error occurs on the Right Spindle, the RPM
      problem is following the Table Servo Module (#4). Hence, bad
      Module! If the error doesn't occur on the Right Spindle,
      Strasbaugh will send in a Field Service Rep this week to
      replace the Table Servo Amp free of charge.
    
                                                              John.
                                                                   

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.173                          CMP.A1                           173 of 356
ASDG::POIRIER                                        16 lines  11-APR-1995 14:59
                       -< Table DC Servo Module Failure >-
--------------------------------------------------------------------------------
    
    * DC Servo Module (Axis #4 - Table):
    
      The "RPM TOO LOW" error has followed Module #4 from the Table
      to the Right Spindle. This verifies that the Module is failing.
      We now know that the Module is definitly part of the problem
      and hopefully all of the problem. Strasbaugh is shipping us
      a replacement Module via UPS RED. It should be on the 2:00 
      P.M. truck tomorrow. After the Module is replaced, we will
      continue to monitor the Table RPM performance to ensure the
      Table Servo Amp is operating properly. Due to this latest 
      development, Strasbaugh will not be sending a Field Service 
      Tech this week to replace the Servo Amp.
    
                                                            John.
                        

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.176                          CMP.A1                           176 of 356
ASDG::POIRIER                                        11 lines  12-APR-1995 20:07
                       -< Table Stalling Problem/Update >-
--------------------------------------------------------------------------------
    
    
    * Table Stalling Problem/Update:
    
      The replacement DC Servo Module for Axis #4 arrived this afternoon
      and was installed on the Motion Control Board. Due to the long
      periods of time between failures, it was determined that a 500 
      wafer VERIFY test should be run. This test will begin on 2nd shift
      tonight and continue thru 3rd and 1st shifts tomorrow. 
    
                                                              John.
5.52TABLE STALL OUTSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3714
ASDG::POIRIER                                        11 lines  12-APR-1995 20:07
                       -< Table Stalling Problem/Update >-
--------------------------------------------------------------------------------
    
    
    * Table Stalling Problem/Update:
    
      The replacement DC Servo Module for Axis #4 arrived this afternoon
      and was installed on the Motion Control Board. Due to the long
      periods of time between failures, it was determined that a 500 
      wafer VERIFY test should be run. This test will begin on 2nd shift
      tonight and continue thru 3rd and 1st shifts tomorrow. 
    
                                                              John.
5.53TABLE STALL-OUTS, SLURRY PUMP CALIBRATIONSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3836
ASDG::POIRIER                                        33 lines  13-APR-1995 19:30
          -< Slurry point-of-use Filters/ Table stalling prob/update >-
--------------------------------------------------------------------------------
    
    
    * Slurry Pump line:
      
      The Tygon tubing on Slurry Pump #1 blew-off this morning making
      a mess. It took a few hours to replace the tubing and clean-up 
      the slurry.
    
    
    
    * Slurry flow calibration/slurry filter problems:
    
      It is standard procedure to recalibrate the slurry flow after the 
      the tubing is changed on either slurry pump. After the tubing was
      changed, I could not calibrate the pump due to inadequate slurry
      flow to the pump. This condition was present on both Strabaugh
      polishers. The flow was restricted by the (point-of-use) filters
      that were installed in the tools last week. These filters are 
      currently under evaluation. Since there were only 2 replacements
      available, (2 per tool), CMP.A2, which is running product, recieved
      the replacement filters. CMP.A1, which is running a 500 wafer
      verification test, is operating without any point-of-use filters.
    
      NOTE: The above mentioned filters are in addition to the slurry
            filters on the bulk delivery system.
    
    
    
    
      Table Stalling Problem/update:
     
      There have been NO Table failures since the SC Servo Module was
      replaced yesterday.    
5.54HUB AND SPINDLE ALIGNMENTS DUE TO BAD UNIFORMITYSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3938
PATE::GEE                                            35 lines  18-APR-1995 23:29
                       -< Uniformity, alignment checks >-
--------------------------------------------------------------------------------
         Do to poor uniformity it was determined that we would check
        all the mechanical adjustments. We checked the table, both
	spindles and hubs, all are within spec (+/- 002).
         Replaced the pad with one that came in this afternoon and
        conditioned it. We started running wafers and ran into a 
        couple of problems.

		1. Right spindle sticking and wafers being left on the
                   table at the end of the polish cycle.

		Fix:

	           We swapped the conditioner disk and holder from
		   CMP.A2. This helped but still loosing wafers every
                   few runs on the right spindle. Disassembled the 
                   right carrier plate, cleaned and replaced the two 
                   o-rings. The problem went away.
                   
		2. Uneven back pressure between spindles. Both at Idle
                   read left 13.4 psia right 14.6 psia. 

                Fix:

                   Disassembled the left carrier plates, replaced the two
                   o-rings. Also, drained the water out of the two filters
                   on top of the spindles, they were full. Now both 
                   spindles read 14.7 psia at idle.


	    Note: 20 wafer had been run before the first mqc test. The
                  left was 14% and the right was 15% "better". We are
                  running 16 more before the next test.
 
  						2nd Shift
        				
5.55ROLLERS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:398
ASDG::POIRIER                                         5 lines  25-APR-1995 19:18
                         -< Cleaning Station Rollers >-
--------------------------------------------------------------------------------
    
    * The Cleaning Station Rollers & Shoulder Screws were replaced 
      on 1st shift today.
    
                                                                John.
5.56SHUTTLE MOVEMENT REAL SLOWSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:3912
FABSIX::R_POIRIER                                     9 lines   1-MAY-1995 23:45
                              -< Shuttle problem >-
--------------------------------------------------------------------------------
 

Shuttle moving to slow. Problem suspected to be
lack of grease on rail. John Poirier and I cleaned
and re-greased rail. Shuttle moving better but is
still slow. Unable to do any more work on the shuttle
because Operations needed A1 to run a HOT lot.

					2nd Shift
5.57MANY ROBOT END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:42240
ASDG::POIRIER                                        55 lines   3-MAY-1995 19:30
         -< Robot Problems: Homing Failure & Pneumatic Sensor Failure >-
--------------------------------------------------------------------------------
    
    
    * ROBOT PROBLEMS:
      --------------
    
       Problem:   #1) Robot "Bottom Arm" will not retain home position
                      location.
    
    
         Cause:   Moisture in the Home Position Sensor.
                            
    
        Action:   Disassembled robot and found "Bottom Arm" optical sensor
                  contaminated with moisture. We also found a small hole
                  in the Robot Bellows. The underside of the Bellows was
                  totally permiated with D.I. water. Maybe the water wicked
                  into the fabric. The Bellows appears to made from some
                  type of fabric with a rubberized coating on the outer 
                  surface. A rush order was placed with Strasbaugh for a
                  replacement Sensor and Bellows. We expect both parts to
                  arrive by tomorrow afternoon. A quantity of 2 each have
                  also been ordered and have been added to the Strasbaugh
                  BOM in the Fab 6 stockroom. 
    
    
     
      Problem:    #2) Robot End Effector failing to retrive wafers from the 
                      Load Elevators consistantly. (Vacuum does not enable
                      at the appropriate time)
    
    
        Cause:    Robot Pneumatic sensor not maintaining adjustment settings.
    
    
     
       Action:    The following items check out O.K.:
    
                  1) Vacuum leaks. (#1 cause according to Strasbaugh) 
                     a) Tubing: from Regulator to End Effector Nozzle.
                     b) No leak at Nozzle/End Effector Arm interface.
                     c) No blocked ports in Nozzle.
                     d) No cracked vacuum fittings in End Effector Arm.
    
                  2) Sol #44  (End effector vac/wafer blow off) functioning
                     properly.
    
                  3) Control system functioning properly.
    
                  
                  We have aquired the ordering information and will place a 
                  rush order tomorrow morning for a replacement sensor.   
    
    
                                                     John, Bobby & Jim T. 
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.205                          CMP.A1                           205 of 356
ASDG::POIRIER                                        41 lines   4-MAY-1995 17:53
                    -< Robot Problems - Sensors & Encoders >-
--------------------------------------------------------------------------------
    
    * Robot Problems:
      --------------
    
    1) Robot Top Arm failing Home Sequence:
    
      A new Robot Top Arm Home Position Sensor arrived and was installed.
      The Robot was reinstalled into the machine and the homing sequence
      performed. The All axis homed properly.
    
    
    2) Robot Bellows leaking:
    
      A replacement Bellows arrived today and was installed on the Robot.
    
    
    3) Robot pneumatic sensor not repeating causing vacuum pick-up errors:
    
       A new sensor was ordered and will be shipped FedEx overnight. It
       should arrive at the Fab6 dock on the 9:30 A.M. truck.
    
     
    4) After all axis sucessfully home, an error message appears on the
       CRT. (Arms not reaching zero):
    
       This is most likely a servo encoder problem with the Robot Top Arm
       servo motor encoder. We have had similar problems in the past. Roger P.
       is placing a call to Strasbaugh to confirm before he exchanges encoders
       tonight. We do stock this part. 
    
     
                                                                      John
    
    
    
    
    
     The Robot homed properly then the system errored for "Arms not reaching
      zero". 
      
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.206                          CMP.A1                           206 of 356
FABSIX::R_POIRIER                                    27 lines   4-MAY-1995 23:59
                              -< Robot problems >-
--------------------------------------------------------------------------------

Talked to Todd Forydice of Strasbaugh about problem with
Robot. Error message  "Arms not arriving at zero". He
suggested that the problem may be either a broken wire
on J3 (Robot Arm Encoder connector) or a bad Encoder. 

Removed Robot from CMP.A1 and checked for broken wires,
and shorts on connector J-3. Found a Violet wire that wasn't 
connected to a pin. Uncertain if wire was ever connected. 
Doesn't look like sheild had ever been stripped back.

According to schematic 09 SHT 14, Voilet wire should be
going to J3-5. There was no wire in J3-5, but there was a
pin in the socket. It's possible that the wire does go to
pin 5. All other un-used wires were cut short and covered
by shrink tubing. Recommend calling Strasbaugh in the morning
and finding out about the Violet wire.

Also, new Encoder from Stockroom 6 needs 2 connector pins.
And Jim Tooker does not want Robot put back in A1 because we
will be receiving the Wafer Present sensor tomorrow that 
needs to be put in.

 


					2nd Shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.207                          CMP.A1                           207 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              17 lines   5-MAY-1995 05:58
                           -< encoder wire repaired >-
--------------------------------------------------------------------------------
     
    
        Checked out schematics that Roger pointed out and it's
    safe to say that the wire that was dangling should be in the
    D-Sub connector.  Some female D-submini pins were found and 
    the wire was repaired.   The robot was then placed in the 
    machine and re-connected.  The machine booted up and homed 
    fine without any problems.  The robot was exercised in different 
    positions and it went to home everytime it was told.  No "arms 
    not at zero" errors popping up.  If however, the "arms not at
    zero" error comes back, the encoder should be the next victim.
    The robot was taken back off because Jim T. wants to replace 
    the wafer detector sensor.
                
    
                                                             BJ 
               

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.208                          CMP.A1                           208 of 356
PATE::GEE                                            13 lines   5-MAY-1995 23:36
                        -< robot , carriers , cleaning >-
--------------------------------------------------------------------------------
	The Robot end effector wafer sensor was installed and the 
	robot put back into the system today on first shift. We
	verified the repair by cycling 50 wafers using the "ZERO"
	recipe ( no touch downs onto the pad). 

	The system was then put down for new carriers, pad and a 
	cleaning (scheduled down). 4 Carriers are built and the pad
	removed.

	The system needs to be cleaned, brushs replaced and a pad
        installed for sunday night startup.

					2nd shift

              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 5.62                    CMP.A1 EE/PE PASSDOWNS                     62 of 62
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              21 lines  10-NOV-1995 03:40
--------------------------------------------------------------------------------
          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.201                          CMP.A1                           201 of 356
PATE::GEE                                            14 lines   3-MAY-1995 00:40
                           -< Robot vacuum problem >-
--------------------------------------------------------------------------------
	Problem: The robot wouldn't pick-up wafers from either send
	         station.

	Fix: Found and air leak on top of the end effector valve block.
	     Repaired the leak with RTV and cleaned the end effector.
             Still wouldn't pick-up wafers, so we performed the robot wafer
             sensor adjustment, still wouldn't pick-up wafers.

	     Checked the solenoid and all hoses for leaks, none found. 
             Reperformed the robot wafer sensor adjustment and got it to work
	     for 6 wafers, then it stopped again. Brian is working on it now.
         
	     
				2nd Shift

              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 5.63                    CMP.A1 EE/PE PASSDOWNS                     63 of 63
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              17 lines  10-NOV-1995 03:41
--------------------------------------------------------------------------------
          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.211                          CMP.A1                           211 of 356
ASDG::POIRIER                                        10 lines   8-MAY-1995 18:33
                     -< Robot Pnuematic Sensor Adjustment >-
--------------------------------------------------------------------------------
    
    * Robot pnuematic sensor adjustment:
    
      The air pressure at the End Effector Nozzle was reset to 5 psi.
      This was followed by a recalibration of the Robot Pnuematic sensor.
      A 55 wafer test run was then sucessfully completed and the system was
      returned to "UP" status in RTC.
      
    
                                                                 John.
5.58MORE END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4215
          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.212                          CMP.A1                           212 of 356
PATE::GEE                                             8 lines   9-MAY-1995 00:12
                         -< Robot End Effector Vacuum >-
--------------------------------------------------------------------------------
	System came up with an "End Effector Vacuum not found" and
	"Left spindle vacuum not found" errors tonight.

	Found a piece of teflon tape sticking out of the bottom of the 
        robot end effector valve block. Removed the tape and cleaned it up,
	All set.
 
					2nd shift
5.59CONDITIONER BACKLASH PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4351
ASDG::POIRIER                                        48 lines  11-MAY-1995 20:39
                  -< Conditioner Backlash/ RObot "Z" Height >-
--------------------------------------------------------------------------------
    
   *  Conditioner Backlash problem:
    
      The bracket connecting the Conditioner Assembly to the Actuator
      was loose. Loctite 242 was applied to the bracket screws before
      they were installed back into the bracket. The "Rod Ends" that
      mount on both ends of the Actuator are begining to show signs
      of wear. These Rod Ends are also contributing the backlash.  
      We'll contact Strasbaugh for replacements.
    
    
    
    * Robot "Z Height" problem:
    
      The Robot experienced pick-up problems at both "Send" elevators this 
      afternoon. The Robot did not complete its full travel in the "Z"
      direction. The problem was resolved by loosening the front cover on
      the Robot base. It appears that the precision slide inside the Robot
      ,that controls the linear "Z" motion, was experiencing a mechanical
      bind.
    
      The thru holes in the Robot covers are not symetrical. The opposite
      side of the cover may have been mounted facing outwards after the
      Robot was reassembled. We'll identify the covers for proper alignment
      in the future.
      
    
     
    * Air Bubbles in SLurry Line:
    
      The Tygon tubing in Slurry Pump #1 had pulled-up partially out of the
      ferrel in the compression fitting. This caused it to draw-in air. The
      tubing was replaced with a slightly longer piece and the pump flow
      was recalibrated.
    
    
    
     Machine Cleaned:
    
     The system was given a thorough cleaning this afternoon.
    
                                                                 
    
                                                                    John.
    
                                                                         
    
     
5.60MANY ROBOT END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4378
ASDG::POIRIER                                        23 lines  13-MAY-1995 12:18
                          -< Wafer Pick-up Problems >-
--------------------------------------------------------------------------------
    
    * Wafer pick-up problems:
    
      I rechecked the Robot Pnuematic Sensor adjustment. It checked-out
      O.K. The pnuematic part of the sensor appears to be changing states
      correctly because the green LED is in the correct state as the End
      Effector Air Nozzle passes over wafers. However, digital input #21,
      (End Effector Wafer Sensor), is not seeing 24 volts when the pnue-
      matic sensor is tripped. (Green LED off = tripped)
    
      I removed the Robot front cover and disassembled the connector
      containing the leads from the pnuematic sensor. I measured 24v
      going into the sensor and 0 volts comming out no matter what
      state the sensor was in. I checked the continuity in the con-
      nector and it was was O.K. As I see it, the question at this
      point is; why is there no output signal from the pnuematic
      sensor in either state? Brian will pick up the search on
      Sunday night. System will remain DOWN/Repair in RTC.
    
      Note: The system can be run in "manual load" mode if necessary.   
     
    
                                                             John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.222                          CMP.A1                           222 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               8 lines  15-MAY-1995 06:14
                           -< end effector problem >-
--------------------------------------------------------------------------------
    
          Continued working on end effector sensor.  We traced
      everything and it all comes back to the $900 sensor that
      was just replaced.  We swapped out CR18, #5 driver card 
      and OPTO-21/23 and swapped back after it was verfied as
      not being the problems.  All of the wiring was checked
      for shorts. Still no progress...
                                                  3rd shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.223                          CMP.A1                           223 of 356
ASDG::POIRIER                                        11 lines  15-MAY-1995 19:13
                        -< End Effector Sensor Problem >-
--------------------------------------------------------------------------------
    
    * Robot End Effector Sensor:
    
      The diagnosis is that the new pnuematic sensor for the End Effector
      is failing to send a signal thru it's output consistantly. We called
      the Strasbaugh Field Service Rep, explained ths situation, and he
      agreed that the sensor is most likely defective. Strasbaugh is
      shipping us a replacement via FedEx. It should arrive tomorrow
      morning on the 9:30 A.M. truck.
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.227                          CMP.A1                           227 of 356
ASDG::POIRIER                                         9 lines  16-MAY-1995 18:49
                       -< Pnuematic sensor replacement >-
--------------------------------------------------------------------------------
    
    * Robot End Effector Pnuematic Sensor Problem:
    
      The sensor was replaced and adjusted per the specs in the manual.
      A 50 wafer verify run was sucessfully completed and the system status
      was returned to UP.
    
                                                             John & Doc.
      
5.61SLURRY LOOPS CLEANEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4415
SUBPAC::LANDRY                                       12 lines  16-MAY-1995 17:47
                             -< plumbing cleaned >-
--------------------------------------------------------------------------------
    While waiting for the vacuum detection switch.........
    
    We got a chance to check out the slurry loop, 3-way valve and pH lines
    for both pumps.  They were dismantled and checked for clogs.  All lines
    were in fine shape, with no solidified slurry anywhere.  This verifies
    that the system is air tight.
    
    We cleaned and reassembled the plumbing.  Turned slurry back on and
    checked for leaks/air bubbles.  All is tight!
    
    						Doc  /  Jim
     
5.62BRUSH AIR MOTORS REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4515
ASDG::POIRIER                                        12 lines  18-MAY-1995 19:22
                           -< Air Motor replacement >-
--------------------------------------------------------------------------------
    
    
    
      * Cleaning Station Air Motor:
    
        The Left Cleaning Station Air Motor was replaced today.
        Both Stations now have new Air Motors.
    
    
     
    
                                                                    John.
5.63WAFER BREAKAGE PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4637
PATE::GEE                                            34 lines  23-MAY-1995 23:38
                         -< Wafer Breakage problems. >-
--------------------------------------------------------------------------------
	Problem: Starting a wafer breakage trend, 3 times in 3 work days.

	
	Observations:
              
               *  Ops had problems with carrier vacuum today.
                  
                  Checked the vacuum at the spindle hubs, no problems with
                  the system vacuum. Checked the check valves and they were
		  good also, could have been a bad o-ring.

	       *  The Right spindle was vibrating and the left was not.
		 
                  Checked the spindle to hub gap and found it between
                  .027 and .072 We performed the spindle to hub alignment
		  and brought it to .035 and .065 the Vibration went away.
		  
		Note: With the way the spindle and hub are made there is no
		      way to maintain a gap of .060 and also keep them aligned 
                      properly.
                      This is NOT part of the monthly P.M. and needs to be
                      added. 

	       *  Tonight we had wafers sticking to the pad very badly, this
		  was a new Speedfam pad.

		  We increased the conditioning frequency and changed the 
	          conditioning disk, nothing helped, one time we almost 
		  broke wafers. We decided to change the pad with one from
		  another lot, this seemed to stop the sticking problem.
		  The pad info is in the EST notes file.

						Second Shift
 
5.64ROBOT Z AXIS PROBLEMS, EMO DELAYSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4644
PATE::GEE                                            22 lines  24-MAY-1995 23:59
                         -< Robot top position / EMO >-
--------------------------------------------------------------------------------
	       
	      *  Broke a wafer on the pad today. 

		 Replaced pad and carriers and started running with a 
		 Strasbaugh pad. We ran into robot problems, it would
		 not move up to the top "Z" position. 
	
	      * Robot not finding top "Z" position. It will not move to top
		position.

		 We checked all the wire connections in the robot, but
		 still no go. We are in the process of removing the robot
		 from the system. Will pass on to third shift.

	      * EMO problems in front and on the top right panel.

		 Found a loose set screw on the front panel EMO button 
		 that was hitting the outside casing of the EMO. Doc
		 tightened it. Found the top right panel EMO button 
		 not traveling smoothly, so it was replaced.

						Second Shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.236                          CMP.A1                           236 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              11 lines  25-MAY-1995 07:51
                           -< Z-axis drive problem >-
--------------------------------------------------------------------------------
    
        After Roger puled the robot, it was discovered that the belt
      came off of the Z-axis and the gear that rides the shaft was all
      loosened up.  It was put back in place and re-assembled.  The
      machine homed fine.  When the robot went over to pick up the wafer,
      the Z-axis worked fine and the END EFFECT switch was changing states
      but the vacuum didn't come.  The output on the switch might have 
      blown...  again...  Doc is putting the robot back in after we checked
      the cables out again.
                                             Roger, Doc, BJ
                          
5.65ROBOT HARDWARE PROBLEMS (CORROSION, ENCODERS)SUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:47190
ASDG::POIRIER                                        17 lines  25-MAY-1995 21:53
                  -< Low Output Voltage on Pnuematic Sensor. >-
--------------------------------------------------------------------------------
    
    
    * Robot End Effector (Low Voltage Output) Problem:
    
      24v was present at the input of the sensor and only 12v
      at the output when the sensor was tripped. 12v was also
      measured at relay 18. In addition, the clipping diode
      on relay 18 was verified as GOOD.
    
      With the sensor output lead removed from the connector, the
      output voltage returned to 24v at the sensor and at relay 18.
      I suspect a short somewhere in the connector. Bob is going to
      check-out the connector tonight. 
    
    
                                                               John.
PATE::GEE                                             5 lines  26-MAY-1995 04:56
                              -< Bad connectors >-
--------------------------------------------------------------------------------
	Leon removed the robot and is in the process of rebuilding
	the shorted connectors. Found the connector for the bottom
	arm encoder had the red 5v wire broken also.

					Third Shift
          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.239                          CMP.A1                           239 of 356
SUBPAC::LANDRY                                       48 lines  26-MAY-1995 17:56
                        -< and the beat goes on...... >-
--------------------------------------------------------------------------------
    What a fiasco this has turned into.............
    
    *  Upon arriving, Leon was in the midst of finishing re-doing a couple
       of the robot connectors that had seen better days.  Upon re-instal-
       lation, it was discovered that the new problem with the robot was
       now NO power to the unit.  The connector that houses said wires was
       NOT one of the connectors re-done.  So, out the robot came again.
    
       Leon continued on with this beast while I scurried over to A2.  What
       Leon found was a broken wire within the connector.  So, he re-did
       the three wires that pertained to this problem.
    
       Again, the robot was re-installed.  Alas power was present.  We then
       tried to HOME all components.  We didn't get far.  The Z axis of the
       robot was non-functional.  Checked all associated wiring and all
       checked out fine.  You guessed it, out comes the robot yet again.
       At this time Leon departed after many hours work.
    
       Found a mechanical flaw this time.  The lead screw/bearing assembly
       was a mess.  It was wedged at the end of the lead screw.  The bear-
       ing assemblr also pulled through the bearing retaining block.  More
       relief help shows up in the person of Brian Provencher.  We dis-ass-
       embled the bearing assembly, removed some rust/corrosion build up,
       did a lube job and then re-assembled the unit.  By Brian, thanks for
       the help.  Karl Koch arrives.  He re-fortifies some connectors as I 
       re-install the bearing assembly.  The robot is re-installed and we
       try to HOME the robot.  No luck.  John Poirier now joins in.  The
       robot is removed and the assembly was re-adjusted to allow a more
       free moving actuation.  It is now Z'ing fine.
    
       Now we're back to a vacuum problem with the robot picking up wafers
       out of the cassette.  John will take over here with Debbie V and Joe
       N to trace down a voltage problem associated to the vacuum problem.
    
    *  A rep will be onsight Tuesday morning to address this issue as well
       as the A2 robot problem.  John Poirier will be contacting Strasbaugh
       with a list of items needed to repair the robots.  Among those items
       are the following:
    
    	-  New connectors (hopefully, some that are water tight)
    	-  A new lead screw for the robot (the one we have is bent)
    	-  Bearings for the lead screw (as ours are rusting)
    	_  John to request additional parts
    
    Thanks go to Brian P, Karl K, Debbie V, and Joe N for their help with
    the arduous task today.
    
    						Doc   //   Leon

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.240                          CMP.A1                           240 of 356
ASDG::POIRIER                                        21 lines  26-MAY-1995 22:25
               -< Robot Vacuum Failures / Connector Replacement >-
--------------------------------------------------------------------------------
    
    
    * Robot End Effector Vacuum Problem:
    
      Intermittent voltage readings were measured at the Robot connector
      and at relay #18. At this point, the Robot connectors are so bad
      that it's nearly impossible to cycle the tool without an error.
      The connectors have been handled so many times that their ability
      to make a connection has been severely compromised, not to mention
      the corrosion or bent leads.
    
      The current consensus is that the connectors must be replaced before 
      we can proceed any further, and that the connectors are most likely
      responsible for the End Effector vacuum failures.
    
      A Strasbaugh Field Service rep will arrive on Tuesday morning with
      upgraded (mil spec/water tight) connectors. All Robot connections
      on both CMP.A1 & CMP.A2 will be replaced. If vacuum failures persist,
      he will assist us until this issue is resolved. 
                                             
                                                                  John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.241                          CMP.A1                           241 of 356
ASDG::POIRIER                                        19 lines  30-MAY-1995 20:39
                 -< Robot RS-232 Connectors/Robot Lead Screw >-
--------------------------------------------------------------------------------
    
    Robot RS-232 Connector Problems:
    
    The local Strasbaugh Field Service Rep arrived this afternoon. We had
    hoped that he would have a replacement Robot in his possesion since
    this tool has a bent Robot lead screw in addition to the connector
    problem. Unfortunatly, Strasbaugh experiencd an internal communication
    problem this morning the rep never recieved instructions to bring a
    replacement with him. The Rep has returned to the Conn Field Service
    Office tonight to pick-up the replacement Robot. He will return
    tomorrow morning, robot in hand, to install the new upgraded conn-
    ectors and swap lead screws, bearings and the ball nut between Robots.
    
    Replacement connectors, insersion, extraction, alignment and crimping
    tools will arrive FedEx tomorrow morning at 9:30. (Shipment confirmed at
    4:00 this afternoon)
      
    
                                                                   John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.242                          CMP.A1                           242 of 356
ASDG::POIRIER                                        11 lines  31-MAY-1995 21:28
                         -< Connector Problem Update >-
--------------------------------------------------------------------------------
    
    
    Connector Problem Status:
    
    The Strasbaugh Rep arrived this morning with a new Robot. The new style
    pins and connectors were installed on the Robot on 1st shift today. The rep
    will finish the connector replacement on the machine harness tomorrow 
    morning. In the mean time, all effort from 2nd & 3rd shift will be
    focused on bringing up CMP.A2.
    
                                                                     John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.243                          CMP.A1                           243 of 356
ASDG::POIRIER                                        25 lines   1-JUN-1995 20:19
                    -< Robot Replacement / New Connectors >-
--------------------------------------------------------------------------------
    
    
    * New Robot/Connectors:
    
      The replacement Robot was installed today along with new connectors 
      on both the Machine cables and Robot wires. The system is now running
      verification test monitors and is in DOWN/VERIFY status in RTC.
      
    
    
      F.Y.I.: The rep discovered water inside 2 of the cables during 
              disassembly. The outer insulation was worn thru due
              to chaffing on both cables. The bad areas of cable were
              cut-off and undamaged cable was pulled-up thru the cable
              track before the connectors were crimped-on. There may
              still be moisture futher down inside the cables.We may want
              to consider replacing the cables during the Robot upgrade
              activity in the near future.
    
            
    
                                                                  
                                                                John.
    
                                                                  John.
5.66ROBOT END EFFECTOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:4828
SUBPAC::LANDRY                                       25 lines   2-JUN-1995 17:40
                     -< End effector vacuum switch again >-
--------------------------------------------------------------------------------
    Just when you thought the robot problems were gone..........
    
    Right about 11:45 this morning, the robot end effector came up with an
    error (wafer detected on the end effector) when in all actuallity, no
    wafer existed.  So, I took off the robot end plate, and sure enough, no
    green LED was on.
    
    Not sure what actually caused this fault.  Thw cables/connectors were
    not stressed.  The tygon tubing going to the switch was a bit
    restricted, but even when the restriction was removed, it still
    detected a wafer.  So, I performed the wafer detection calibration once
    again.  Found the pressure to be at 7 PSI, adjusted it to it's proper
    setting of 5 PSI.  Performed the normal adjustments and all was well.
    
    About the 7 PSI I found........I'm not sure if it's related or not at
    this time, but this error coincided with the CMP room incoming N2 
    pressure boost.  It's possible the switch could have been set up when
    the room N2 pressure was on the low side (@ 70 PSI this morning).  The
    pressure after the boost was @ 100 PSI.  Quite a substantial change.
    The only way to verify it would be to regulate the pressure back down
    to 70 PSI and see what affect it has on the wafer detection pressure at
    the robot.  Food for thought!!!!
    
    						Doc
    
5.67MORE CORROSION PROBLEMS WITH THE ROBOTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:53263
ASDG::POIRIER                                        56 lines   5-JUN-1995 22:05
           -< Robot Cables, Connectors & Termiator Boards Corroded >-
--------------------------------------------------------------------------------
    
    * Cracked tank on Left Send Elevator:
    
      Strasbaugh was called and all ordering information was forwarded
      to our Materials dept for processing. The Tank will be added the
      the Strasbaugh Fab6 Stockroom BOM. A min stocking level of 2 was
      requested.
    
    
    
    * Robot Z axis malfunction:
    
      The Robot was reinstalled and tested for Z-axis functionality.
      The z-axis movement was normal thru the full stroke and the home
      switch functioned properly. 
     
    
    
    * (NEW CONDITION) Robot Top Arm not homing:
    
      The Robot Top Arm would not home apon placing the Robot back into
      the system. The diagnostics screen was pulled-up and input #41 was 
      observed. No input was present as the flag passed thru the switch.
      A 24v input was measured at the switch. Next, the switch was replaced
      with a KNOWN GOOD switch, still no input #41. We then took a look at
      CR #23, the contacts did not actuate as the flag passed thru the Top
      Arm switch. We then followed the cables down to the terminator
      boards at the bottom of the electrical cabinet. (13R1 & 14R1)
      There we found corrosion, water & slurry inside the cables,
      connectors & terminator boards.
    
      Jim Tooker was called at home and made the decision to replace all
      affected cabling, connectors and boards. Strasbaugh was then called
      and agreed to a warranty replacement of all the above mentioned
      parts. They shipped the parts FEDEX Priority #1 at 4:00 pst. The
      items will arrive tomorrow either at 9:00 a.m or by 2:00 P.M. the 
      latest.
    
      Power to the system has been Locked Out and the disassembling of the 
      cables connectors and boards has been started by Bob and myself. 
      The corrosion in the all these parts is likely to blame for the
      intermittent failing of various Robot functions over the last few
      weeks.
      
    
    
     * Conditioning Arm Amplifier Cable:
    
       While skwerming around inside the base of the machine, I noticed
       that the cable for the Conditioning Arm (Down Force) Amplifier
       was damaged. It appears that the PC mounting rails have scrapped
       the insulation off the cable exposing the shielding. We'll call
       Strasbaugh in the morning for a replacement.
    
    
                                                           John & Bob.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.249                          CMP.A1                           249 of 356
ASDG::POIRIER                                        25 lines   6-JUN-1995 20:08
            -< Cable & Board Replacement / Robot Z-Axis not moving >-
--------------------------------------------------------------------------------
    
    * Robot Cable & Terminator Board replacemnt:
    
    
      The parts arrived this morning on the 9:00 FEDEX truck. The cables
      and terminator boards were replaced. The cables came with RS-232
      connectors on both ends. We left the connectors on the ends of the 
      cables that connect to the boards. The connectors were removed from
      the ends that connect to the Robot and replaced with the Water Tight
      variety. 
    
      The Robot was reinstalled and tested. One problem was encountered
      during testing. The Robot Z-axis failed to move. We tested the 
      Robot Z-Axis Home Switch while the system was powered-up. With the
      diagnostics screen pulled-up, no signal was present at input #43  
      as a flag was passed thru the switch.
    
      Bob is taking over on 2nd shift and is working with the Strasbaugh
      via the phone to continue the troubleshooting activities. If the Z-Axis
      movement issue is not resolved by the end of 2nd shift, Strasbaugh will
      send a rep here in the morning.
     
                                       
                                                    John, Tim & Bob.
                                     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.250                          CMP.A1                           250 of 356
PATE::GEE                                            34 lines   6-JUN-1995 23:57
                                   -< Robot >-
--------------------------------------------------------------------------------
	Problem: After the connectors and cables were replaced the robot
		 wouldn't move up to home. It would jam in the down position.
		 We checked the diagnostic inputs and found that neither
		 the up #43 or down #44 switches activated when tripped.

	Action: Called field service (Mike Nochols).

		* He had us check the DCX user to see if board #2 Counter #8 
                  changed when the up down motor was moved, it did, telling 
                  us that the problem was not the DCX Module for that motor.

		* Checked inputs #43 and #44 in the power distribution 
		  cabinet while toggling the switches and nothing happened.
		  #43 stayed on and #44 stayed off, this is an indication
		  that there is a wiring problem.

		* We disassembled the 9 pin connector in the power distribution
		  cabinet and found that the wires didn't match the old 
		  connector. We rewired the cable at the military connector
		  and robot did find up/down home.

		  The Robot mechanical is working fine.

	 New Problem : The Robot won't pickup wafers everytime. The Vacuum 
		       Switch will need to be looked at and also the 
		       elevator heights.

	*** As for now the rep. is planning on being here around 10:30am 
	    tomorrow. He will call before he leaves for here in the morning.
	    He also said that he is available tonight by telephone.
	    Beeper 1-800-514-1603
	

					Second Shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.251                          CMP.A1                           251 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              13 lines   7-JUN-1995 03:54
                        -< Effector dead, PM completed >-
--------------------------------------------------------------------------------
    
          Completed weekly PM while it was down....
    
    
          The problem with the Z-axis was resolved by the end of 
       the 2nd shift.  But...  the robot wasn't picking up the wafers.
       The end effector switch was way out of adjustment.  After I brought
       it back in, the robot picked up a couple of wafers but died again
       just after midnight.  We tried ajusting the switch again but it 
       didn't want to re-calibtrate.  There are no more switches in stock.
       A shortage requsest was sent in.
    
                                              BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.252                          CMP.A1                           252 of 356
ASDG::POIRIER                                        19 lines   7-JUN-1995 19:56
                    -< Robot Pnuematic Sensor Replacement >-
--------------------------------------------------------------------------------
    
    * Robot Pneumatic Sensor Problem:
    
      
      The Strasbaugh rep has been working on this problem with us since
      2:00 p.m. We verified that the wiring in the 25 conductor Robot 
      cable is correct. The problem appears to be a malfunctioning
      Pnuematic Robot Sensor again. The rep thinks that the corroded
      connectors may have allowed current to leak across the pins
      and blow the Output Transistor in the sensor. The corroded
      connectors may have been responsible for the low voltage
      readings as well. 
    
      A replacement sensor was ordered from Strasbaugh and is expected to
      arrive tomorrow morning at 9:00 A.M.  The rep will be back at 9:00 A.M.
      to install the sensor and evaluate its performance during the verifi-
      cation run.
      
                                                       John, Bob & Roger.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.253                          CMP.A1                           253 of 356
ASDG::POIRIER                                        19 lines   8-JUN-1995 18:46
                      -< Robot Pnuematic Sensor Replaced >-
--------------------------------------------------------------------------------
    
    * Robot Pnuematic Sensor replacement:
    
     The Strasbaugh rep arrived this morning and the sensor was replaced.
     A verification run was started but the sensor failed after 30 wafers.
     By "failed" I mean the sensor stopped changing states and thus didn't 
     send an output signal. A slight tweeking of the sensor pot returned
     the sensor to full functionality. 
    
     The problem we now face is SENSOR DRIFT. Noone, including the
     Strasbaugh rep understands why the sensor can't maintain a setting.
     The pnuematic hoses in the effected circuit were checked for leaks
     and none were found. The Strasbaugh rep, and possibly an assistant
     will return tomorrow to continue with the troubleshooting. 2nd
     & 3rd shift will continue to run the system and record the results.
    
    
                                                             John. 
                                                                  

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.254                          CMP.A1                           254 of 356
PATE::GEE                                            23 lines   8-JUN-1995 23:42
                            -< Robot Verification >-
--------------------------------------------------------------------------------
	Problem: Need to verify the robot pneumatic vacuum switch repair.

 	Action: 

		* Started running wafers, on the 22 wafer the robot saw
		  the wafer but didn't turn vacuum on to pick it up.
		  We turned the "SET " pod 1/2 turn clockwise at the
		  request of D.Sullivan The procedure used to set the switch,
		  (Strasbaugh Manual) and the procedure sent with the switch,
		  (Nippon Pneumatics) are not the same.

		* We are now up to 168 wafers with no problems. Third
		  shift will take over.

	FYI:
		* The rep. has indicated that we should be very careful
		  tightening the nut that holds the "P ADJ" set screw
		  over tightening may damage the switch, finger tight
		  is good.
		  
			    	

					Second Shift

              <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
                                    -< CMP >-
================================================================================
Note 5.72                    CMP.A1 EE/PE PASSDOWNS                     72 of 72
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              16 lines  10-NOV-1995 03:49
--------------------------------------------------------------------------------
          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.247                          CMP.A1                           247 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               9 lines   5-JUN-1995 06:14
                           -< Z-axis on robot lost >-
--------------------------------------------------------------------------------
    
        
           The robot started crashing into the elevator when it went
        to pick up a wafer.  The machine kept erroring with ROBOT DID
        NOT GO TO LOWER LIMIT, UPPER ect...  Homing it didn't work.
        Pulled the robot to see if there was any obstructions but none
        were found.  We are continuing to look into sensor and binding
        possibilities.
                                            3rd
5.68CONDITIONER ARM FROZESUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:5490
SUBPAC::LANDRY                                       17 lines  10-JUN-1995 11:43
                                 -< arm froze >-
--------------------------------------------------------------------------------
    Late this morning as Jim was running touch-up wafers......
    
    	The conditioner arm froze up over the middle of the table (was
    still in the UP position.  Tried re-homing but wouldn't move.  Through
    the Diagnostics menu, we could get the arm to lower and raise, but it
    would not got back or forth.
    
    	We powered down then rebooted the system.  The arm was down on the
    table once power was shut off.  Upon the HOMING sequence, it said that
    the "arm would not move up or down".  It required an answer of yes or
    no.  Of course answered no, raised the arm up using D32, then homed
    from there.  All went well.  No problems occured for the balance of his
    running time.
    
    						Doc
    
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.260                          CMP.A1                           260 of 356
PATE::GEE                                            11 lines  12-JUN-1995 08:10
                              -< Conditioner Arm >-
--------------------------------------------------------------------------------
	Problem: Conditioner arm would not move to center position,
		 but would move up and down.

 	Action:  Checked all wire connections and mechanical hardware,
		 everything looks good. Re-booted the system and it
		 is working. Made a slight adjustment on the home sensor,
		 the arm was very close to the right panel when it homed.

				  
								    	


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.261                          CMP.A1                           261 of 356
ASDG::POIRIER                                        43 lines  12-JUN-1995 19:28
       -< Conitioning Arm Travel Problem / Robot X-Axis Travel Problem >-
--------------------------------------------------------------------------------
    
    * Conditioning Arm Travel Problem:
    
      The bracket that is located between the Conditioning Arm Assembly and
      the Actuator had moved slightly out of position. This caused the head
      of the Bracket Mounting Bolt to interfere with the Jam Nut on the
      Actuator Rod End.
    
      The Bracket mounting screws were cleaned and the threads in the mounting
      holes were chased with a 1/4 - 20 tap to remove the old loctite.  
      The Bracket was reinstalled using fresh loctite.
    
      Also, the Rod Ends that are connected to both ends of the Actuator
      were both rotated 2 turns . This provided an extra 1/8" of clearance
      between the head of the Bracket Mounting Bolt and the Jam Nut on the
      Actuator. This will prevent the Assembly from binding-up if the 
      Bracket lossens-up in the future. 
      
    
    
    
      Robot X-Axis Movement Problem:
    
      This problem is related to the (Rail/Lead Screw) assembly that the
      Robot is mounted on top of and NOT the Robot itself. We called
      Strasbaugh and the recommendation was a thorough cleaning and 
      relubrication of both the Lead Screw and both Rails. Roger has
      begun this activity and expects to be finished in a few hours. 
      If a significant improvement is observed, a Pad & Carriers will
      be installed on the tool and MQC's will begin. If not, the entire
      assembly will have to be removed (at some future date), stripped 
      on the bench and thoroughly cleaned/lubricated to attain smooth
      motion. 
     
      Strasbaugh has recommended that we change lubricants on all Lead
      Screws and guide rails. They want us to stop using NSK #2 light 
      duty grease and start using Mobil Velocite #6 light oil.
      
    
                                                            John & Roger.
    
    
    
5.69CONDITIONER PROBLEMS, ROBOT X-AXIS PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:5567
ASDG::POIRIER                                        43 lines  12-JUN-1995 19:28
       -< Conitioning Arm Travel Problem / Robot X-Axis Travel Problem >-
--------------------------------------------------------------------------------
    
    * Conditioning Arm Travel Problem:
    
      The bracket that is located between the Conditioning Arm Assembly and
      the Actuator had moved slightly out of position. This caused the head
      of the Bracket Mounting Bolt to interfere with the Jam Nut on the
      Actuator Rod End.
    
      The Bracket mounting screws were cleaned and the threads in the mounting
      holes were chased with a 1/4 - 20 tap to remove the old loctite.  
      The Bracket was reinstalled using fresh loctite.
    
      Also, the Rod Ends that are connected to both ends of the Actuator
      were both rotated 2 turns . This provided an extra 1/8" of clearance
      between the head of the Bracket Mounting Bolt and the Jam Nut on the
      Actuator. This will prevent the Assembly from binding-up if the 
      Bracket lossens-up in the future. 
      
    
    
    
      Robot X-Axis Movement Problem:
    
      This problem is related to the (Rail/Lead Screw) assembly that the
      Robot is mounted on top of and NOT the Robot itself. We called
      Strasbaugh and the recommendation was a thorough cleaning and 
      relubrication of both the Lead Screw and both Rails. Roger has
      begun this activity and expects to be finished in a few hours. 
      If a significant improvement is observed, a Pad & Carriers will
      be installed on the tool and MQC's will begin. If not, the entire
      assembly will have to be removed (at some future date), stripped 
      on the bench and thoroughly cleaned/lubricated to attain smooth
      motion. 
     
      Strasbaugh has recommended that we change lubricants on all Lead
      Screws and guide rails. They want us to stop using NSK #2 light 
      duty grease and start using Mobil Velocite #6 light oil.
      
    
                                                            John & Roger.
    
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.272                          CMP.A1                           272 of 356
ASDG::POIRIER                                        13 lines  27-JUN-1995 19:03
          -< Conditioning Arm Sweep Problem - Servo Module Replaced >-
--------------------------------------------------------------------------------
    
    * Conditioner Arm Sweep Problem:
    
      The DC Servo Control Module for Axis #15 was replaced this afternoon. 
      We ran a few conditioning cycles as a test with good results.
      However, this problem has a history of being intermittent in nature
      so we'll need a few shifts of operation before can have any level
      of confidence with the fix. 
    
      The old module is being returned to Strasbaugh for a $630.00 warranty
      credit.
    
                                                             John & Roger.
5.70CONDITIONER PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 03:5592
ASDG::POIRIER                                        17 lines  13-JUN-1995 20:03
                       -< Conditioning Arm Adjustments >-
--------------------------------------------------------------------------------
    
    
    * Conditioner Arm Home Switch & Mechanical Travel Adjustments:
    
      The Actuator Rod Ends were returned to their origional positions
      due to interference with the Conditioner Assembly Sleeve. This 
      contact with the side of the sleeve caused the Conditioning Arm
      Home Switch to loose its adjustment intermittently. The Rod End
      adjustment was changed as a preventitive measure to ensure that 
      the linkage wouldn't bind-up if the Mounting Bracket lossened up
      again.
    
      It turns out that were better-off leaving the Actuator/Rod End 
      adjustment alone. We'll have to add an inspection of the Assembly
      Bracket to the Monthly P.M. based on its inability to stay tight. 
    
                                                                  John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.265                          CMP.A1                           265 of 356
ASDG::POIRIER                                        14 lines  14-JUN-1995 19:12
        -< Conditioner Arm Motion Problem/ Control Module Replacement >-
--------------------------------------------------------------------------------
    
    
    * Conitioner Arm Motion Problem:
    
      The Conditioner Arm Motion is continuing to malfunction at the
      rate of once per shift. When this occurs, a system power down
      and reboot is necessary. Strabaugh was called and recommended
      replacing the DC Servo Control Module for the Conditioner axis.
      They have agreed to send us a warranty replacement module and
      will send it FEDEX Priority #1. We expect it to arrive tomorrow
      morning at 9:00. We'll swap out the boards at an appropriate
      time tomorrow.
    
                                                               John. 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.270                          CMP.A1                           270 of 356
ASDG::POIRIER                                        21 lines  26-JUN-1995 19:41
               -< Conditioner Arm Sweep/Table Stalling Problems >-
--------------------------------------------------------------------------------
    
    * Conditioner Arm sweep problem:
    
      A replacement DC Servo Control Module for Axis #15 is in the Grey
      cabinet. Please install this module on the Motion Control 
      Board in the PC as soon as the tool becomes available.
    
     
    
    * Table Stalling problem:
    
      This problem came back today after disappearing 2 months ago. We 
      thought this issue was dead after replacement of the axis #4 
      control module. Strasbaugh was called and wants us to observe
      one more failure before sending a Field Service tech. The next 
      logical step is to replace the axis #4 Motor Controller. Strasbaugh
      has agreed to replace the Motor Controller if one more failure
      occurs, even if it happens after the August 26 expiration of our
      warranty. This is reassuring since this controller costs over 5K.
    
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.272                          CMP.A1                           272 of 356
ASDG::POIRIER                                        13 lines  27-JUN-1995 19:03
          -< Conditioning Arm Sweep Problem - Servo Module Replaced >-
--------------------------------------------------------------------------------
    
    * Conditioner Arm Sweep Problem:
    
      The DC Servo Control Module for Axis #15 was replaced this afternoon. 
      We ran a few conditioning cycles as a test with good results.
      However, this problem has a history of being intermittent in nature
      so we'll need a few shifts of operation before can have any level
      of confidence with the fix. 
    
      The old module is being returned to Strasbaugh for a $630.00 warranty
      credit.
    
                                                             John & Roger.
5.71WAFERS BROKE IN SEND CASSETTESSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:0334
PATE::GEE                                            10 lines  27-JUN-1995 07:27
                            -< Right Send Station >-
--------------------------------------------------------------------------------
	Problem: Broke a wafer tonight, two wafers came out of the right
		 send cassette, #24 and #25 slots. #24 broke on the robot 
		 body.
		 

 	Action: Releveled the elevator so it has a very slight pick towards
                the back. Adjusted Calabration variable #81 to 2.97 from
		3.04 				  
								    	


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.272                          CMP.A1                           272 of 356
ASDG::POIRIER                                        13 lines  27-JUN-1995 19:03
          -< Conditioning Arm Sweep Problem - Servo Module Replaced >-
--------------------------------------------------------------------------------
    
    * Conditioner Arm Sweep Problem:
    
      The DC Servo Control Module for Axis #15 was replaced this afternoon. 
      We ran a few conditioning cycles as a test with good results.
      However, this problem has a history of being intermittent in nature
      so we'll need a few shifts of operation before can have any level
      of confidence with the fix. 
    
      The old module is being returned to Strasbaugh for a $630.00 warranty
      credit.
    
                                                             John & Roger.
5.72SEND ELEVATOR PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:059
PATE::GEE                                             6 lines  28-JUN-1995 04:34
                    -< Right Send station/Cal.Variable #81 >-
--------------------------------------------------------------------------------
    	Had to adjust calibration variable #81 from 2.97 to 2.99 
        (axis 6 offset to slot 0). The robot was having trouble 
    	picking wafers from the right send station. Ran 24 wafers
    	with no polish and another 36 breaking in the pad and 
    	doing MQC's, no problems. 
    	
5.73EMO DELAYSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:0568
FABSIX::R_POIRIER                                    21 lines  29-JUN-1995 23:40
                            -< Problem with EMO's >-
--------------------------------------------------------------------------------

There is a problem with the EMO'S on A1. Sometimes they
take as much as 5-10 seconds to completely shut down the 
tool. This caused a wafer breakage today when we lost
Slurry for a little while. Jim Sadin tried to EMO the
tool but was unsucessful. He hit the EMO in time but
it failed to shut off the system. 

John call Strasbaugh and they recommended we replace the 
24VAC relay citing that this could be the cause of the problem.
I've tried searching thru the stockrooms and thru Maxcim
and I have been unable to find the correct relay. It appears
we will need to order a relay from Strasbaugh.

All the relays in the other bay of A1 are 24VDC relays. They
won't work for our need. The part we need is a 24VAC 10A relay

I put the old relay back in the tool. The tool is still down.
There is still a safety hazard because we don't know if the 
EMO switches will work when pushed.


          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.279                          CMP.A1                           279 of 356
ASDG::POIRIER                                         7 lines   6-JUL-1995 17:43
                        -< EMO circuit relay replaced >-
--------------------------------------------------------------------------------
    
    * The 24v AC relay for the EMO circuit was replaced this afternoon.
      No delay was observed after several EMO power-off cycles. 
    
    
    
      

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.280                          CMP.A1                           280 of 356
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              21 lines   6-JUL-1995 23:49
                -< table stall, conditioner failed, EMO delay >-
--------------------------------------------------------------------------------
    
    
           The table stalled right in the middle of polishing tonight.
       The error "TABLE SPEED TOO LOW" came up and put the machine into
       finish cycle after 20 seconds into the polish cycle.  After the
       machine was still idle, the table started turning by itself at
       3 RPM's.  The call was to log the system to EXTERNAL and wait for
       the Service Rep. to come in.  A call was placed to Mike Nichols
       but there was no call back.  Strasbaugh will have to be called
       in the morning.  
    
           The conditioner arm failed to sweep and do its conditioning
       again.  The board has already been swapped out per Strasbuaugh's
       request.  The Rep can be filled in as to that problem as well.
    
           The EMO time delay still exists.  I was pushing the front EMO
       button for 4 seconds before the system shut down.
    
           The machine is down until further notice....
    
                                                      BJ
5.74TABLE STALLS, EMO DELAYSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:07125
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              21 lines   6-JUL-1995 23:49
                -< table stall, conditioner failed, EMO delay >-
--------------------------------------------------------------------------------
    
    
           The table stalled right in the middle of polishing tonight.
       The error "TABLE SPEED TOO LOW" came up and put the machine into
       finish cycle after 20 seconds into the polish cycle.  After the
       machine was still idle, the table started turning by itself at
       3 RPM's.  The call was to log the system to EXTERNAL and wait for
       the Service Rep. to come in.  A call was placed to Mike Nichols
       but there was no call back.  Strasbaugh will have to be called
       in the morning.  
    
           The conditioner arm failed to sweep and do its conditioning
       again.  The board has already been swapped out per Strasbuaugh's
       request.  The Rep can be filled in as to that problem as well.
    
           The EMO time delay still exists.  I was pushing the front EMO
       button for 4 seconds before the system shut down.
    
           The machine is down until further notice....
    
                                                      BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.281                          CMP.A1                           281 of 356
SUBPAC::LANDRY                                       26 lines   7-JUL-1995 17:52
                 -< Table speed/amps problem, help on the way >-
--------------------------------------------------------------------------------
    Here's where it be.........
    
    The polish table bogging down problem has officially been elevated to
    Strasbaugh for them to come in and have a look.  It was quite the
    experience today trying to get some action from them.  After many phone
    calls, many beeps, many hours, we finally got confirmation that an
    abled body will be here Monday morning.  He will also have parts
    shipped here to use in trouble shooting the system.  He'll have at
    least an axis module for the computer and a motor controller/amplifier.
    
    I neglected to mention it to them, but the EMO delay time still exists
    dispite the newly installed 24 VAC relay Strasbaugh suggested we start
    with.  Good time to have them take a gander at it while on site.
    
    Today, a new "old style" Strasbaugh pad was installed along with the
    brushes that were removed to stay wet.  Installed two new carriers
    thanks to BJ and Jim S building them.  Once all was installed, we found
    that a wafer pickup problem existed at both elevators.  So, the "robot
    height adjustment procedure" was performed.  Variable number 130 was
    changed from .23 to .16.  Finally, late first shift, we began to run
    wafers.  They are to be run to try and re-create the bogging down
    problem.  If problem occurs, try and gather as much info from the
    status screen as possible (amps, speeds, pressures, etc).
    
    						Doc  /  Jim T
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.282                          CMP.A1                           282 of 356
SUBPAC::LANDRY                                       24 lines   8-JUL-1995 18:20
                        -< The Tables draggin' me down >-
--------------------------------------------------------------------------------
    Well,  I finally witnessed the polish table stalling senario.  As was
    the case when BJ saw it Thursday night, the following happened/was
    noticed:
    
    	*  Table AMPs went to approximately 30
    	*  Table RPM dropped to 3-4 range (stayed rotating at 3 RPM even 
           after "quitting out" of run mode).
    	*  The spindle down force readings remained at set point (313).  No
           spikes or dips occured.
    	*  Spindle amps:  Left reading ranging from 1.0-1.5 amps.  Right
           reading never moved off of -.1 (???).  It reads this number when
           idle of spinning.  Somethings not right here!
    	*  Back pressure stayed at 14.8
    	*  Slurry flow was at 303 (no restrictions or "pulsing" flow
           noticed)
    	
    	Check the computer connections yet again.  All looked fine. 
    	Hopefully, the patrs arriving can conquer the problem.  
    
    						Doc
       
    By the way, Lightning storms swept through the area this afternoon. 
    Several times the system dumped out to DOS.  Fortunetly, no wafer
    breakage occured during these occurrances.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.283                          CMP.A1                           283 of 356
ASDG::POIRIER                                        27 lines  10-JUL-1995 20:00
         -< Table Servo Amp Replaced - Conditioner Arm Sweep Problems >-
--------------------------------------------------------------------------------
    
    * Table Stalling Problem:
    
      Todd from Strasbaugh arrived this Morning. We replaced the Axis #4
      DC Servo Control Module and the Table Servo Amp. We observed no
      Table rotation failures during a test run of 25 wafers. Since the
      problem has been intermittent in nature, we'll need a few days of 
      production on the tool to be confident with this fix.
       
    
    
     
    * Conditioner Sweep Problem:
    
      Although Strasbaugh has sent us a replacement Conditioning Arm 
      Actuator/Encoder Assembly, Todd recommends that we first replace 
      the "Rod-Ends" that are mounted on both ends of the actuator assembly.
      Both rod-ends are worn out causing excessive play in the assembly. 
      Todd has observed this conditioner problem at other customer sites 
      and replacing the rod-ends solved the problem. The excess play may 
      be causing the Actuator Reed switch disable when the Conditioner
      Arm moves up.  Strasbaugh is sending 2 new rod ends ASAP. In the mean
      time, we'll have to run under warning status.
    
    
    
                                                               John.
5.75CONDITIONER ARM PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:0913
PATE::GEE                                            10 lines  12-JUL-1995 09:21
                      -< Conditioner Arm Motion problems >-
--------------------------------------------------------------------------------
      * Replaced the conditioner arm Actuator/Encoder and repaired a poor
	connection on the Reed switch. This didn't correct the conditioner
        arm sweep problem, it failed to sweep about one out of every five
        times between wet mode cycles. This was checked by running 2 minutes
        of wet mode between pad condition cycling. Ran wafers in auto mode 
        with no conditioner problems.
        Checked the wire connections all the way back to the Axis #14 module,
        inside the computer, they all look good. Swapped Axis #14 module with
        unused module #13. 

5.76CONDITIONER ARM PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:0923
ASDG::POIRIER                                        20 lines  13-JUL-1995 19:21
                  -< Conditioner Problem / Emo Delay Problem >-
--------------------------------------------------------------------------------
    
    Conditioner Arm Sweep Problem:
    
    Strasbaugh was in Today to address this reoccuring problem. The
    servo control module for axis #14 was replaced and the system was
    tested under various conditions without any failures. (Auto cycles
    & conditioning cycles). We are currently performing a verification 
    run. Warnings #'s 1 & 2 have been removed and the status has been
    changed to VERIFY.
    
    
    
    
    EMO Power-Down Delay Problem:
    
    The Strasbaugh rep changed the system main contactor (CON #1) this
    afternoon. Repeated power-down tests revealed no delay. Continued
    testing will take place at appropriate intervals during the verifi-
    cation run tonight.
                       
5.77CONDITIONER DELAYSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 04:0930
ASDG::POIRIER                                        27 lines  10-JUL-1995 20:00
         -< Table Servo Amp Replaced - Conditioner Arm Sweep Problems >-
--------------------------------------------------------------------------------
    
    * Table Stalling Problem:
    
      Todd from Strasbaugh arrived this Morning. We replaced the Axis #4
      DC Servo Control Module and the Table Servo Amp. We observed no
      Table rotation failures during a test run of 25 wafers. Since the
      problem has been intermittent in nature, we'll need a few days of 
      production on the tool to be confident with this fix.
       
    
    
     
    * Conditioner Sweep Problem:
    
      Although Strasbaugh has sent us a replacement Conditioning Arm 
      Actuator/Encoder Assembly, Todd recommends that we first replace 
      the "Rod-Ends" that are mounted on both ends of the actuator assembly.
      Both rod-ends are worn out causing excessive play in the assembly. 
      Todd has observed this conditioner problem at other customer sites 
      and replacing the rod-ends solved the problem. The excess play may 
      be causing the Actuator Reed switch disable when the Conditioner
      Arm moves up.  Strasbaugh is sending 2 new rod ends ASAP. In the mean
      time, we'll have to run under warning status.
    
    
    
                                                               John.
5.100TABLE STALL OUTS STILL WITH USSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 05:549
    
         Table stalling out again due to high current.  Wafers
       were getting out and trashing the retaining rings as well.
       We noticed several items changed in the polish steps from
       the previous shift but since the stalling started two weeks
       ago and is still with us, these changes obviously haven't
       changed anything.  The system is still down in workstream.
                          
                                             D shift
5.101Wafer Fly-out experimentASDG::POIRIERFri Nov 10 1995 19:4715
    
    
    Wafer Fly-out Experiment:
    
    The system continued to fly-out wafers and experience high amperage
    conditions today. These conditions occured even after the force
    ramp time was increaced to 15 sec and the slurry flow was increaced
    to 500 mil/min.
    
    
    D Shift, a Speedfam pad has been mounted on the table. This Pad is
    from an old lot that was shipped to us many months ago. Please build
    2 new carriers and follow the instructions from Eric Bodensieck.
    
                                                             John. 
5.78BRIDGE TRAVELED IN WRONG DIRECTIONSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:478
    
        The bridge was all done polishing and after the carriers 
    lifted off the pad, the bridge moved towards the conditioner
    arm and hit the over-travel limit.  Homing the bridge wouldn't
    move it even after a power down.  With the power down, the bridge
    was moved off of the limit switch manually by rotating the pulley.
    
                                                  BJ
5.79PAD CONDITIONER SWEEP PROBLEMSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:4719
ASDG::POIRIER                                        16 lines  17-JUL-1995 19:19
                     -< Pad Conditioner / Poor Uniformity >-
--------------------------------------------------------------------------------
    
    * Pad conditioner sweep problem:
    
      Control modules for axis # 14 & 15 were changed in the PC last
      Thursday. There were no failures until this morning at 6:00 a.m.
      The system was re-booted and no failures have been observed since.
      Strasbaugh was contacted and have recommended changing the motor
      controller board for the conditioner. They will ship a replacement   
      ASAP.
    
    
    * MQC Failure:
    
      The system failed MQC's this morning for poor uniformity. The pad was
      replaced, conditioned & break-in wafers run. MQC's are now being run. 
     
5.80ROBOT LEAD SCREWSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:488
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               5 lines   5-AUG-1995 00:30
                                -< Lead screw >-
--------------------------------------------------------------------------------
    
         The robot wasn't making it to the right load cup.  
      After the lead screw was cleaned and lubed, it traveled
      fine.
                                           BJ and Doc
5.81ROBOT LEAD SCREW TRAVELSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:4827
ASDG::POIRIER                                        24 lines  14-AUG-1995 20:00
      -< Robot Leadscrew sticking/send 1 & recieve 1 elevators off-line. >-
--------------------------------------------------------------------------------
    
    * The robot leadscrew assembly has been sticking in the same
      manner as CMP.A2's did last week. We have therefore changed
      calibration variable #105(Rec'v 1 robot position) to -35.5
    
      This makes Send Elevator #1 & Recieve Elevator #1 unusable
      until the leadscrew position is changed back. The leadscrew
      assembly will be removed, cleaned & readjusted as soon as
      we have an opportunity to do so. This will most likely be
      after completion of the Speedfam upgrade activities.  
    
      The Strasbaugh rep has suggested that we reduce the length of
      Robot bellows on the Elevator end of the leadscrew assembly.
      The current length is unecessarly long and only compounds
      our problem with excess friction in the assembly. The 
      bellows ends-up getting compacted tighly between the Robot
      base and the side of the machine base.
      
    
                                                           John.
    
     
      P.S. The old (rec'v 1 robot postion) has been stored in an
           unused variable, #109. 
5.82CONDITIONER TRAVEL/SWEEP PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:4836
ASDG::POIRIER                                        33 lines  15-AUG-1995 20:30
                     -< Conditioning Arm Position Changes >-
--------------------------------------------------------------------------------
    
    * Conditioner "not arriving at center position" problem:
    
      This problem was investigated with the Strasbaugh rep 
      this afternoon. The following problems were discovered:
     
      1) The Conditioner Assembly bracket screws were dragging
         on the assembly sleeve during its up/down motion.
         This was causing an over-current condition with the 
         motor which halted the movement of the conditoner axis.
    
     2) An engineering update, published by Strasbaugh concerning 
        this issue, was never sent to us. It requires entering an  
        additional command into the (init1.dan) file.
    
      
      Action Taken: 
    
        Tony entered the required command into the init1.dan file.
        This allows a negative value to be entered into the conditioner
        arm position variable. The park station position is now -2500
        instead of o. The center and edge positions are also 2500 less
        than they were previously.
    
        The conditioner assembly sleeve was rotated slightly to permit
        clearance on both sides of the bracket. The end result is that 
        the conditioning disk now postions itself 2500 counts off center
        in the park station resivoir. However, this reduction in sweep
        allows the conditioner assembly to travel in both directions
        without any mechanical iterference with the bracket.
    
    
                                                               John.
5.83HIGH PARTICLESSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:4911
PATE::GEE                                             8 lines  16-AUG-1995 08:41
                       -< High particles / brush change >-
--------------------------------------------------------------------------------
	Failed MQC's for high particles after a pad change.

	We found that the load cup guide rings and the PVA brushes
	had a black substance on them. We couldn't remove it so
	we replaced the brushes and load cup guide rings. We also
	cleaned everything inside the system with IPA wipes. 
	
	Ran a second particle check and it passed.
5.84CONDITIONER DISK SIZE CHANGEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:4911
PATE::GEE                                             8 lines  22-AUG-1995 07:26
                  -< 4" Conditioning Disk / variable changed >-
--------------------------------------------------------------------------------
	We replaced the conditioner disk assembly with the 4" type
	(no 6" in stock), do to poor performance. 
	There has been a rush order placed on the 6" conditioner
	disks since last week.
	We had to change calibration variable #126 (center position)
 	from 113000 to 108000 and #127 (edge position) from 139750
	to 142000.
					Roger & Bob
5.85CONDITIONER ARM AIR CYLINDER LINESSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5113
ASDG::POIRIER                                        20 lines  22-AUG-1995 18:58
    
    * Conditioning Arm Air hose failure:
    
      The air line that supplies the Conditioner down force air cylinder  
      blew apart today. A union fitting was spliced-in as a temporary
      repair. Strasbaugh recommends using 1/4 " Poly-Flow hose. Teflon
      can also be used but is not necessary. The use of Teflon will also
      require replacing all the fittings. The repair will require approx 10
      ft of hose. It begins at the air cylinder under the conditioning
      assembly and terminates at the air cylinder at the end of the
      conditioner arm. 
       
5.86LEAK AT ROTARY JOINTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5260
PATE::GEE                                            11 lines  23-AUG-1995 07:16
                               -< Rotary joint  >-
--------------------------------------------------------------------------------
	Found a water leak at the top of the left spindle rotary
	joint, looks like the top inner seal is bad. It only leaks
	when the wafer blow-off occurs. This needs to be monitored 
	until time can be scheduled with production to fix it.
	
	The front spindle housing cover has been re-installed to
	prevent any contamination from reaching the table.   
	
	We need around 6 hours to replace both left and right spindle
	rotary joint.
					Leon & Bob

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.305                          CMP.A1                           305 of 356
ASDG::POIRIER                                        38 lines  23-AUG-1995 19:41
             -< Rotory Joint Replacement / Conditioner Disk Mount >-
--------------------------------------------------------------------------------
    
    * Left Spindle Rotory Joint:
    
      The Left Spindle Rotory Joint was replaced this afternoon. This was
      the first time we've replaced one. The replacement joint was a newer
      rev than the joints that came with the tool. Also, we experienced
      problems with the joint that we first attempted to install. The 
      tappered pipe threads in the mounting nut were not tapped to the 
      correct depth. Fortunately, we had more than one in stock and
      were sucessful with the 2nd joint we tried. 
    
      Be careful of the following items if you install a new Rotory Joint
      on the Strasbaughs:
    
      1) The new rev joint is a 2 stage design and is much higher. Make
         sure the back pressure transducer doesn't interfere with the 
         sheet metal air filter supports on the top of the tool. Move
         the Bridge manually to ensure clearance.  
    
    
      2) The Barbed fitting that protrudes from the Bottom of the Spindle
         Quill has been replaced with a 90 deg plastic fitting. The new
         Rotory Joint pipe is slightly longer than the old style.
    
    
      3) Do not reuse the old pipe it, is too short. The new pipes are
         in the bottom draw of Jim Tooker's file cabinet.
                                     
    
    
      Conditioner Disk Mount:
    
      The Spacer Sleeves in the Conditioner Disk Mount were replaced with
      slightly shorter ones to allow better gimbling of the assembly.
      This was done to reduce the rotational hesitation we've been seeing
      as the disk rotates on the Pad.
    
                                                     John, Tim & Brendan.
5.87BRIDGE VACUUM FILTERSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:528
FABSIX::R_POIRIER                                     5 lines  29-AUG-1995 23:37
                           -< Bridge vacuum filters >-
--------------------------------------------------------------------------------
    Replaced Bridge Vacuum Filters in CMP.A1 during pad change.
    
    
    						Bob,Roger,BJ
    
5.88EMO CONTACTOR REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:529
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"               6 lines   8-SEP-1995 07:06
                        -< Loud EMO contactor replaced >-
--------------------------------------------------------------------------------
    
        Replaced the EMO contactor with a new one.  The old one
      is on the tool box.  The new one does not hummmmmmm at this
      time.
    
                                         BJ
5.89RMP'S IN QUESTIONS, CHECKED OUTSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5315
SUBPAC::LANDRY                                       12 lines  13-SEP-1995 15:01
             -< rotation speeds confirmed, slurry filter changed >-
--------------------------------------------------------------------------------
    While waiting on new pad shipment, we checked both spindles for proper
    spin speed.  23 RPM was programmed, 23 RPM was the tach readout on both
    spindles.  The table speed was also checked.  Again, 23 was the
    programmed speed andd 23 was the tach readout.
    
    Changed the pimary slurry filter, as the old one was clogged.  Just to
    make note of, when this filter clogged, the pressure gauges on the
    other systems rose about 10 PSI.  Makes sense, seeing they are all on
    the same feed loop.
    
    						Doc  /  JP
    
5.90ELEVATOR NOT HOMINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5564
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              22 lines  21-SEP-1995 23:31
    
                 -< Elevator and robot problems, PM completed >-
--------------------------------------------------------------------------------
    
    
         The left SEND elevator wasn't homing.  At first, since we
      had the problem last week with the encoder, we were going to 
      go right after the encoder.  When we pulled the elevator away
      and discovered the 3/8" line to fill the tank was all crimped
      and sitting near the home switch.  It looks like the line was 
      caught under the elevator and was blocking the it rom getting 
      to home.
        
           The robot wasn't picking wafers from the right SEND.  The
      Axis 6 to slot zero parameter was changed to -3.07. Everything
     ran fine after that.
    
    
      ORIGINAL PROBLEM---     -< LEFT SEND ELEVATOR dead >-
--------------------------------------------------------------------------------
    
        DI water has once again caused problems.  The left send
     elevator encoder appears to be the culprit after (3) hours
     or checking things out.  The elevator and tub are on the table
     in the chase for ENG inspection.  
    
================================================================================
Note 55.318                          CMP.A1                           318 of 356
ASDG::POIRIER                                        26 lines  22-SEP-1995 19:57
      -< Left Send Elevator Motion Problem / Right Send Tank Replacement >-
--------------------------------------------------------------------------------
    
    * Left Send Elevator Motion Problem:
    
      The Left Send Elevator Motor, Encoder & related wiring were removed
      from the tool and allowed to dry-out overnight. The corroded
      connectors & lugs were cleaned and re-tightened. The Elevator
      assembly was then reinstalled and tested. Its working fine now.
      Leon took the motor apart and cleaned the rotor & stator. They
      were rusted together.
    
      It appears that D.I. Water was splashing onto the Elevator Mounting
      Base and running over to the Elevator Motor, Encoder & Connectors.
      The Elevator Tank Drain Hoses were extended to prevent water from
      splashing on the Elevator Mounts in the future.
    
    
                                                         John/Leon/Brian
    
    
    
    
    * Right Send Elevator Tank:
    
      The cracked Tank was replaced this afternoon. While the elevator was
      out, corroded connectors were found on the assembly. They were
      cleaned and retightened.  
                                                        
          The outstanding intems for the annual including the slurry
    lines and robot positioning was completed.   The annual PM has 
    finally been completed.
    
                                                 D - shift 
5.91ROBOT WAFER PICKUP PROBLEMSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5642
ASDG::POIRIER                                        11 lines  13-OCT-1995 18:48
                          -< Robot pick-up problems >-
--------------------------------------------------------------------------------
    
    * Robot Sensor Adjustment:
    
      The Robot experienced wafer pick-up problems around 5:00 this
      afternoon. The sensor was readjusted and a verification run was
      completed. It seems to be working fine now. We're not sure why
      the sensor is continually going out of adjustment. we'll have
      to keep a close watch on the Robot for the time being.
    
    
                                                              John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.340                          CMP.A1                           340 of 356
FABSIX::R_POIRIER                                     7 lines  15-OCT-1995 20:36
                        -< Robot Pick-up problems (A1) >-
--------------------------------------------------------------------------------
    * Robot Sensor Adjustment
    
      Had problems with wafer pick-up on A1 Saturday nite. Re-adjusted
      sensor and ran fine after that.
    
    						B Shift
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.341                          CMP.A1                           341 of 356
FABSIX::R_POIRIER                                     5 lines  16-OCT-1995 04:16
                        -< Robot End -effector CMP.A1 >-
--------------------------------------------------------------------------------
    * Robot Sensor Adjustment
    
      The problem occured again tonite. Re-adjusted sensor and
      robot is working fine since.
    
5.92TABLE STALLINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5617
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              14 lines  20-OCT-1995 09:03
                       -< table stalled, wafer breakage >-
--------------------------------------------------------------------------------
         The pad was replaced earlier in the shift and was broken
      in using the BREAKIN recipe.  As soon as wafers were started
      on the tool for MQC's, the table stalled out during polishing 
      after the amps got over 40.  The system was EMO'ed before 
      anything crashed.  we started the tool back up and we started 
      to run some wafers to see if the table would stall again.  As
      we were running the tool, a wafer under the left spindle broke
      THEN slid out.  The table did not over current.  The pad was 
      replaced and Barry break-in the original way.  He did not
      experience any problems.  MQC's are being run and the machine 
      is being closely watched.
 
                                                    D shift
       
5.93TABLE STALLING AGAINSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5650
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              14 lines  21-OCT-1995 06:50
                           -< Table stalled out.... >-
--------------------------------------------------------------------------------
    
          The table stalled out but Gwen saved it from crashing
        by EMO'ing the system.  12 wafers were without any problems.
        New carriers were installed due to poor vacuum problems so
        MQC's needed to be done.  The MQC wafers ran fine but when
        (4) sender heads were sent, the table stalled.  There was 
        a 15 minute wait between the MQC's and the sender heads.
        For now, run (4) send heads before you run product.  Have
        the 1st two wafers set at 200 down force and the 2nd two 
        at 313.  The last time we started seeing the stalling problem,
        the Reps suggested looking at a bigger ratio in the gear box.
        I will speak with John P. in the a.m. regarding this....
    
                                                     BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.351                          CMP.A1                           351 of 356
ASDG::POIRIER                                        25 lines  26-OCT-1995 19:32
                    -< Power-Down Problem / Wafer Breakage >-
--------------------------------------------------------------------------------
    
    * Mysterious Power-off Problem:
    
      The system powered itself down 3 times this morning while idle
      and once this afternoon while polishing. I checked the EMO circuit
      including all door interlock switches and all 4 EMO button contacts.
      All connections were o.k. 
    
      Strasbaugh field service was called and they advise checking the
      following after the next incident:
    
      a) Main Contactor (CON #1) Coil may be bad.
    
      b) 24v relay in EMO circuit. 
    
      c) check continuity on all door interlocks and EMO buttons.
    
    
    
      Wafer breakage:
    
      The system just broke wafers during the verification run for
      the power-down problem. The status should be changed to VERIFY 
      after the Pad and Carriers are replaced.
    
5.94CONDITIONER PROBLEMSSUBPAC::BJUBINVILLENFD Fire - RescueFri Nov 10 1995 23:5736
SUBPAC::BJUBINVILLE "NFD Fire - Rescue"              11 lines   1-NOV-1995 05:30
                     -< Conditioner assembly disassembled >-
--------------------------------------------------------------------------------
    
    
          The roll pin that connects the drive shaft to the 
      white ring which in turns connects to the conditioner
      disk worked its way out and ripped the bellows apart.
      The bellows are not a stocked item so at least one will
      have to be rushed in.  The machine is still usuable but
      the bellows should be put back on the machine asap.
      The Strasbaugh 24-hour hotline was called but I got no
      answer.  
                                           D shift

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.356                          CMP.A1                           356 of 356
ASDG::POIRIER                                        14 lines   3-NOV-1995 19:47
       -< Conditioner Bellows Replacement / Cond Drive Belt Idlers Adj. >-
--------------------------------------------------------------------------------
    
    Conditioning Disk Bellows:
    
    The replacement bellows arrived this morning and was installed
    on the Conditioning Arm Disk Assembly.
    
    
    
    Conditioning Disk Hesitation:
    
    A hesitation was observed as the Conditioning Disk was rotating. It
    seemed to be getting worse over the last few weeks. We adjusted the
    idlers on the Conditioning Arm Belt Assembly an the hesitation ceased.
    
5.1A1 FIT-UP AND MECHANICAL ACCEPTANCESUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 11 1995 00:44748
================================================================================
Note 55.1                            CMP.A1                             1 of 356
SUBPAC::LANDRY                                       12 lines   6-JUN-1994 16:34
                                -< 6/6 update >-
--------------------------------------------------------------------------------
    Monday Installation Update...........
    
    There wasn't much activity today.  I got with the contractors (KSI & 
    NEVCO) and clarified a couple of questions they had concerning plumb-
    ing (DI & GCW) and slurry supply and return.  Also, since the systems
    had moved a little due to the pedistal installation, they wanted to
    reverify with me where they could and could not run the hookups.
    
    They are in the process of running pipe under the floor at this time.  
    Not much to see! There was no other activity with this system today.
    
    					Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.2                            CMP.A1                             2 of 356
SUBPAC::LANDRY                                        8 lines   7-JUN-1994 15:53
                                -< 6/7 Update >-
--------------------------------------------------------------------------------
    	Today, KSI hooked up the two GCW drains to the system.  The also
    hooked up the GCW drain on the vacuum cabinet resevoir.  Both of these 
    tasks required substantial installation work under the floor (running
    headers and sticks).  Don't be alarmed as you look through the raised 
    floor and notice some water, they had to cut into an existing line that
    contained some water.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.3                            CMP.A1                             3 of 356
SUBPAC::LANDRY                                       18 lines   9-JUN-1994 15:40
                                -< 6/9 Update >-
--------------------------------------------------------------------------------
    Today, KSI has the main system and vacuum cabinet drain lines under 
    pressure testing.  Alls well so far!
    
    KSI had a question concerning the slurry supply and return line hookup.
    They were under the assumption that the incoming and outgoing lines
    needed to be double contained plumbing.  After confering with George
    Hynes and Tom Richards, it was determined that the lines would NOT have
    to be double containment.  This will help this portion of the install
    go alot quicker.  The double contained wall would have housed a supply
    and a return line within each of the two pump hookups.  They would have 
    had to fabricate a manifold between the slurry valves and the valve
    manifold box to separate the two prior to entering/exiting the VMB's in
    order to maintain double containment integrity.
    
    Also, the electricians are beginning to address what is needed on the
    system.  Some conduit has been run (minus the wires).
    
    						Doc 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.4                            CMP.A1                             4 of 356
SUBPAC::LANDRY                                       18 lines  16-JUN-1994 15:49
                                -< 6/16 Update >-
--------------------------------------------------------------------------------
    Installation status:
    
    The electrical hookup is complete.
    
    The GCW drains have been leak/pressure checked to satisfaction.  The
    lines were then connected up to the main system.
    
    The TCU installation has begun.  There is a rack mounted that houses
    the incoming/outgoing process cooling lines.  They still need to be
    tied into the TCU.  Also, need the hookup from the TCU to the main
    system to be performed.  The electrical hookup is complete.
    
    The vacuum cabinet has the GCW drain hooked up.  The electrical is also
    complete.  Waiting for the CDA hookup to be performed.
    
    No slurry hook ups have been performed as of yet.
    
    						Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.5                            CMP.A1                             5 of 356
SUBPAC::LANDRY                                       15 lines  24-JUN-1994 14:05
                         -< 6/24 Installation Update >-
--------------------------------------------------------------------------------
    KSI personnel ran and hooked up the slurry lines on the system today.
    The lines are a single PVA hook up (NO secondary containment plumbing
    is required).  The lines are supported by a unistrut rack.
    
    KSI also is making good progress with the GCW drains.  The main system
    hookup is COMPLETE!  The vacuum cabinet drains are COMPLETE as well!
    Plumbing work for the TCU is on-going.
    
    The system still needs the following hook ups:
    
    	CDA, UHP DI, Cooling water in/out, vacuum cabinet tied into the
    	main system, vacuum cabinet CDA, ALL TCU plumbing requirements.
    
    						Doc
     

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.6                            CMP.A1                             6 of 356
SUBPAC::TOOKER                                       37 lines   8-JUL-1994 10:59
                         -< Hook-Up Status for 7/8/94 >-
--------------------------------------------------------------------------------
                CMP.A1 Tool Hook-Up Status For 7/8/94



COMPLETED ITEMS:
----------------

    *  Bulkhead 

    *  DI Water connection except for pressure gauge.

    *  All GCW drain connections

    *  Eelectrical connections

    *  NESLAB TCU soft water, process cooling water, and drain
       connections

    *  Slurry line from valve manifold box to system


OPEN ITEMS:
-----------

    *  DI Water pressure gauge

    *  DI water return line final connection

    *  CDA final connection to main system and vacuum cabinet

    *  N2 final connection

    *  All exhaust work

    *  HEPA filter connection

    *  Slurry valve manifold box valve work

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.7                            CMP.A1                             7 of 356
SUBPAC::LANDRY                                        6 lines  12-JUL-1994 13:39
                         -< 7/12 Installation Update >-
--------------------------------------------------------------------------------
    Today, the three HEPA filters for the system were installed.  This
    will allow NEVCO to begin installation of the AC hookup to the top of
    the system.  Thanks to Jim and Earl for their astute assistance in this
    installation!
    
    							Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.8                            CMP.A1                             8 of 356
SUBPAC::TOOKER                                       11 lines  26-JUL-1994 09:17
                             -< CMP.A1 punch list >-
--------------------------------------------------------------------------------
                CMP.A1 Tool Hook-Up Status As Of  7/25/94


PUNCH LIST ITEMS
----------------

             ITEM                                      RESPONSIBLE PERSON
   --------------------------------------------------------------------------
  1) Complete exhaust work.                            D. Atkinson/NEEVCO

  2) Uni-strut guard around plumbing area.             D. Atkinson

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.9                            CMP.A1                             9 of 356
SUBPAC::BJUBINVILLE                                   7 lines  29-JUL-1994 15:46
                  -< CMP.A1 is all set for the vendor on 8/1 >-
--------------------------------------------------------------------------------
   
        Doc and I completed N2, CDA and vacuum facility hook-ups
    this afternoon.  It's all set for the vendor on Monday.  We are
    hoping to be able to pick up the drip pans and install them on
    Monday.
    
    BJ

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.11                           CMP.A1                            11 of 356
ASDG::POIRIER                                        64 lines   4-AUG-1994 20:51
                            -< CMP.A1 Mech Accept >-
--------------------------------------------------------------------------------
    
    Power, vacuum, CDA and D.I. water have all been supplied to the tool.
    The D.I. water pressure is still too low. Its currently at 25 p.s.i.
    up from 15 p.s.i. yesterday but still not at the required 40 p.s.i.
    
    There is currently no slurry available. We're hoping the slurry system
    will be operational by next Tuesday for the planned 500 wafer run.
    If not, then we will perform the 500 wafer run with water only. The 
    limitation of doing this is "touching down" on the polishing pad
    momentarily instead of oscillating on the pad.
    
    
    Here is a highlight of some of the work completed thus far on CMP.A1
    
    * The table, spindles and chucks have been indicated and adjusted to the
      proper specifications.
                  
    * The shuttle has been leveled and adjusted in both the forward and
      rear positions. 
    
    * The shuttle speed has been adjusted.
    
    * The shuttle hard stop positions have been adjusted.
    
    * The load stations and cleaning stations have been set to the proper 
      positions.
    
    * Calibration values have been set for Table Min Pressure, Spindle Down
      Soft and Spindle Down Hard.
    
    
    Plans for tomorrow:
    
    * Adjust spindle down force.
    
    * Adjust conditioner down force.
    
    * Calibrate slurry PH meters.
    
    * Calibrate robot positions.
    
    * Calibrate Elevator positions.
    
    * Install elevator tubs.
    
    * Cycle machine.
    
    
    Problems:
    
    * The load cell that was shipped with the machine doesn't work. The rep
      went to Conn. tonight to get another one.
    
    * Some of the poly welds cracked during shippment.
    
    * Rust was discovered in the weldment area.
    
    * The Thompson bar way bearings that support the robot are noisy.
    
    * The lead screw bearing on SEND elevator #2 is binding.
                                            
                                          
                                                      
    * 

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.12                           CMP.A1                            12 of 356
ASDG::POIRIER                                        17 lines   5-AUG-1994 17:51
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
   * We attempted to calibrate the down force on both spindles today.
    
   * The spindle lever arms were not adjusted properly at the factory
     which caused the air cylinder housings to contact bridge box frame.
     This in turn put a side load on the teflon air cylinder bushings which
     caused them to move out of they're proper locations. 
     
   * The left spindle air cylinder was replaced with the spare that came with 
     the machine. 
    
   * The spindle lever arms have been adjusted to the proper specs. This
     will prevent this situation from occuring in the future. 
    
   * Four new Air cylinders were shipped today from Strasbaugh. We'll replace
      them on Monday morning.   
    
   

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.13                           CMP.A1                            13 of 356
ASDG::POIRIER                                        36 lines   8-AUG-1994 22:48
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The four spindle air cylinders that were ordered last Friday arrived
      this morning. One cylinder was pulled from stock on Friday and
      installed on the left spindle. One of the four that arrived today was 
      fitted with a load cell and mounted on the right spindle assembly.
      Down force calibrations were then sucessfully performed on both spindles. 
    
    * Down force calibration was also performed on the pad conditioner
      assembly.
    
    * All four elevator tubs were installed and are flowing D.I. water.
    
    * Both left and right spindle hub o-rings have been replaced with
      the new silicone style o-rings.
    
    
    Problems:
    ---------
    * The A23 pad conditioner down force output was incorrectly labled on
      the print. Strasbaugh was notified and will send a corrected print.
    
    * The teflon tubing which was disconnected from the elevator tubs would
      not fit onto the the valve fittings. Strasbaugh sent a flaring tool
      and we resized the tubing.
    
    * Broken poly welds were discovered near the send elevators. Strasbaugh 
      is sending a small hand held poly welder so we can rework the welds.
    
    * The D.I. water pressure continues to be a problem. The water pressure
      at the Speedfam tool holds steady at 35 p.s.i., we're only 10 ft down
      line and the pressure at the CMP.A1 fluctuates from 5 to 20 p.s.i.
      We need 40 p.s.i. of D.I. water pressure at CMP.A1.  This will be a 
      gating item for the planned 500 wafer run tomorrow. We can't touch-
      down on the pad without a sufficient amount of water flowing.
      We could most likely get by with 20 p.s.i. for tomorrow's 500 wafer
      run as long as the pressure holds STEADY.   

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.14                           CMP.A1                            14 of 356
ASDG::POIRIER                                        52 lines  10-AUG-1994 09:29
                           -< mechanical acceptance >-
--------------------------------------------------------------------------------
    
    * The status of CMP.A1 in RTC has been changed from "red tag sign off"
      to "mechanical acceptance". All other CMP tools have been changed 
      as well.
    
    * The noisy robot travel problem has been solved. The sheet metal
      mounting bracket for the robot servo motor was loose. This forced
      the lead screw coupling to bind and caused the noise. The bracket
      has been tightened and the robot is running quiet now.
    
    * The shuttle support brackets have been replaced with the new stronger
      ones that arrived from Strasbaugh on Monday.
    
    * The robot to elevator locations will be re-checked with the new 
      Empak cassettes. The tool was origionally set-up with pre-production
      sample cassettes in January. The material of the new cassettes is
      slightly different from the earlier ones and may have a different rate 
      of shrinkage thus different distances between slots. 
    
    * The machine was cycled and a 25 wafer run was completed with 1
      assist. The spindles did not lift up off the pad when they were
      supossed to. The problem was corrected by changeing the "Min Lift
      Off" variable.
    
    
    * Fine tuning needed before 1000 wafer run:
    
      1) Set motor spindles so they operate together. They don't touch down
         or lift off at exactly the same time.
    
      2) Fine tune load station.
    
      3) Fine tune the conditioner functions.
    
      4) Install new version of software.
    
    
     * Problems:
      
      1) Neoprene door gaskets are already tearing.
      
      2) Water is accumilating on the poly deck behind the recieve stations
         and dripping on the floor. 
    
      
      Misc issues:
    
      1) We could really use an access door behind the recieve elevators.
         This tool was not origionally designed for a bulkhead 
         configuraton. This leaves no access to the send/recieve elevators
         from the service chase. I recommend we repace the plexiglass
         window behind the recieve elevators with an access door.  

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.15                           CMP.A1                            15 of 356
ASDG::POIRIER                                        35 lines  11-AUG-1994 02:50
                                    -< ex >-
--------------------------------------------------------------------------------
    
    * The 500 wafer run was completed at 2:30 a.m. with no operator assists.
    
    
    The following items will be addressed prior to Monday's 1000 wafer run:
    
    1) Right load station ring doesn't fit properly.
       Strasbaugh will ship new one tomorrow.
    
    2) Right carrier spray nozzel hose is too short.
       New hose will be installed tomorrow.
    
    3) Left load station cylinder has a slight hesitation. 
       It will be cleaned and inspected.
    
    4) Left recieve elevator is making noise.
       We will clean and inspect the lead screw bushing.  
       Lube if necessary.
    
    5) Brush height on both cleaning stations is to high.
       Height will be adjusted.
    
    6) Elevator locations need to be set to four new Empak cassettes.
       Currently set to pre production cassettes.
    
    7) The down force count calibration needs to be reset.
       This ensures that the spindles come down evenly. This calibration
       will not effect the actual down force during the polishing cycle.
    
    8) The new version of software needs to be down loaded.
    
    9) The leak in the poly shroud needs to be sealed.
       We'll seal it with RTV.
   
    * We are ready to polish as soon as slurry becomes available.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.16                           CMP.A1                            16 of 356
ASDG::POIRIER                                        27 lines  11-AUG-1994 20:40
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The new load station guide rings arrived today and were installed in
      both CMP.A1 & A2.
    
    * The pad conditioner arm was fine tuned. The uneven vertical motion
      turned out to be a slight mechanical interference in the drive
      mechanisim. Eric and I adjusted the mechanisim and the motion is 
      now smooth. The home position location was also changed, this allows
      the conditioning head to rest in the center of the park position
      tank. 
    
    
      Misc items that need attention:
    
      1) Finish leveling machine.
     
      2) Install containment trays under tool.
    
      3) Have 11" dia plexiglass plates fabricated. These plates will fit
         into the load stations and act as fixtures.They will aid in setting up
         the brush heights. They will also provide a convenient surface to 
         place the indicator base while indicating the chucks. This is
         especially helpful when a fresh pad is on the table.
    
    
       * (1000 wafer run in planned for Monday)
                                               

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.17                           CMP.A1                            17 of 356
ASDG::POIRIER                                         6 lines  15-AUG-1994 22:50
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
  * The 1000 wafer run was postponed due to a malfunctioning robot (upper
    arm) motor encoder. The encoder was replaced and the robot is now
    functioning normally.
    
  * A second attempt at the 1000 wafer run will comence tomorrow.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.18                           CMP.A1                            18 of 356
ASDG::POIRIER                                        15 lines  16-AUG-1994 21:39
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The 1000 wafer run has been delayed once again. This time because of  
      robot and conditioner problems. Replacing the robot assembly last night 
      rendered most of the elevator locations to be incorrect. All of the 
      locations have been corrected except for the elevator Z-heights. That
      work should be finished-up in the morning. 
    
      The other time consuming hit was the conditioner arm. It died at 8:00
      p.m.. Upon disassembly, a broken weld was discovered on the main
      shaft bushing. This bushing is welded to the shaft near the top of
      the conditioning arm. Strasbaugh was called and a new shaft/bushing
      assembly will be shipped tomorrow. We expect the assembly to arrive
      Thursday morning.
        
                                                                 J.P.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.19                           CMP.A1                            19 of 356
ASDG::POIRIER                                        12 lines  17-AUG-1994 21:59
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The elevator Z-height postions for both the send and recieve
      elevators were calibrated today. These locations were lost 
      when the robot was dis-assembled to replace the servo encoder.
    
    * A 300 wafer run was completed at 9:00 p.m. with no assists. 
      This test was run in preparation for tomorrow's 1000 wafer
      marathon run. The Strasbaugh rep will start the run at 8:00 a.m.
      and I will finish it at approx 12:00 midnight. (16 hrs)
    
    
                                                               John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.20                           CMP.A1                            20 of 356
SUBPAC::LANDRY                                       15 lines  18-AUG-1994 15:33
                            -< Almost slurry time! >-
--------------------------------------------------------------------------------
    
    Slurry happenings today.......
    
    A decision was made to have one supply and one return line supply both
    pumps.  Therefore, one supply and one return line were removed from the 
    system/rack/VMB.  We (Earl and myself) then cut into the remaining two
    lines and "tee'd" off them.  
    
    Also, the VMB connection was made to complete the slurry line
    installation.  We are now ready for slurry, pending the room eye wash /
    safety shower signoff.  Said signoff work is in process.
    
    Thanks go out to Earl for his support and work with this project.
    
                                                  Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.21                           CMP.A1                            21 of 356
ASDG::POIRIER                                        12 lines  19-AUG-1994 03:44
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The 1000 wafer marathon run was sucessfully completed at 3:32 A.M. 
      No assists were necessary.
    
    * Power, vacuum & D.I. water were turned off.
    
    
      MISC:
    * I checked the slurry valve boxes for leaks during 2nd shift. No leaks
      were found. 
    
                                                                    John.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.22                           CMP.A1                            22 of 356
SUBPAC::LANDRY                                        9 lines  19-AUG-1994 16:11
                   -< nice try....but totally unacceptable >-
--------------------------------------------------------------------------------
    
    *  The Neslab TCU has been found to be improperly plumbed up.  The
       resevoir drain was hard-plumbed without including a shutoff valve.
       This was discovered as we attempted to fill the resevoir.  The
       ethylene glycol ended up going straight to the GCW drain.  The 
       problems has been addressed with the proper people to have the
       situation fixed.  Stay Tuned!!!
    
                                                 Doc

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.23                           CMP.A1                            23 of 356
ASDG::POIRIER                                         7 lines  19-AUG-1994 22:37
                                 -< Leveling >-
--------------------------------------------------------------------------------
    
    * CMP.A1 was leveled. The design & layout of the six leveling feet
      leaves a lot to be desired. What a pain in the neck, back and just
      about everywhere else.
    
    
                                                                       J.P.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.24                           CMP.A1                            24 of 356
ASDG::POIRIER                                         8 lines  22-AUG-1994 12:09
                        -< Slurry is flowing in CMP.A1 >-
--------------------------------------------------------------------------------
    
    FYI: Slurry is flowing in CMP-.A1 as of 12:00 noon today. 
         Please be aware of the protocol restrictions now that 
         slurry is flowing in the tool.
         
    
                                                            John.
    

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.25                           CMP.A1                            25 of 356
ASDG::POIRIER                                        33 lines  22-AUG-1994 18:54
                           -< mechanical acceptance >-
--------------------------------------------------------------------------------
    
    * The left spindle air cylinder was replaced today. The reason was that 
      the spindle would not retract the last 1/4 " of its travel. The
      cylinder replacement required us to recalibrate the spindle down force
      as well.
    
    * The conditioning arm assembly was reinstalled after the replacement 
      of the mounting shaft. This shaft suffered broken a weld last
      Thursday. All adjustments and locations were set today.
    
    * We opened the valve in the slurry valve box and slurry is now flowing
      in the tool. The slurry flow rate was also calibrated today.
    
    * The send and recieve elevators were recalibrated. This was done
      because the wafers were dragging out of slots # 23, 24 & 25.
      Since we are mounting the cassettes H-BAR up, I'm not sure if were
      referencing from a surface on the cassette that is held to tight 
      tolerences. Also, since the problem seems to come and go, I'm not
      confident that the cassettes are uniform to each other. I'll take a 
      look at the Empak print when I get a chance.
    
     Plans for tomorrow:
     -------------------
     1) Change polishing pad.
    
     2) Install latest rev software.
    
     3) Start setting up CMP.A2.
    
     F.Y.I.:
     ------
     A new phone has been installed in the CMP service bay, the # is 5804.
                                                          

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.26                           CMP.A1                            26 of 356
ASDG::POIRIER                                        22 lines  23-AUG-1994 18:39
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    
    * The conditioner arm travel was re-adjusted. We're trying to set up
      IN-SITU conditioning. We need more travel to cover the wafer track.
      Strasbaugh is sending us a 6" conditioning disk to replace the 4"
      one we are currently using.
    
    * The process folks polished wafers for the first time this afternoon.
    
    * We discovered that the suction between the wafer and the pad is strong
      enough to prevent the spindle from lifting under certain conditions.
      We may be able to solve this problem by changing our process
      parameters or maybe Strasbaugh can come up with a hardware change 
      on the tool. Stay posted. 
    
    Planned activities for tomorrow:
    
    1) Calibrate slurry PH meters.
    
    2) Replace conditioning disk.
    
    3) Polish wafers.

          <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
                                -< Films Only! >-
================================================================================
Note 55.27                           CMP.A1                            27 of 356
ASDG::POIRIER                                        41 lines  24-AUG-1994 20:11
                           -< Mechanical Acceptance >-
--------------------------------------------------------------------------------
    
    * The cleaning station brush heights were checked today using the new
      lexan fixture made by Al in the carpentry shop. It worked great.
    
    * The air brush deflector that blows air onto the pad does not work
      very well. Gregg Myers came up with an idea for a new one with an 
      angle instead of the current flat plate. I'm going to ask Al in the 
      carpentry shop to fabricate us a prototype.
    
    * The new 6" dia conditioning disk arrived today. It does allow 
      full coverage of the wafer track on the pad but not without 
      a few problems. The increased friction of the larger disk puts
      additional strain on the conditioning arm. This increased strain
      is causing the base of the machine, where the conditioner mechanisim
      is bolted, to flex approx 1/8". This 1/8" of movement translates into
      2" of oscillation at the conditioning disk which is cantilevered over 
      2 feet from the base. If it is determined that the 6" disk is going
      to be used (it's needed for IN-SITU conditioning) then the base will
      have to be strengthened with a support bracket or plate.
    
    * Do to a verbal miscommunication during a software adjustment, we
      suffered of minor crash. The bridge plowed into the conditioning arm.
      The arm did not get the full force of the bridge, thank goodness. If 
      it did there would be nothing left to it. An alert individual hit the
      EPO with just seconds to spare. This allowed time for the bridge to
      slow down before impact resulting is a small bend in the conditioner
      arm. The arm was disassembled, reshaped and reassembled. We then
      made all the necessary adjustments, both software and mechanical, and
      ran a test run. The conditioner is now working fine.
    
    * The problem with the spindles not retracting UP has been identified.
      The IC 1000 pads combined with an 8 inch wafer cause a lot of suction.
      Eric Bodensieck has contacted Rodel for a new pad that will incorporate
      concentric grooves and thus reduce the suction. The temporary solution
      is to block off different vacuum/pressure ports in the the carriers.
      This works but it also constrains the process folks from optimizing
      the backpressure on the wafer which is critical to polishing results.  
    
      
    
                                                                     John.  
5.102Things are looking better....SUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 11 1995 07:109
    
      The night went rather well....a pad that was in Larry's office 
    for (6) months was installed on C shift.  We broke the pad in per
    Eric's direction.  After seeing results of what happened to A2
    (see 7.last), we let the machine sit for an hour and tried running 
    more wafers at 313 lbs.  The table was fine the entire time.  
    It appears the pad is the culprit.  MQC's are going to be run.
                      
                                               D shift
5.103Robot vacuumFABSIX::R_GEESun Nov 12 1995 07:0214
	* The robot wouldn't pickup wafers out of the send stations.

	   - Cleaned the robot, end effector and air nozzle, no good.
	
	   - Checked the air pressure at the end effector, it read 9 psi.
	     turned the air pressure back to 5 psi, no good.

	   - Adjusted the vacuum switch, all set.
		
	   					B shift


	   

5.104pass downFABSIX::B_FINNMon Nov 13 1995 19:169
	System had been Idle. Upon start up it threw a wafer from the left 
        carrier.   System just came up from MQC.


					A shift

	

5.105Gear box oil levelSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 16 1995 07:2717
    
          We checked the level of the gear box and we couldn't
      even get a level detection from the top fill/full level
      port.  The oil is thicker than molasses and we don't a
      pump so we are playing doctor and using the intravenus
      method.  We are working on a gollon right now and we
      are not full yet.  We called the Strasbaugh 24-hotline
      to verify and he said the top hole (the only port hole)
      was the fill/full level port.  He also agreed that the 
      spec was correct as far as the quantity, 19 pints.  Both
      style gear boxes are the same except for the gear ratio, 
      the oil level is the same.  At different times during the
      morning, we were firing the table up to 30 RPM's and it
      already sound nice and quiet.  The oil was not spilling 
      out of the hole at this time.  We will continue to fill
      the oil until it hits its level.
                                        D shift
5.106Table cal. value changed, monthly PM completedFABSIX::B_JUBINVILLEThu Nov 16 1995 11:0016
    
    
    
      also...  the carriers on the machine were just going to
      be used for the grooved pad.  After that, they are to be
      pulled off the machine due to high wafer count.  The machine
      was left in wet idle, however, change the table RPM value
      back to 15 RPM's in WET IDLE in the calibration menu.  We
      wanted to keep the pad and carriers wet without having the
      table spin.
    
            The monthly PM was completed and the checklist will
      be posted tonight when I come in.  SUBPAC is down right now
      and I can't access the filled in checklist with all of the
      results.                               
                                           BJ
5.107Wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueSat Nov 18 1995 08:308
    
    
         Tool crashed an exercise wafer from the right
      carrier.  The carrier housing is the new style
      housing with the shims rather than the jack screws.
      Tool was cleaned up and is running product again.
    
                                              D shift
5.108CMP.A1 Status 11/18/95ASDG::POIRIERSat Nov 18 1995 19:2512
    
    System Status:
    -------------
    
    * Hot Lot completed at 6:30 P.M.
    
    * Pad being removed after 205 wafers for poor removal rate.
    
    * Carriers still in good shape with only 85 wafers run.
    
    
                                                                  John.
5.109Broke wafersFABSIX::R_GEESun Nov 19 1995 07:1311
	* Broke wafers on the pad tonight.

	  Carriers need to be built and a pad installed. One carrier is on
          the table witrht the holes completed.

		
	   					B shift


	   

5.110Wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 23 1995 06:286
    
         A wafer slid out of the left carrier and crashed into
      the right carrier.  We stripped the machine and put the
      brushes in the sink and cascading.
    
                                        D shift
5.111Air Motor ReplacedASDG::POIRIERFri Nov 24 1995 19:329
    
    Cleaning Station Air Motor:
    --------------------------
    
    The Left Cleaning Station Air Motor failed to rotate this morning.
    The Air Motor was replaced and the Brush height was readjusted.
    
                                                              John.   
             
5.112comments whilst performing the Weekly PMSUBPAC::LANDRYSat Nov 25 1995 18:3612
    The weekly PM was performed today, hence fullfilling the weekly quota.
    
    While performing the PM, we ahd to cut back the tubing on the DI spray
    gun due to a severe kink in the line.  Also, the jam nut on the inner
    conditioner lift cylinder was tightened up, as it had some slop to it.
    Slurry pump # 2 was out of whack a bit.  it was running at ~10% more
    flow than desired.  Adjusted the calibration variable to tweak in flow.
    The PM was taken one step deeper with the shuttle assembly being
    cleaned and lubed.  No other problems were found with the system.
    
    						JP  /  Doc
    
5.113Broken wafer FABSIX::R_GEEMon Nov 27 1995 07:0814
	* Broke a wafer on the pad tonight.
	 
	- First 2 wafers after a 45 min. idle period
	
	- 54 wafers on the pad.

	- 252 wafers on the carriers.

	
	   					B shift


	   

5.114passdownSTRATA::FINNMon Nov 27 1995 19:038
	No slurry flow , Ashland's pump had shut off. Contacted ashland
to restart pump.  Changed filter it was clogged.  Verified slurry flow.
System was down due to problems with the main slurry pump and other 
two systems.  System is up and running product.			

			
					Tim/Doc/Brendan
5.115Wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueThu Nov 30 1995 08:377
    
          A wafer from the left spindle slid out and crashed.
      It was discovered that the barbed fitting for the hub
      was loose.  It's not related to the crash however it
      was noted.  The machine was cleaned up and a new pad was
      installed.
                                                D shift
5.116Robot vacuum lineFABSIX::R_GEEMon Dec 11 1995 06:507
	Replaced the 1/8 " robot end effector vacuum line.
	
	   					B shift


	   

5.117CALIBRATION VARIABLES AS OF 12/21/95SUBPAC::BJUBINVILLENFD Fire - RescueThu Dec 21 1995 00:35203
  1.    BRIDGE HOME POS.= -33.751       
  2.    BRIDGE LOAD POS.= -33.709       
  3.    BRIDGE POLISH POS.= 0          
  4.    BRIDGE MIDDLE POS.= -11        
  5.    TEMP IN SCALE FACTOR= 1        
  6.    TEMP OUT SCALE= 1
  7.    TEMP IN OFFSET= -.3
  8.    TEMP OUT OFFSET= 1
  9.    BRUSH SPEED SCALE FACTOR= .719 
  10.   SLURRY 1 FLOW FACTOR= .83
  11.   SLURRY 2 FLOW FACTOR= 1.29
  12.   TABLE SHUNT LOE TEMP= 0
  13.   TABLE SHUNT HIGH TEMP= 0
  14.   BRIDGE LEFT CURTAIN POS.= -12
  15.   BRIDGE RIGHT CURTAIN POS.= -25
  16.   NOT USED
  17.   NOT USED
  18.   NOT USED
  19.   NOT USED
  20.   SLURRY 2 FLOW FACTOR= .75
  21.   NOT USED
  22.   NOT USED
  23.   NOT USED
  24.   NOT USED
  25.   NOT USED
  26.   NOT USED
  27.   NOT USED
  28.   NOT USED
  29.   NOT USED
  30.   LEFT PRESS. OUTPUT FACTOR= 6.1
  31.   LEFT PRESS. OUTPUT OFFSET= 975
  32.   LEFT PRESS. MIM.= 1100
  33.   LEFT FORCE SCALE FACTOR= .98
  34.   LEFT SPINDLE DEAD WEIGHT= 150.26
  35.   LEFT SPINDLE REDUCE RATIO= 10
  36.   LEFT SPINDLE CURRENT OFFSET= 0
  37.   LEFT PRESSURE MINIMUM, TBL2= 0
  38.   SPINDLES DOWN SOFT VALUE= 1000
  39.   SPINDLES DOWN HARD VALUE= 1000
  40.   RIGHT PRESS. OUTPUT FACTOR= 6.1
  41.   RIGHT PRESS. OUTPUT OFFSET= 975
  42.   RIGHT PRESS. MIM.= 1100
  43.   RIGHT FORCE SCALE FACTOR= .98
  44.   RIGHT SPINDLE DEAD WEIGHT= 142.37
  45.   RIGHT SPINDLE REDUCE RATIO= 10
  46.   RIGHT SPINDLE CURRENT OFFSET= .1
  47.   RIGHT PRESSURE MINIMUM= 0
  48.   SPINDLES DOWN EXTRA SOFT VALUE= 0
  49.   NOT USED
  50.   PRESS. CONTROL LOOP GAIN= .75
  51.   MAX. PRESSURE STEP DIVISOR= 30
  52.   FORCE LOOP DERIVATIVE GAIN= 56
  53.   FORCE LOOP NEG DER GAIN= 0
  54.   DOWN FORCE INTEGRAL GAIN= 0
  55.   LEFT FORCE SCALING= 1
  56.   RIGHT FORCE SCALING= 1
  57.   BACK PRESS. CONTROL GAIN= 4
  58.   NOT USED
  59.   NOT USED
  60.   6DS REPEAT TEST (0-NO, 1-YES)= 0
  61.	TABLE TEMP CONTROL GAIN= 35
  62.	TABLE REDUCER RATIO= 7.5
  63.	TABLE CURRENT OFFSET= .9
  64.	FILTER CONST.(0-UP/FAST-SLOW)= 1
  65.	BACKPRES UNITS(0-PSIA, 1-PSIG)= 0
  66.	LT BACKPRESS OFFSET= -1.5
  67.	LT BACKPRESS SCALING= 1
  68.	RT BACKPRESS OFFSET= -..255
  69.	RT BACKPRESS SCALING= 1
  70.	COND FORCE LOOP GAIN= 6
  71. 	COND FORCE DEAD WEIGHT (LBS)= 2.49
  72.	COND FORCE SCALE FACTOR= .726
  73.	COND FORCE 0-LBS, DIA (mm)-PSI= 0
  74.   NOT USED
  75.	PUMP 3 FLOW FACTOR= 0
  76.   NOT USED
  77.   NOT USED
  78.   NOT USED
  79.   NOT USED
  80.	AXIS 5 OFFSET TO SLOT 0= 3.02
  81.	AXIS 6 OFFSET TO SLOT 0= 3.04
  82.	AXIS 7 OFFSET TO SLOT 0= 3.02
  83.	AXIS 8 OFFSET TO SLOT 0= 3.02
  84.	CASSETTE SLOT SPACING= .2488
  85.	LEFT LOAD RINSE, 0-DI, 1-CHEM= 0
  86.	RIGHT LOAD RINSE, 0-DI, 1-CHEM= 0
  87.	LEFT CLEAN RINSE, 0-DI, 1-CHEM= 0
  88.	RIGHT CLEAN RINSE, 0-DI, 1-CHEM= 0
  89.	CURTAIN RINSE, 0-DI, 1-CHEM= 0
  90.	CURTAIN RINSE (1-ENBL, 0-DSBL)= 0
  91.	CURTAIN RINSING TIME (SECONDS)= 0      
  92.	FORCE 0-LBS, WAFER DIA(mm)-PSI= 0
  93.	TEMP UNITS (O-degC, 1-degF)= 0
  94.	SHOW COND FORCE (0-NO,1-YES)= 1
  95.	SHOW PUMP 3 (0-NO, 1-YES)= 0
  96.	SHOW pH (0-NO, 1-YES)= 1
  97.	TBL RNS W/CUR RNS (0-NO, 1-YES)= 0
  98.   NOT USED
  99.   NOT USED
  100. 	SPINDLE FLUSH (0-NO, 1-YES)= 1
  101.	FOUNTAIN RINSE (0-NO,1-YES)= 1
  102.	ROBOT LEADSCREW VALUE= 8000
  103.	SEND 1 ROBOT POSITION= -35.9
  104.	SEND 2 ROBOT POSITION= -24.99
  105.	REC'V 1 ROBOT POSITION= -35.5
  106.	REC'V 2 ROBOT POSITION= -25.08
  107.	LEFT LOAD ROBOT POSITION= -9.965
  108.	RIGHT LOAD ROBOT POSITION= -.51
  109.  NOT USED
  110.	ROBOT ARMS POSITION SEND 1= -3030
  111.	ROBOT ARMS POSITION SEND 2= -3090
  112.	ROBOT ARMS POSITION REC'V 1= 2880
  113.	ROBOT ARMS POSITION REC'V 2= 2900
  114.	ROBOT ARMS POSITION LEFT LOAD= 3030
  115.	ROBOT TOP ARM POSITION RT LD= 5075
  116.	ROBOT Z TOP POSITION= .31
  117.	ROBOT Z MIDDLE POSITION= .44
  118.	ROBOT Z BOTTOM POSITION= .48
  119.  NOT USED
  120.	ROBOT BOTTOM ARM POS'N RT LD= -3120
  121.  NOT USED
  122.  NOT USED
  123.  NOT USED
  124.  PROG CONDITION (0-NO,1-YES)= 1
  125.  PROG COND PARK POSITION= -2500
  126.  PROG COND CENTER POSITION= 110000
  127.  PROG COND EDGE POSITION= 143000
  128.  PROG BRUSH REDUCER RATION= 15
  129.  NOT USED
  130.	ROBOT Z OVER LOAD POSITION= .185
  131.	ROBOT Z LEFT LOAD POSITION= .39
  132.	ROBOT Z RIGHT LOAD POSITION= .37
  133.  NOT USED
  134.  SWEEP START (0-EDGE, 1-CENTER)= 1
  135.	SWEEP STYLE (0-1 WAY, 1-2 WAY)= 0
  136.	IN SITU COND (0-DSBL, 1-ENBL)= 0
  137.  NOT USED
  138.  NOT USED
  139.  NOT USED
  140.	CONTROLLED SWEEP (0-NO, 1-YES)= 1
  141.	COND SCALAR= 0
  142.	COND PERIOD WIDTH= .5
  143.	COND HORIZONTAL SHIFT= 0
  144.	COND MINIMUM COMMAND= 0
  145.	COND VERTICAL OFFSET= 0
  146.  NOT USED
  147.  SEAT WAFERS @ UNLD (0-NO, 1-YES)= 0
  148.  VAC W/FOUNT RINSE (0-NO, 1-YES)= 0
  149.  WET IDLE TABLE SPEED= 15
  150.	AUTO WET IDLE (0-NO, 1-YES)= 1
  151.	WET IDLE TABLE ON (secs)= 40
  152.	WET IDLE TABLE OFF (secs)= 60
  153.	WET IDLE CARRIERS ON (secs)= 15
  154.	WET IDLE CARRIERS OFF (secs)= 60
  155.	CARRIER PAD RINSE TIME (secs)= 5
  156.	LOAD RINSE TIME (secs)= 5
  157.	FOUNTAIN RINSE TIME (secs)= 8
  158.  NOT USED
  159.  NOT USED
  160.	2nd TABLE (0-NO, 1-YES)= 0
  161.  NOT USED
  162.	2ND TABLE REDUCER RATIO= 0
  163.  NOT USED
  164.	SCRUBBER PRESENT (0-NO, 1-YES)= 0
  165.	USE SCRUBBER (0-NO, 1-YES)= 0
  166.	SCRUBBER LOAD POSITION= 0
  167.	SCRUBBER UNLOAD POSITION= 0
  168.  NOT USED
  169.	WAFER BLOWOFF TIME (secs)= 10
  170.	APPROX WAFR BLOWOFF (psi)= 30	
  171.	MIN LIFT OFF VACUM (psia)= 8
  172.	WAFER PRESENT ERROR VAC (psia)= 7
  173.	APPROX FLUSH BLOWOUT PRES (psi)= 12.5
  174.	WAFER ABSENT ERROR VAC (psia)= 7
  175.	COOL WATER DETECT (GPM, 0-DSBL)= 0
  176.	TBL START W/SLRY (RPM, 0-DSBL)= 10 
  177.	ROTATE RESTART (0-NO, 1-YES)= 1
  178.	LOAD RINSE (0-INTRVL, 1-STEADY)= 0
  179.	ELEV TANK FILL TIME (sec)= 0
  180.	DUNK SEND (1-EN,0-DIS) RESET= 1
  181.	FILTER ROUTINE Kf= 10
  182.	FILTER ROUTINE Kl= .1
  183.	SERVICE ALARM (0-DSBL, 1-ENBL)= 0
  184.	CASS RANDOM/MATCH (0,1) RESET= 0
  185.	IDLE DELAY MINUTES (5 MINIMUM)= 5
  186.	CONSERVE LOAD RNS (0-NO, 1-YES)= 0
  187.  AIR WAND 90-OFF, 1-POL, 2-COND)= 0
  188.  NOT USED
  189.  NOT USED
  190.	SECS BAUD RATE= 9600
  191.	SECS DEVICE ID= 1
  192.	SECS T1 TIMEOUT= 5
  193.	SECS T2 TIMEOUT= 5
  194.	SECS T3 TIMEOUT= 15
  195.	SECS T4 TIMEOUT= 15
  196.	SECS RETRY LIMIT= 3
  197.	SECS COMM PORT (1 OR 2)= 1
  198.	REPORT INTERVAL (SECONDS)= 10
  199.	REPORT MODE (1-ENBL, 0-DSBL)= 1
  200.	PASSWORD (4 DIGIT NUMBER)= 

                                         
5.118Start-up for 12/29SUBPAC::BJUBINVILLENFD Fire - RescueFri Dec 29 1995 15:0912
    
    
        The DI came on but is still out of spec so we didn't want
      to keep it to the tool so we shut off the DI feed and opened the
      by-pass.  Facilities still have some by-passes opened, that's
      why we don't have very high pressure.
    
        The slurry lines are all back to normal and slurry is flowing.
      We put the filter on but but we didn't run the slurry through
      the machine because of the DI situation.
     
                                                       Mitch,John,Leon,BJ
5.119Shuttle Lubed & Adjusted / Right Cleaning Station Air Motor NoisyASDG::POIRIERSat Dec 30 1995 15:2111
    
    * Shuttle:
    
      The Shuttle Rails were cleaned and lubed today. The Air Flow Needle
      Valve Controllers were adjusted to increase the travel rate of the 
      Shuttle. This will shave-off a few seconds from every Auto Cycle.
      
      We also noticed that the Right Cleaning Station Air Motor was noisy.
      A shortage request was generated in SGMR for 4 more Air Motors.
    
                                                           John & Mitch.
5.120CMP.A1 START-UP COMPLETESUBPAC::BJUBINVILLENFD Fire - RescueSun Dec 31 1995 14:3143

                    CMP.A1 STARTUP CHECKSHEET
                    -------------------------

                                                       Shift       Date
                                                       -------     -------
  1.  Ensure that all necessary facilities are         B, C + D     12/31/95  
      available and ready to use.                      

  2.  Power up each system.                            B, C + D     12/31/95  
                                                       
  3.  Power up each vacuum pump cabinet.               B, C + D     12/31/95  

  4.  Power up each Neslab TCU.                        B, C + D     12/31/95  

  5.  Open incoming DI Water Valve to each             B, C + D     12/31/95
      system.                                       

  6.  Close UDR valve at each system.                  B, C + D     12/31/95

  7.  Install wafer carriers with new films.           B, C + D     12/31/95

  8.  Place cassettes in left and right send
      and receive stations to fill DI Water
      tanks.                                       

  9.  Re-connect slurry inlet and outlet lines         B, C + D     12/31/95
      to filter housings and install new filters.    

 10.  Change RTC status to SCHEDULED DOWN and
      perform a Weekly PM on each system.           
  
 11.  Install new polishing pads.               

 12.  Re-install PVA scrub brushes.                    B, C + D     12/31/95

 13.  Place systems in wet idle mode.                  B, C + D     12/31/95

 14.  Change RTC status to VERIFY and notify
      Operations that the systems are ready
      for MQC.                                        
                                                        
5.121vacuum problems / rotory unionFABSIX::R_GEETue Jan 02 1996 07:4431
	


	* Calibrated the slurry flow for both pumps.

	* The system errored for bad vacuum on the right spindle. We
	  found the right rotory union was leaking. 

	  - Removed the old rotory union.
	
	  - Assembled the new rotory union.

	Things left to do:

	  - get and install a new o-ring for the top of the rotory union
	    rod.

	  - clean spindle shaft.

	  - install the rotory union assembly, it's already assembled.

	  - install and align the hub.
  
	  - install carriers and test, sorry no carriers built.


	   					B shift


	   

5.122rotary union installed (right spindle)SUBPAC::LANDRYTue Jan 02 1996 19:3610
    The rotary union was installed this morning on the right spindle. 
    Vacuum performance was checked and it was fine.  The hub was
    re-installed and alignment was attempted.  Didn't go well at all.
    Backed up and checked the spindle alignment.  Found it to be out of
    tolerance (was .004, not to far out).  Adjusted this to .001.  Re-did
    the hub alignment and brought that in to spec as well.  The system is
    currently in VERIFY, awaiting a new pad / carrier installation and then
    to run breakin wafers.
    
    					Doc  /  Jim T  /  Kyle
5.123DOWNFORCE PROBLEMS, LEAKING HOUSINGSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 05 1996 01:1755
From:	FABSIX::R_GEE "Bob Gee, Beeper #184  03-Jan-1996 0808"  3-JAN-1996 08:12:00.39
To:	@CMP_ALL
CC:	R_GEE
Subj:	Passdown, the notes file is down. Info for CMP.a1,CMP.B1,CMP.B2

	CMP.A1
	------

	High downforce was noticed on the display screen, 500 lbs
	from the right spindle, the system was emo'ed. 

		- Found the right filter housing was leaking. 
		  Water had leaked onto the load cell amplifier,
                  cleaned the water up and dried the amplifier.

	
		Production is running mqc's. 
		
					B shift


	   

	* The system was down until 3am due to scratches (?)
	  found on the mqc wafers. A second mqc was run with no 
	  scratches found.

	* At 6 am the system powered itself down in the middle 
	  of polishing for no apparent reason, crashing all 4 wafers.

		- we found that carrier #1 had been polishing
		  a broken wafer for some time. The wafer detection
		  didn't detect the crash due to the small size of 
		  the particles. 

		We thing the system powered itself down due to high
		amps. from  either the table or head #1. 
	
	      					
					B shift
	High downforce was noticed on the display screen, 500 lbs
	from the right spindle, the system was emo'ed. 

		- Found the right filter housing was leaking. 
		  Water had leaked onto the load cell amplifier,
                  cleaned the water up and dried the amplifier.

	
		Production is running mqc's. 
		
					B shift


	   

5.124Robot "at Left Load" Position Changed.ASDG::POIRIERFri Jan 05 1996 19:385
    
    * Robot Load Position Changed:
     ----------------------------
     The "Robot at Left Load" Position was changed today.
     Calibration variable #107 was changed from -9.965 to -9.90. 
5.125Robot Pnuematic Sensor.ASDG::POIRIERSat Jan 06 1996 19:3315
    
    * Robot Pnumatic Sensor:
      ---------------------
      The Robot Sensor acted-up again Today. The system kept booting out 
      of auto cycles after completing 1 cycle. No alarms or error messages
      appeared during this situation. When I pressed the Cycle Start 
      button to resume, an error message appeared along the bottom of
      the screen reading "Vacuum present at robot end effector". I took
      the cover off the robot and found the green LED flickering. A 1/4
      turn of the fine adjustment on the pnuematic sensor fixed the problem.
    
      Keep an eye on the sensor over the next few days but don't knock
      yourselves out to adjust it as were replacing the robot on Monday.
    
    
5.126Rbot wafer detection vacuum sensorFABSIX::L_SHKOLNIKOVFri Jan 12 1996 06:396
                                                                                            
          Machine would not continue cycling after each run and would not
          go into wet mode. Robot vacuum sensor has been adjusted and
          problem disappeared.                             
    
          L.S.
5.127ROBOT SENSOR ERRORSFABSIX::B_WESTPHALFri Jan 12 1996 19:384
CMP.A1 - Down for problems w/ the robot wafer sensor.  The same was adjusted
	 and the tool was given back to production.
   
5.128ROBOT SENSORFABSIX::B_WESTPHALSat Jan 13 1996 19:426
	 
CMP.A1 - Down For robot sensing errors.  Attempted to adjust the sensor by 
	 turning the needle valve until the lED flickered but it never did.
	 Made several attemps  at adjusting both tthe needle valve and pot.
	 but could not get it in.  Suspect that the switch itself is bad.
	 
5.129cal of wafer detect sensorFABSIX::B_FINNSun Jan 14 1996 17:569

	B shift replaced Robot Wafer Detect Sensor.  We completed the repair by 
        calibrating said sensor.  Ran test wafers aok. Returned the system
        manufacturing.


					A shift

5.130vacuum FABSIX::B_WESTPHALThu Jan 18 1996 19:515
    CMP.A1 - Down for dropping the carriers of the spindle.  This happened
    at the end of the shift so I didn't have much time to look at it but 
    believe the pump to be the problem.  Noticed that A1's pump was running
    alittle louder than A2's.  The outside cabinet housing the pump was
    very warm to the touch and hot air could be felt coming from fan.
5.131Spindles dropping off hubsSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 19 1996 06:4925
    
    
        The spindles were dropping like bird turd last night.
      With the assistance of Eric B., we lowered the "new"
      software calibration for spindle flush pressure from 20 
      to 15 psi.  The extra 5 psi was pushing the carriers off
      of the hubs.  The o-rings however are another story.  
      They were the stock line line type o-ring and not the
      white proper o-rings.  Eric replaced the four o-rings
      in the hubs with new ones from stock.  
    
          When the right carrier fell, it had perfect timing.
      It fell on the robot just after the robot arm was picking
      up the wafer out of the right load cup.  The pedistal took
      a hit but seems to be ok.  We now have an error.  After the 
      wafer in the right cup is blown off and is ready for pick-up,
      The system alarms saying there is no vacuum in the cup.  The 
      alarm clears after 4 or so "rings" and continues on its way.
      The pedistals were swapped and that wasn't it.  The vacuum
      setting might of gotten knocked off  or the pedistal isn't
      coming up enough to form a good seal.  Along with the other
      problems going on, we lets them run it since it wasn't causing
      a problem.  I logged a warning on it and informed D shift OPS.
      Todd from Strasbaugh is suppose to be on site today.  I will
      communicate this problem to C shift and ENG.
5.132Clogged drain linesSUBPAC::BJUBINVILLENFD Fire - RescueFri Jan 19 1996 07:0915
    
         Another thing....  while we were working on the
      load cups, we noticed the shuttle area had 3" of
      DI water and slurry.  It wasn't draining.  We traced
      it back to the union for the GCW waste where it comes
      out of the machine and before it goes through the floor.
      It was clogged beyond that.  An action-call was placed 
      to security and site services brought in the roto-roter
      and cleaned out the 2" line all the way to the main 4"
      line.  In a distance of 15ft, why are there soooooooooooo
      many elbows.  That is sure to cause a problem in the
      future.
                    All's a flowin' now.......
                         
                                                  D shift 
5.133vacuum regulator leakFABSIX::B_FINNFri Jan 19 1996 19:0810
	Manufacturing ran a couple of lots, before system failed for carrier 
vacuum problems. Rep traced the problem to vacuum regulator valve, when 
back pressure would come on it depleted the vacuum leaking by the valve seat. 
Replaced valve.    System back up to manuf.



					A shift

5.134Robot Z heightSUBPAC::BJUBINVILLENFD Fire - RescueSat Jan 20 1996 05:2811
    
    
         The robot Z-height must of gotten a bit out position
      when it was nailed.  The robot wasn't coming down enough
      to pickupt the wafers.  It was missing them from time to 
      time on both send elevators and was dropping them into 
      place in the receive elevators.  The Z-bottom was adjusted
      from 1 to 1.05.  Its been running great all morning without
      any problems but...  keep an eye on it.
    
                                              D shift
5.135The new airmotors have arrivedSUBPAC::TOOKERMon Jan 22 1996 14:048
CMP Equip. Techs,

I have just received 4 new air motors for the cmp.a1/.a2 clean stations.  These
need to be installed as soon as possible.  I will leave them on my desk in a
plastic bag.

Thanks,
Jim Tooker
5.136New robot calibration proceduresSUBPAC::TOOKERMon Jan 22 1996 15:059
CMP Equip. Techs,

I have placed clean room copies of the procedures for calibrating the new
robot on cmp.a1 in the very back of Strasbaugh manual #2.  Please refer to
these new sections when performing any robot position calibration at an
elevator or load station.  

Thanks,
Jim
5.137Left SEND elevatorSUBPAC::BJUBINVILLENFD Fire - RescueWed Jan 24 1996 06:4611
    

          The left SEND elevator was removed and everything was
       cleaned up and lubed.  There was an excessive amount of 
       grease on the rails and lead screw.  The motor assembly
       showed signs of rust and was cleaned up as well.  OPS
       are running the tool right now so it was not tested as
       far a repeatability is concerned but it homes fine.  The 
       WARNING will stay on the tool until it is verified. 
              
                                           BJ
5.138pm in progressFABSIX::B_FINNWed Jan 24 1996 19:3814
	Kyle saw a 400 A' difference in removal rate between head 1 + 2, 
he asked us to check out the system ie downforce.  This was great since we 
wanted the the system for a monthly. Trained brad on hub alignments.  Right 
spindle was 9mils out.  Have been adjusting it since, real pain, came as close 
as 2 mils but lost it. Sorry Brian and Leon.   Check sheets for weekly and 
monthly have been forward to you. Jim would like the air motors to go in also.
Could not find regulator setup to check ep pressure regulators. I got a fluke 
adapter from Jim Tooker and brad added a tee, but we did not get past the 
spindle angle.   Screws for covers are on black pad cabinet.


					A shift

5.139AIR MOTOR REPLACEDSUBPAC::BJUBINVILLENFD Fire - RescueThu Jan 25 1996 05:527
   
         The air motors were replaced with new ones during
     the monthly PM.

                                         BJ

5.140it's hose was pinched!SUBPAC::LANDRYThu Jan 25 1996 12:288
    Was called to look at why slurry pump two was not flowing when it was 
    commanded to.  Found that the hose length was a tad bit long, causing 
    the head rotation to pinch the tubing off on the output side.  Cut back 
    about an inch on the hose, then reconnected.  Performed slurry pump 
    calibration on pump two ONLY.
    
    					Dr.
    
5.141Robot problem back.FABSIX::L_SHKOLNIKOVFri Jan 26 1996 07:134
        
       Robot didn't pick up the wafer from right send elevator.
      Elevator was adjusted but problem appeared again after 6-8 wafers
      run. Run machine with left send elevator.  
5.142system checks in progressSUBPAC::LANDRYFri Jan 26 1996 18:2343
      The system has been experiencing abnormal polishing performance
      for the past couple of shifts.  It appears to have just begun
      these problems PRIOR to the PM that was performed mid-week.
      Uniformity, removal rates, etc are bad.  Therefore, Jim and I 
      began to check on a few components in hope of locating an obvious 
      cause.  Details thus far are below.........

      -  Checked the E-P regulators (Fairchilds) for both the left and
         right spindles.  The high end (80 PSI) was off a tad bit on
         both spindles, but not enough to cause any problems.  We used
         the new pressure transducer/fluke meter to read the pressure.
         A neat little tool!  The pressures can be read more accurately
         than the gauges we were using previously.

      -  Performed Spindle Down Force calibration.  Both were within spec.
         We reran the left spindle a second time to tighten the margin.  
         The right was run once.  Again, fairly certain this is not 
         causing the disparities.  

      -  Next, the hub alignment was confirmed in spec for both spindles.
         So, this is ruled out as the cause.

      THEREFORE.........

      -  The spindle alignment needs to be confirmed next.  

      -  Maybe look at the polish table alignment, which is never out???

      -  You might want to call Jim T at home (before the night gets too 
         late, of course)

      AFTER THAT........

      -  Punt??  No....maybe run a wafer or two to see if any little 
         tweaks helped.  

      FYI department.....
    
      -  Both carriers and both scrub brushes are soaking by the sink.
      -  A new pad will need to be laid down when ready to exercise wafers.
    
      					Jim T  /  Doc
                                                                           
5.143left spindle pressure transducer / RobotFABSIX::R_GEETue Jan 30 1996 07:1232
      * We have been having carriers sticking to the pad intermittently
	do to poor vacuum on the left carrier. The system won't allow
	the spindles to raise without proper vacuum. 
	
	   Found the problem to be a broken wire at the connector for
	   the left spindle pressure transducer. Repaired the broken
	   wire and ran for several hours with good vacuum.
----------------						
      * The Robot wouldn't pickup wafers from either send cassette.
	The vacuum seems to be working fine. It looked like the robot
	bottom and middle positions are to low. It picks up out of
	the load cups fine.

	   Cleaned the robot and end effector, no good.
	
	   Checked the robot to send cassette level it looked o.k.

	   Adjusted the robot middle position (CV#117) from .96 to .85

	   Adjusted the robot bottom position (CV#118) from 1.05 to .98
	
	   Adjusted the left send offset to slot "0" from 3.02 to 3.1

	The height alignments look good but still not working. Our 
	unfamiliarity with the software and robot made the job very 
	time consuming and frustrating.	
	
					B shift


	   

5.144robot vacuum switch calibrationSUBPAC::TOOKERTue Jan 30 1996 19:1325
    This morning I calibrated the robot vacuum switch as per the new robot
    calibration procedures added to the back of service manual #2.  I found
    this procedure very easy to perform in comparison to the calibration
    procedure for the old robot vacuum sensor.  Anyone attempting this new
    procedure must use 2 mirrors in order to correctly read the digital
    readout on the vacuum switch.  I did this by wedging the mirror from
    the toolbox in front of the vacuum switch and then holding a wafer in
    my left hand to reverse the readout image.  The switches for setting
    the vacuum switch can be reached and pressed with the short end of an
    allen wrench.  After completing the calibration, I cycled 50 wafers
    through the system (25 per station).  The robot picked up all wafers 
    without a problem.  However, there is still a problem with the left 
    receive cassette station.  When the robot loads a wafer into the cassette,
    the robot arm is dragging the wafer partially out of the cassette.  Some
    wafers are coming back out further than others.  Once or twice, I saw a
    wafer get caught on the top of the water tank and lift the cassette up.
    This could break wafers and if the sensor arm raises up, it will some-
    times abort the auto cycle.  I believe the problem can be fixed by re-
    calibrating the elevator and robot positions, but there was not enough
    time today to do this.  In the interim, I have placed a warning on cmp.a1 
    to NOT use the left send and receive cassettes.  Tomorrow morning I will 
    try the new calibration procedure for this and see how it goes. 
    
           Jim 

5.145Robot to elevator alignment / wafer presentFABSIX::R_GEEWed Jan 31 1996 06:0037
	Aligned the robot arm to the left send and left 
	receive elevators. Ran over 100 wafers using only
	the left send and left receive elevators, everything
	looked very smooth. Removed the warning in RTC.

		Calibation variables changed:
						 old 	new
		#80 left send offset to slot 0   3.1    3.07

	        #82 left recv offset to slot 0   3.15   3.12

	       #116 robot up/down top position    .83    .8


	* Still having problems with the robot thinking there
	  is a wafer on the end effector before it moves to the 
	  send elevators, you must push cycle start to continue.
	  
	      We believe the problem is that before the robot
	      moves to send it blows the water out of the end 
	      effector line. This creates a puddle of water
	      between the end effector vacuum hole and the top 
	      cover. When the robot checks for a wafer present,
	      before moving the send cassette it sucks up water 
	      and thing there is a wafer present. This was confirmed
	      by blowing the water out from under the end effector
	      manually before it looks for a wafer, the error was 
	      gone. This was tested many times.

	The error is still there and needs to be addressed. A
	warning has been logged in RTC.

					B shift


	   

5.146"wafer present on robot" error problemSUBPAC::TOOKERWed Jan 31 1996 16:1023
The last few days we have been seeing a problem on cmp.a1 where the "AUTO
CYCLE" will abort due to a "WAFER FOUND ON ROBOT" alarm.  This has been
occuring at times prior to the robot getting wafers from the send cassettes.
Prior to getting wafers from a send cassette the system checks for any wafers
already present on the robot.  We are getting this alarm because the air
pressure to the robot vacuum switch is too low.  In talking to Strasbaugh,
they feel that there is a kink somnewhere in the airline running from 
solenoid 44 to the robot air pressure regulator.  This is the same regulator
previously used for the old robot vacuum sensor pressure.  The pressure at
the regulator should be 15 psi and the pressure on the robot vacuum switch
readout should be 5.0 psi without a wafer present.  Currently the robot
vacuum switch readout is ~1.4 psi.  This is not enough pressure to blow the
water out from the vacuum line after unloading a wafer.  When the system
checks to see if a wafer is present prior to getting another wafer from the
send cassette, the system thinks there is a wafer present due to the water
in the vacuum line.  If this alarm occurs and the "AUTO CYCLE" aborts, you
will need to press "START" to continue the "AUTO CYCLE".  When time allows,
maintenance will need to trace the lines to check for any kinks.  Please call
me at home at 508-874-0839 if there are any questions.

Thanks,
Jim

5.147Vacuum problem correctedSUBPAC::BJUBINVILLENFD Fire - RescueThu Feb 01 1996 05:458
    

         Found a slit in the 1/8" end effector line.  
     Most definitely a potential cause for the vacuum
     problems.  Replaced line and checked with transfer 
     program.  Will remove WARNING when everything else 
     looks ok. 
                             BJ
5.148Robot End Effector Pressure too lowASDG::POIRIERThu Feb 01 1996 19:2719
    
    * Robot Air Pressure:
      ------------------
      With the Robot End Effector Pressure Regulator set at 15 psi, the
      pressure at the Robot should be 5 psi (Dynamic). Unfortunately, its only
      1.2 psi. This condition is causing the system to error-out of
      the Auto Polish Cycle after every run. (Software checks for wafer on
      End Effector before picking-up wafer at the Send Cassette).
    
      We visually inspected the air hoses that run from SOl 44 to the Robot
      and found no kinks. We removed the barbed fittings from Sol 44 and
      discovered them clogged with slurry. They were cleaned-out and re-
      assembled. The pressure at the Robot increaced only slightly.
      
      We suspect that the solenoid valve itself may be clogged with slurry
      causing the pressure drop. "D" shift, we have ordered a replacement 
      valve fron the stockroom. Please replace sol valve #44 and check
      the air press at the Robot.
                                                  "C" Shift
5.149Robot amendmentSUBPAC::BJUBINVILLENFD Fire - RescueThu Feb 01 1996 19:289
    
          Due to the fire alarm this morning, my reply 
       was rushed.  It was suppose to read that I had replaced 
       the 1/8" line and was about to start the transfer
       program before the alarm went off.  I spoke with EE 
       in the morning and they understood what I did and were 
       going to go ahead and recalibrate the robot..
                        
                                   BJ
5.150Robot Vacuum Sensor CalibratedASDG::POIRIERFri Feb 02 1996 19:156
    
    * Robot Vacuum Sensor Cal:
      -----------------------
    
      The Robot Vacuum Sensor was calibrated to reflect the change in air
      pressure at the End Effector. 
5.151load station alignmentFABSIX::B_FINNMon Feb 05 1996 18:4915
	Broke a wafer form left spindle on the pad.  Checked alignment and found
	the wafer was intermittently not being placed in the carrier at 
        load correctly. Wafer was intermittently bridging the black exposure 
        ring.

	Cleaned up broken wafers, (left broken, Right wafer scratched and 
        cracked), Replaced one Black extension ring, Changed Pad.

	Realigned left Load Station to the lefty carrier, tested aok.


						A shift


5.152crash & valveSUBPAC::LANDRYTue Feb 06 1996 19:1713
      A wafer spit out on the table from the left carrier early
      this morning.  Jim T's trigger finger on the EMO save one
      of the carriers.  The left load station alignment was verified.
      Mess cleaned up, new pad, one new carrier installed, a new
      diamond disk installed and we were ready to run again.  Breakin' 
      wafers were successfully run.  MQC's passed.  Turned the system 
      back into manufacturings hands.

      While the mess was being cleaned, the teflon valve on the "dog
      dish" was replaced with a new one from stock (listed in the
      Strasbaugh parts inventory).  

      						A Shift
5.153pm / pcSUBPAC::LANDRYTue Feb 13 1996 19:288
    The monthly PM was performed and completed today.  I'm short on time,
    so the chechsheet will be completed and entered tomorrow.
    
    On a good note......Jim T got the PC back in shape.  The down force
    program works just fine!
    
    						A Shift
    
5.154Rotbot Vac Switch in need of adjustmentASDG::POIRIERThu Feb 15 1996 19:307
    
    * Robot Vacuum Switch:
      -------------------
      The system started booting-out of Auto Cycle again this afternoon.
      Its doing this on every 2nd or 3rd cycle. The Robot Vacuum switch
      needs another adjustment after completion of the current run. The
      run should be completed within the next few hours.
5.155Vacuum switch recalibratedSUBPAC::BJUBINVILLENFD Fire - RescueFri Feb 16 1996 07:286
     
        The vacuum sensor in the robot was tweaked again.  The
      tool kept stopping after polish cycles without any warnings.

                                                  D shift
5.156Software bugSUBPAC::TOOKERFri Feb 16 1996 14:3515
On cmp.a1 and cmp.a2 there is a bug in the software that will occasionally
cause the autocycle to abort and go to idle.  This occurs after the robot
has unloaded wafers and before trying to load more wafers.  At this point,
the system checks to see if there are any wafers present on the robot before
trying to take any wafers from a load cassette.  Strasbaugh says that the
time delay in the software is too short and this is causing the system to
eroneously sense the presence of a wafer.  Next week, Strasbaugh is sending
a new software Rev. to us that will correct this problem.  If the auto cycle
does abort to idle, hit "start" to resume the autocycle.  Maintenance techs.,
please do not adjust the vacuum switch as a fix for this problem.  Only adjust
the vacuum switch when the robot misses picking up a wafer from a cassette.
(i.e. the robot moves into the cassette and comes back out without a wafer)

Thanks,
Jim T.
5.157Cut down a speedfam padFABSIX::R_GEEMon Feb 26 1996 06:4320
	We completed the weekly p.m. and replaced the pad with a
	cut down speedfam pad. This was done because the engineering
	and strasbaugh pads seem to be failing for poor uniformity
	after only a very few wafers. The last Strasbaugh pad, (new
	shipment) replaced had only 225 wafers and we had problems 
	qualifing the one on cmp.a2. The Engineering pad had to be 
	replaced after only 98 wafers. The speedfam pad is from an old
	lot of pads and hopefully will last longer.  
    
                            TOP     CR QJMR0313A,P/II 
                                    S-IV, 32", II        
                                    LOT# 6AP-JP2
                                    
			bottom      cr suba iv,28",ii
			            lot# 6av-da1-4            

	Ran 12 breakin wafers and 2 mqc wafers, it passed mqc's.		


					B shift				 
5.158air motor replacedFABSIX::B_FINNWed Feb 28 1996 19:328
	

		Robot lost position, would not continue needed to be 
                re home. Air brush motor on left clean station failed.
		Replaced air motor, tested motor aok


						A shift
5.159Robot Pick-up Problems on both Send StationsASDG::POIRIERThu Feb 29 1996 19:3827
      
    * Robot pick-up problems at both Send Stations:
      --------------------------------------------
      The Robot was having wafer pick-up problems at the Right Send Station
      this afternoon. The Robot End Effector was contacting the edges
      of the wafers. The Elevator was re-homed and NO improvement was
      observed. We noticed that the Elevator was vibrating at its upper-
      most position. We performed a system reboot and the Elevater started
      functioning normally again. I heard that this condition appeared 
      recently on another shift. Anyone have any thoughts? We're suspecting
      the DCX Module.
    
      
    * Robot not arriving at "Left Send" error:
      ----------------------------------------
      This error came-up repeatedly all afternoon. We noticed that the 
      Robot X axis position was 35.990 while the error was appearing. We
      checked the calibration variable for the Left Send and found the 
      value to be 36.150. The Robot was being mechanically restrained
      from reaching its set position. Thus, the error "Robot Not Arriving
      At Left Send."
    
      WE cut six folds from the Robot lead screw bellows to eliminate
      the bellows from wedging between the side of the tool & the Robot.
      A cassette of wafers were run on both Send Stations without any
      errors.
      
5.160New wafer push bars installedSUBPAC::TOOKERFri Mar 01 1996 15:399
Today we installed new style wafer push bars in cmp.a1 and cmp.a2.  These
are the bars located inside each elevator rinse tank to help keep the wafers
from floating out of the cassettes.  The new bars are thicker than the old
ones and should do a much better job of keeping the wafers in the cassettes.
Please ensure that the wafers are completely in the cassette when placing a
cassette of wafers onto a cassette station.

Jim T.

5.161uniformity/scratchesFABSIX::B_FINNMon Mar 04 1996 19:3012
	Problem: Bad uniformity and Scratches.

	Engineering made several changes. Lowered Platen temp to 
        25 degrees,  removed suba IV.  Failed uniformity at 15%
	Trying now with ic 1000 and suba IV, but temp still low.
	If it fails again Temp will need to be set back to 37 degree's


					A shift


5.162wafer breakageSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 07 1996 23:177
 
       System crashed two wafers.  Upon cleaning up that carriers,
   Leon found a bad chack valve inside one of the carreirs.  The
   tool was cleaned up and new pad and carriers installed.
                
                                          D shift
5.163Wafer breakageFABSIX::B_JUBINVILLEFri Mar 08 1996 07:0012
    

           During the running of the break-in wafers
       a ~crash~ consisting of two wafers and their    
       respective carriers took place.  The small pieces 
       of film that was as a result of not being blown 
       out during assembly was evident on the brushes, 
       the wafers, pad and check valves.  After everything 
       was cleaned up, new carriers and pad were put on the tool.

                                      BJ
                                            
5.164scratches and brushesSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 09 1996 04:5220

       A1 was down for scratches at the end of C shift.  The
   brushes were off the tool soaking.  After inspection, a
   few small pieces of wafer debris were embedded in the brushes.
   The brushes were ordered but to my lovely luck this week, the
   brushes are in VSR1 in stockroom 6 and it is down until Monday.
   I called around and there were now brushes in hiding.  Then to
   my luck again, A2 crashes so I took the brushes off of A2 and
   put them on A1 since there was already a new pad on A1.  the tool
   was MQC'ed and passed on everything, great!  But...  the right
   spindle was sticking at lift off.  The conditioner was looked at
   and it was ok.  While running product, a wafer slid out of the
   right carrier but did not crash.  My partner removed the right 
   spindle and performed open heart surgery.  After the carrier
   was rebuilt and put back on the tool.  A quick MQC was done and
   it failed for MQC's and scratches.  Off goes the pad...  The system
   came back up after that.....

                                                 BJ
5.165Conditioner Head / Right Send Elevator - Software BugASDG::POIRIERSat Mar 09 1996 19:2916
    
    * CONDITIONER MOTION:
      ------------------
      The Conditioner was hesitating and making a growling noise during its 
      sweep.  One of the socket head cap screws had loosened-up and was
      preventing the disk holder from gimbling. The screw was tightened and
      the problem disappeared.
    
    
    * Right Send Elevator/ Software Bug:
      ---------------------------------
      The Right Send Elevator Continues to Loose its position. This causes
      the Robot End Effector to miss a wafer at pick-up & hit the side of the 
      next wafer. We had to EMO the system twice today to recover from this
      situation. Jim Tooker is waiting for Strasbaugh to send us a fix for
      this bug.
5.166robot end effectorFABSIX::R_GEEMon Mar 11 1996 03:306
	* The robot would pickup the wafer out of the send cassette 
	  and then error for no wafer on the end effector. Cleaned the
	  end effector but it didn't help. Since CMP.A2 was down waiting
	  for parts we swapped the end effectors, the problem went away.
	  	
					B shift				 
5.167vac line tornSTRATA::FINNMon Mar 11 1996 19:198

	End effect not picking up wafers. Replaced 1/8 inch vacuum hose, which
        was torn. Vacuum was poor at wafer. Cleaned bur from end effector. 
        Test AOK

						A shift
        
5.168left receiveSTRATA::RIDLONEliminate the obviousTue Mar 12 1996 20:345
    Left receive elevator not working. Found the cassette sensor wire
    crimped under the elevator preventing it from making the home sensor.
    Once it was removed it homed fine.
    
    						B shift
5.169software changed to ver 4.16.3FABSIX::B_FINNWed Mar 13 1996 19:3825
    
    
                     Version 4.16.3 installed today.
	System was placed down for system robot halting during wafer 
load.  We have been working the last three weeks to get a new version 
of software in to fix this bug.  We picked it up at the dock this 
morning and installed it.  Config values and error limits in dos were 
changed as well as several on the config/parameter page.  We also had 
to make mechanical adjustment to upper and lower limit switches on 
both right and left spindles.  this is due to the new way the software 
reads the switches.

System is running breakin wafers Right now, to test the complete 
software.  when this is complete the "handling" program must be run 
until d Shift is satisfied there are no hang ups.  

Also had to stone the end effector to  remove burs.  Its running well 
now.


				Brendan/Hawk



5.170Auto-cycle failure, slurry filter issueSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 14 1996 04:5217
    
         After cycling ~250 wafers on the tool, the robot had
     failed to continue auto-cycle (4) times.  Since this is
     why we were verifying the tool, I stopped the run.  B2
     had its slurry filter clogged and there were none in stock
     so I used the one from A1.  I tried calling PE to get the 
     ok to run the tool without the filter but I couldn't reach
     him.  OPS are running MQC's to at least get that part done.
     If PE gives it the thumbs up, at least it will be ready
     to be turned over to production.
 
         The slurry upgrade for the peristaltic pumps has been
     completed.  Everything is flowing well and the calibration
     of the flow was completed.

                                                 BJ
5.171Robot bottom position @ L Send/ Robot Load Pos ChangedASDG::POIRIERThu Mar 14 1996 19:4328
                                                                   
    * Load Positions Changed:
      ----------------------
      The following calibration variables were changed due to poor loading
      at both load stations:
    
                              #107 from -9.6 to -10
    
                              #108 from -.2 to -.3
    
                              #115 from 50100 to 50600
    
      These changes work well when wafers are picked-up from the right send 
      elevator. However, these settings do not work well with the left send
      station because the elevator needs calibration. The system was put
      back UP with a warning on the left send elevator.
    
    
     * Robot Bottom Position @ Right Send:
       ----------------------------------
       The Robot (Bottom) position at the Right Send Station was changed
       due to excessive deflection of the wafers. This may have been the
       cause of the intemittent wafer pick-up problems at the Right Send
       Station. The excess deflection of the wafer may have been causing
       the Robot End Effector vacuum seal to break. If this happens, the
       Robot will back out of the cassette, the elevator will index up by
       one slot and the End Effector will crash into the side of the next 
       wafer. Calibration variable #118 was changed from .98 to .97
5.172Left send elevatorSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 14 1996 22:2314

        Previous shift had put a WARNING on the left send
    elevator due to positioning and handling issues.  Two
    full boats of wafers were run using the HANDLING recipe 
    and we experienced no problems at all.  With the elevator
    that is...

        One thing I did notice is a "Both up and down sensors
    on the left spindle were on" error pop up for a second.  These
    switches were adjusted during the software upgrade.  These
    might have to be adjusted if this becomes an issue.

                                          D shift
5.173handling troubleFABSIX::B_WESTPHALFri Mar 15 1996 19:409
CMP.A1 - Tool is down for robot pickup/handling problems. The robot vacuum
         switch was recalibrated and the handling problems disappeared for
         a few hours only to reappear later in the shift. 

         Strasbaugh was contacted and they suspect we have a vacuum leak 
         somewhere between sol #44 and the robot end effector. Trace hose
         # 94 for leaks.

         NOTE: A new slurry filter was installed on the tool today. 
5.174Robot problem went away???FABSIX::L_SHKOLNIKOVSat Mar 16 1996 07:3411
              The robot wouldn't pickup the wafer out of the send cassette.
          Vacuum line between sol#44 and the robot was bypassed.
          No vacuum leak was found. Vacuum was checked at the end effector
          point looks good 11 mm/m - on  other side of vacuum line 12.
          Recalibrated the vacuum sensor and run 70 wafers - OK.
          After that robot failed again - didn't pick up wafer from slot 14.
          Checked sensor calibration - fine. Looks like  elevator getting
          out of alignment after robot arm reaches  certain slot.
          Otherwise vacuum sensor doesn't work properly.

                                                          D - shift
5.175Robot Vac Problems / Load Station Drop-Off PositionsASDG::POIRIERDuuuhhhSat Mar 16 1996 19:3725
    
    * Robot Vacuum Problems:
      ---------------------
      Vacuum line #94 that runs from SOl#44 to the Robot End Effector
      was monitored for leaks, per Strasbaugh's instructions,
      and none were found.
    
      The Robot Vacuum Switch has been repeatedly recal'ed over the last few
      days. The robot will handle wafers properly for 1 to 18 hrs before
      failing. (We are observing a direct "cause & effect" after the vacuum
      switch is adjusted). "B" Shift, you may want to continue the
      troubleshooting by checking all the vac connections in the Robot.   
    
      A & B SHift, We've been working this problem with Todd Fordyce 
      at Strasbaugh.
    
    
      Load Station Drop Off Positions:
      -------------------------------
      After the Robot Handling issue is resolved, the robot drop-off
      positions at the 2 load stations need fine tuning.
      This is due to a calibration of the "Robot at Right Send Station"
      adjustment that was performed yesterday. We were not able to 
      follow thru with the Drop-off at Load adjustment due to the 
      Robot handling errors.
5.176robot stuffSTRATA::RIDLONEliminate the obviousSun Mar 17 1996 06:5023
    Checked end effector and found burrs in the end of it from 
    hitting wafers, these were cleaned off with a razor blade (any burrs
    will prevent the wafer from making a good seal which is essential in
    any vacuum system). Checked the
    robot arm for center (paddle end over rectangle black cover) and it was
    out, also from smacking into wafers, this was bent back to center
    (pretty flimsy set up). 
    
    On the right send cassette noticed that the spacing between the wafer
    and end effector wasn't consistent from top to bottom. Changed the
    motor and encoder to correct this. 
    
    The left and right receive cassette offsets were adjusted for smoother
    wafer placement. The Left and right load placement numbers were
    adjusted also for smoother wafer placement.
    
    Transferred a cassette of wafers from the left and right and everything
    looked good. 
    
    Time will tell.....
    
    					B shift
    
5.177Robot alignments / Right send elevatorFABSIX::R_GEESun Mar 17 1996 06:4915
	The system was down due to robot vacuum and alignment
	problems. We found all alignments needed to be adjusted.
	
	The robot to right send elevator adjustment was done. We
	found the alignments looked good at the top of the cassette 
	but the robot hit the wafers at the bottom of the cassette,
	the 25th slot was off by 1/4". We removed the elevator and
	replaced the motor and encoder, both looked bad. We ran 100
	wafers with no problems. 	  	
    
	No vacuum problems found.
    
    
    					B shift				 
    
5.178ghost wafersSTRATA::RIDLONEliminate the obviousMon Mar 18 1996 05:5410
    Alarmed several times for "wafer found on robot". Rather than adjusting
    the vacuum switch we checked the tool to make sure everything was
    working as it should. The only thing we found out of the ordinary was
    that solenoid 44 wasn't turning on when it should. When we checked the
    machine the tool was in wet idle and solenoid 44 was not energized. On
    A2 it was. Whether this is a problem or not is unclear. We cycled some
    wafers through in the hopes it would alarm but it didn't. Gave it back
    to production as they had work for it.
    
    					B-shift
5.179semi annual in progressSTRATA::FINNMon Mar 18 1996 19:326
	Semi annual pm in progress, check sheet sent to B shift


						A shift
        
5.180Purge pressure increased on cmp.a1SUBPAC::TOOKERWed Mar 20 1996 16:2510
Today I increased the robot vacuum line water purge pressure from 25 psi to
33 psi, and recalibrated the the vacuum switch.  After this, 50 wafers were
cycled with the handling recipe without any problems.  I want to see if this
helps the "wafer present on robot" error that we have been getting on cmp.a1.
If anyone sees this problem, please send me mail so that I will know if the
problem is getting better or not.

Thanks,
Jim T.

5.181System upgraded with new load and clean station ringsSUBPAC::TOOKERWed Mar 20 1996 16:268
CMP.A1 and CMP.A2 have been upgraded with new load station rings and new
6 roller clean station rings.  I have checked the alignment between the
spindles and the new rings, and cycled 50 wafers through each system with
the handling recipe without any problems.  Please be aware that if either
system starts to have misloads, that a re-alignment may be necessary.

thanks,
Jim T.
5.182Crash...SUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 21 1996 05:1021

     
          The tool crashed at beginning of shift.  It 
       was a good crash to say the least.  Not your 
      ordinary "slide out, opps, EMO, small mess ect....   
      The tool was cleaned up and a new pad was placed on  
      the tool.  Upon checking out the carrier in question, 
      I found one of the .50 cent el-cheapo check valves
      completely inoperative and closed.  After I replaced
      all of the tubing and check-valves, I actually spent
      some nice quiet quality time rebuilding a set of 
      carriers.  The tool went through pad break-in and
      passed MQC's with no scratches and good uniformity.

                 e-ha...

                      I'm going to Disneyland.....

                                        BJ
             
5.183Wafer Fly-Out & Breakage @ 7:10 p.m.ASDG::POIRIERDuuuhhhThu Mar 21 1996 19:2511
    
    * Wafer flyout & breakage at 7:10 P.M.
      ------------------------------------
      Two wafers broke at Approx 7:10 p.m.  The system was not used on the
      shift until 7:00 p.m. Ten conditioning cycles were run on the Pad
      before any wafers were run. The system was running wafers 5 & 6
      when the fly-outs occurred. We do not know what step of the Auto
      Cycle was in process when this happened as noone was watching at the
      time.   
    
      
5.184Tool had no problemsSUBPAC::BJUBINVILLENFD Fire - RescueFri Mar 22 1996 05:039
        
        This tool has carriers built by ENG with no check valves
     no back pressure in the recipes and just the outer holes open.  
     OPS broke in the pad and ran passed MQC's.  No problems were
     seen however there is only 38 wafers on the pad.

                                        BJ
             
5.18532" Pad cut down & installedASDG::POIRIERFri Mar 22 1996 19:427
    
    * Wafer Fly-out Problems:
      ----------------------
    
      A 32"Speedfam IC 1000 pad was just placed on the tool. The Break-in
      recipe is now running. The Left Carrier is has a vac leak. Please
      remove & repair or build-up a new one.
5.186no crashes but failed MQC'sSUBPAC::BJUBINVILLENFD Fire - RescueSat Mar 23 1996 05:2810

          Tool failed first round of MQC's from being down and 
       there were no over currents or slide outs.  The condition of 
       the tool is currently :

                       DOWN/VERIFY-U
                       32" pad cut down to 28"
                       check valves back in line in the carriers
                       new style housing with spacers, no jack screws.
5.187crash and burnSTRATA::RIDLONEliminate the obviousWed Mar 27 1996 06:199
    Crashed wafers twice tonight. Both times the wafers were just breakin
    wafers. Pad and carriers replaced both times. Weekly pm was completed
    while the machine was down. The second time it crashed the tool kept
    alarming for "wafer found on spindle." This was due to one of the
    carriers having a check valve installed backwards. Both of the check
    valves should be installed so flow is allowed from outside of carrier 
    to middle of carrier.
    
    						B shift.
5.188CMP.A1 shutdown for facilities tie-in completedSUBPAC::BJUBINVILLENFD Fire - RescueThu Mar 28 1996 06:0757

  SHUTDOWN:  5:00 AM ON 3/28/96

  STARTUP:   9:00 PM ON 2/28/96


                    CMP.A1 SHUTDOWN CHECKSHEET
                    --------------------------


                                                        Tech.       Date
                                                       -------     -------


  1.  Change RTC status to SCHEDULED DOWN.               BJ        3/28/96

  2.  Remove all wafer carriers.  If possible
      keep them wet.  If not ensure that there
      are enough carriers with new films for
      startup.                                           BJ        3/28/96

  3.  Remove all wafer cassettes and drain Send
      and Receive DI Water tanks.                        BJ        3/28/96

  4.  Purge both slurry lines with DI Water.             BJ        3/28/96

  5.  Remove and discard all polishing pads.             BJ        3/28/96

  6.  Remove all PVA scrub brushes and submerge         
      in DI water.                                       BJ        3/28/96

  7.  Perform a complete clean of each system
      to remove dried slurry.  Wipe down and dry
      all wet surfaces.                                  BJ        3/28/96

  8.  Remove and discard all slurry filters.             BJ        3/28/96
                                                         
  9.  Disconnect inlet and outlet lines on each
      filter housing and remove housing.  Connect
      inlet and outlet lines together with a 3/8"
      union.  This will completely bypass the 
      filter housings.                                   BJ        3/28/96

 10.  Close main incoming DI Water valve to each
      system and open the DI Water return valve
      (UDR).  This will allow all incoming DI Water
      to flow to UDR.                                    BJ        3/28/96

 11.  Power down each system.                            BJ        3/28/96

 12.  Power down each vacuum pump cabinet.               BJ        3/28/96

 13.  Power down each Neslab TCU.                        BJ        3/28/96
    
    
                                                                       
5.189Strasbaugh Cleaning Station Brush Height AdjustmentASDG::POIRIERThu Mar 28 1996 19:439
    
    Cleaning Station Brush Height Adjustment:
    ----------------------------------------
    Looks like we're going to need a new adjustment plate. The 11" dia
    plexiglass reference plate doesn't fit on the new Cleaning Station
    rollers. We need a new one 9 1/2" in Dia. We'll contact Al LeBlanc 
    and ask him to fabricate a new one. We made  a "Best-Guess"  adjust-
    ment on the brush heights for CMP.A1 & CMP.A2. Try to find something
    flat, clean & 9" to 9 1/2" in Dia. to use as a reference.
5.190WAFER CRASHSTRATA::WESTPHALThu Apr 04 1996 20:055
CMP.A1 - Down for breaking wafers on the polish table.  Broke wafers while 
	 Process Engineering was performing experiments.  Cleaned up the mess 
	 and replaced the pad.  The tool is back to production.

5.191New Strasbaugh maintenance manuals are hereSUBPAC::TOOKERMon Apr 08 1996 13:4312
We have received 2 new sets of maintenance manuals for cmp.a1 & cmp.a2.  These
new manuals are marked ver. 3.  One set has been placed in the gray cabinet
in the CMP service aisle and the other set is in my office.  These new manuals
reflect all the changes brought about by the new software rev. and the new
style robot.  The old manual containing the schematics has been left in the
service aisle.  You may find that in some cases these will need to be 
referenced.  Please be gentle with these manuals and ensure that pages are not
ripped out or lost.  These manuals cost $1200.00 per set.


Thanks,
Jim
5.192water nozzles added STRATA::FINNWed Apr 10 1996 20:155
	Added water nozzles to left and right clean station. Tested AOK

					Brendan/Hawk/Dan
        
5.193No slurry flowSUBPAC::BJUBINVILLENFD Fire - RescueWed Apr 17 1996 05:579

         During the manual procedure of getting slurry in the
     lines, large air bubbles were coming through the line.  
     The slurry system had shutdown and caused this problem.
     Ashland came in and swapped over the filters.  The flow
     was back to normal after that.

                                        BJ
5.194not picking up wafer at load stationsFABSIX::B_FINNSun Apr 21 1996 20:2711
	
	Will not pick up wafer in load cassette left or right. Cleaned 
end effector without positive results. Ran elevator and robot in 
manual to check alignment and it pick wafers up ok. Looking for leak 
and about to check cal of sensor.


					A shift


5.195robot vacuumFABSIX::R_GEEMon Apr 22 1996 08:085
	The robot wasn't picking up the wafer out of either send cassette
	in auto. Checked alignments and hoses for leaks, all looked OK.
	Calibrated the vacuum switch, ran 100 wafers with the handling
	recipe, all set.
					B shift				 
5.196exSTRATA::RIDLONEliminate the obviousTue Apr 23 1996 07:2716
    We had more robot loading problems tonight. Since the switch was just
    adjusted last night we looked around for some other root causes. We
    finally found a cut in the pressure/vac line that supplies the end
    effector. It was cut where the line connects to the robot base module.
    The fitting is quick disconnect but instead of being an elbow like A2
    it was straight fitting. This caused the air line to rub on the flex
    track. The line was cut back and the straight fittings installed. There
    isn't a parts listing in the fab for the Strasbaugh so we didn't look
    up any part numbers. The robot didn't have any more problems loading 
    after the problem was corrected. 
    
    
    						B shift
    
    
    						
5.197robot vacuum switch adjustmentSUBPAC::LANDRYTue Apr 23 1996 19:4717
      Today..........

      The system system experienced poor uniformity in the AM.
      The pad was removed/replaced.  While this was taking place,
      the system went down for scheduled DI water work to be
      performed by the contractors.  The scrub brushes and the
      carriers were removed and placed in water to remain wet
      and intact.

      While the DI work was taking place, the robot vacuum switch
      was checked and re-calibrated.  

      After the system was returned to us, the carriers and brushes
      were re-installed (as well as the new pad).  Breakin wafers
      and MQC's were run.  The system is up and running.

      					A Shift  /  Jim T
5.198Right spindle downforce problemsFABSIX::R_GEETue Apr 30 1996 12:1414
    
	The system was put down for wafer breakage and an error,
	"Right spindle downforce low". We looked at the old errors
	and found that it had errored for that and "Right spindle
	downforce high". 
	We did a quick downforce check at 300 lbs and found that it 
	was very unstable. Found a solder weld in the connector 
	coming out of the amplified was bad. Removed and resoldered 
	all the wires in the connector coming out of the amplifier
	and going into the air cylinder, they all looked bad. Performed
	the downforce calibration on the right spindle and checked
	the left. Installed a new pad and carriers, the system is up.
 
					B shift				 
5.199leaksLUDWIG::WESTPHALWed May 01 1996 20:305
                
CMP.A1 - Down for a leaking slurry line on pump #1 and a DI water leaking from 
	 the spray nozzle.  Replaced the slurry line and repaired the DI
	 nozzle.
     
5.200Downforce calibration performed on cmp.a1SUBPAC::TOOKERThu May 02 1996 16:317
Today John Swenson, Brad Wesphal and I did a complete down force calibration
including E-P regulators on cmp.a1.  We found that both E-P regulators were OK.
The left spindle was out of calibration by around 1% (not much), and the right
spindle was on the money.  After the calibration, a new pad was installed.

Jim T., John S.,& Brad W.
5.201Right load cup FABSIX::R_GEEMon May 06 1996 08:1010
	Broke a breakin wafer on the pad. 

	The Wafer slid out from under the right carrier at touchdown.
	There was 240 wafers on the pad, 285 wafers on the left 
	carrier and 345 wafers on the right. While it was down the 
	weekly p.m. was done and the carrier to load cup alignment 
	checked. We adjusted the right load cup slightly so the
	carrier will seat smoother. 
 
					B shift				 
5.202Fairchild E-P analog counter settings for calibrationsSUBPAC::TOOKERFri May 10 1996 13:2242
To all CMP ET's;

Here are the proper analog counter settings for calibrating the Fairchild E-P
transducers for left and right spindle downforce on cmp.a1/.a2.  These analog
counter settings apply only to the spindle downforce E-P's and do not apply
to the left and right backpressure or conditioning downforce E-P's.  Please
refer to the Strasbaugh manual for the proper counter settings for those E-P's.
Prior to any spindle downforce E-P calibrations, please check the E-P to see 
which model number it is and then use the appropriate analog counter setting
listed below.  The adjustments work the same on both models.  Adjust the
"ZERO" pot for 10.0 PSI and the "SPAN" pot for 80 PSI.  When either the "ZERO"
or "SPAN" is radically adjusted, it is sure to through out the other one, so
be sure to go back and forth between the two until the proper PSI is achieved.

You may have noted that the Strasbaugh manual calls for the "ZERO" pot to be 
adjusted to 5.0 PSI.  We adjust our tools to 10.0 PSI.  We do this (with Stras-
baughs blessing) to bias the spindles with a few pounds of downforce, so that 
when the spindles are manually lowered over the load station or cleaning sta-
tion they don't float upwards over time.

Please contact me with any questions or if anyone needs more one on one training
for this calibration.
================================================================================

      CMP.A1/.A2 LEFT OR RIGHT SPINDLE DOWNFORCE E-P TRANDUCER CALIBRATION
      --------------------------------------------------------------------

        FAIRCHILD E-P                    ANALOG COUNTER SETTINGS
                                         -----------------------
           MODEL #                10.0 +/- .2 PSI        80.0 +/- .2 PSI
        -------------             ---------------        ---------------

         Z18109-3                      580                    3410
        (old style)

         TT6000-06                     601                    3481
        (new style)

    NOTE:  If any of the E-P's are re-calibrated, then the downforce must
           also be re-calibrated.

================================================================================
5.203scratchesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat May 11 1996 06:179
 
      Failed for scratches during MQC.  No-touch showed scratches
   from brushes as well.  After the brushes were replaced, no-touch
   showed no scratches but touch down had a bunch of scratches.  The
   pad was replaced and was broke-in again.  MQC's were run and there
   were no scratches and passed uniformity.

                                       D shift
5.204What's up A1 is down.FABSIX::L_SHKOLNIKOVFri May 17 1996 08:386
                             
  The system went down this morning - left carrier about 1/2 inch lower
  at load and unload then the right one causing robot to hit the edge of left
  carrier when loading the wafer.
      
                         
5.205Robot position over load variable adjustedSUBPAC::TOOKERFri May 17 1996 16:2616
This morning the top of the robot arm was hitting the carrier on the left
spindle.  This problem was caused by two things.  One, the gap between the
top and bottom plates of the carrier vacuum chuck was set a little wider
than usual during the monthly PM hub alignment.  Two, the left spindle
sits a little lower than the right spindle in the up position and always
has.  This was not critical with the old style robots, because the robot
arms were much lower than the new style robot arms.  The short term solution
to the problem was to increase the value of calibration variable 130, robot
position over load.  This sets the robot arm at a lower position when moving
the wafer in over the load chuck and under the carrier.  The long term
solution will be to install new style turnbuckles on the lever arms (an upgrade
already planned) and then adjust the left spindle up position to give a 
greater clearance between the top of the robot arm and the bottom of the 
carrier.

Jim T.
5.206Bad UniformitySUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu May 23 1996 06:598

       Problems all night with uniformity.  After two pad changes,
    we are currently using a pad from lot #6DP-JP1.  The conditioner
    disk was changed as well.  Carrier #1 was swapped out but the 
    uniformity on both were 16% after three tries.

                                               D shift
5.207Bubbles....SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed May 29 1996 08:197
       Huge bubble formed in middle of pad at 162 wafers. 

             replaced pad....

                                          BJ
       
5.208Slurry line blewSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Jun 01 1996 05:256
 
         The slurry line on pump #1 blew and sent slurry everywhere.
      Leon replaced both lines with the surgical style tubing.

                                         BJ
5.209SCRATCHESFABSIX::J_SWENSONSat Jun 01 1996 20:496
    REPLACED PAD.
    REPLACED CONDITIONER PAD. MQC HAD SCRATCHES REINSTALLED ORIGINAL
    CONDITIONER PAD AND REPLACED SCRUBBER PAD.
    MQC'S THEN PASSED.
    
                          DOC AND HAWK  
5.210broken waferFABSIX::R_GEESun Jun 02 1996 07:538
	Broke a wafer on the pad. 

	The Wafer slid out from under the left carrier during a 5 sec.
	t-up run. There were 134 wafers on the pad and 408 wafers on the 
	carriers, the spec is 400 wafers on the carriers. The carriers 
	were to be changed when this lot finished.
   
					B shift				 
5.211High particles and scratchesFABSIX::R_GEEMon Jun 03 1996 08:1616
	The system has been down all night due to high particles
	and scratches. We did a run not touching the pad and the
	particles level was good. We replaced the pad and conditioning 
	disk, still failed twice for particles and scratches, 5000+.
	We changed the pad and conditioning disk again and ran the 
	old breakin recipes to see if the new breakin program with 
	25 lbs of conditioner down force was causing the high particles.
	Still out of spec. no scratches but 3000+ particles edge heavy. 
	Tried another no pad touch and it also looked good, we are still 
    	looking at a pad area problem.
	The slurry delivery system went down at 5:00am for clogged filters,
    	could this be contributing to the high particle counts. We are
    	running another set of mqc's now.
    
    
    					B shift				 
5.212strasbaugh voodoo loungeSTRATA::WOOLDRIDGEPleasure, Spiked With PainMon Jun 03 1996 16:5710
    you were right Mitch/Bob- changed A1 conditioner force back to 13 where
    it was supposed to be.  A2 was correct at 13.  Fould pad particles on
    the pad & has center fast profile (makes sense at this force).  Changed
    pad and we will see what we have. (used a different pad batch for the
    heck of it)-  Good catch guys- hope this fixes it.  mqc'S showed little
    improvement prior to this change by the way (particles in the 1500
    range)    IF this does not work out, do you suppose this has anything to
    do with the broken wafer a day or so ago?  Maybe slicon in the vacuum
    lines?  dunno.
                     Kyle
5.213Particles high...but improvingSUBPAC::LANDRYMon Jun 03 1996 20:317
    The system still has high particles.  The latest run today showed a big
    improvement, but still on the high side.  There is optomism that a few
    more break-ins and another MQC try may bring it in.  Unfortunetly,
    there are NO MQC wafers available in the room.
    
    						Doc
    
5.214Pad & Conditioning Disk ChangedASDG::POIRIERThu Jun 06 1996 20:225
    
    Pad Change:
    ----------
    The Pad and Conditioning Disk were changed at 7:00 p.m.
    
5.215Wafer slid outFABSIX::R_GEETue Jun 11 1996 07:509
	A wafer slid out from under the right carrier, no wafers
	broke but 1 was scrapped due to scratches. It was the 1st
	run of t-ups, 9 sec. polish time. We believe it wasn't seated
	properly in the carrier as it slipped out when the vacuum 
	released. We checked all the load cup to carrier alignments	
	and ran 24 dummies with no problems, gave the system back
	to production. No problems with it the rest of the night.

						B shift
5.216filtersFABSIX::R_GEEMon Jun 17 1996 08:144
	The wafers wouldn't blow off into the unload cups after
	polish, they stayed on the carriers. We replaced the two
	in-line filters, all set.
						B shift
5.217bubbles on padLUDWIG::RIDLONEliminate the obviousMon Jun 24 1996 07:244
    Found 5 bubbles on the pad. Pulled and replaced.
    
    					
    					B-shift
5.218particles and uniformity after pad changeFABSIX::J_SWENSONFri Jun 28 1996 20:125
    System failed for particles on left spindle and uniformity on both
    spindles.
    
    Found scrub brushes not touching wafer surface, adjusted to proper
    height. 
5.219alsoFABSIX::J_SWENSONFri Jun 28 1996 20:122
    more wafers are being run to bring in uniformity.
    
5.220Squeaking Noise In BridgeASDG::POIRIERSat Jun 29 1996 19:5710
    
    * Squeaking Noise In Bridge:
      -------------------------
      We found the source of the squeaking noise to be the Right Lever Arm
      Air Cylinder Pin (top). The Pin and Rod End assembly were lubricated
      with a few drops of light oil. The Left Lever Arm Pins were also
      lubricated. All's quiet for now.
     
    
                                                         C Shift.
5.221carriers changedSTRATA::RIDLONEliminate the obviousSun Jun 30 1996 07:017
    Left carrier sticking to pad. Removed carriers and found the ring on
    the left to worn all the way around and the right not much better. They
    both had only 234 wafers on them. Changed them and the maching passed
    MQC's with out incident. 
    
    				
    					B-shift
5.222Start up A1LUDWIG::SWENSONFri Jul 05 1996 20:363
    POwered tool up with no problems
    New carriers brought into room for install.
    scruber brushes ordered and retrieved
5.223wafer breakage issuesSUBPAC::LANDRYTue Jul 09 1996 20:3218
      Today, the system broke a wafer early in the shift.  No apparent
      reason could be identified.  The mess was cleaned up, new pad and 
      carriers installed.  Breakin wafers were run and MQC's performed 
      and passed.  The lot was then placed back on the system, CRASH.
      Did a spindle flush and a bunch of particles came flying out.
      The particles turned out to be silicon, most likely sucked up
      from the breakage.  Cycled the flush procedure sevral times.
      Installed two new carriers and a pad.  Lets try again.

      Also today, found the chiller reading ~15 C at the TCU.  The TCU
      was stuck in the "cooling" mode.  Pressed the display to view the
      set point and the gauge was fluctuating all over the place.  Gave
      the adjusting knob a bunch of turns and the fluctuating went away.
      It is now stable and the TCU readout is responding properly.
      Looks like the controller may need to be cleaned in the near future.

      						A shift
  
5.224SCRATCHESFABSIX::J_SWENSONFri Jul 12 1996 20:249
    SAME AS CMP. A2.
    
    ENDED UP REPLACEING THE PAD AND CONDITIONING PAD.
    MQC'S NOW BEING PERFORMED.
    
    NOTE: LOOKS AS THOUGH THE DISC'S FROM THE STOCKROOM HAVE PROBLEMS.
    RECOMMEND DO NOT USE.
    
                      C-SHIFT
5.225scratchesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Jul 13 1996 07:0612
       Down for large scratches....  I used a modified
   no-touch program to include a brush cleaning cycle
   and everything was fine.  As soon as a touch-down on
   the pad was involved, scratch city.  The conditioner
   looked ok but seems to be a suspect.  I'm going to pull
   the pad and have OPS break-in the pad and check MQC's
   at their next open opportunity.

                                           BJ
                                  
                                      
5.226Large scratches on wafers / conditioning disks suspectedASDG::POIRIERSat Jul 13 1996 20:4810
    
    Scratches on wafers:
    -------------------
    We are continuing to see large deep scratches on the wafers. Repeated
    conditioning disk and pad changes yield no change. We highly suspect
    the latest batch of conditioning disks are responsible for the
    scratches.
    
                                                    C Shift.
    
5.227High particles and scratchesFABSIX::R_GEEMon Jul 15 1996 08:0810
	Saturday the system was put into Engineering in workstream. 
	This was done because of high particles and scratches believed 
	to be coming from the new shipment of conditioner disks.
 
	We did the monthly p.m. replaced the pad and conditioner.
	Before the conditioner disk was installed it was deburred
	with a small hole file, wire brushed and cleaned. The system
	passed particles and uniformity on the first try, returned
	it to production.
						B shift
5.228stores inventory look ok?STRATA::WOOLDRIDGEPleasure, Spiked With PainMon Jul 15 1996 10:483
    Has anyone looked at the currect inventory in stores to see if we have
    inventory with the "burr" problems we saw some time ago?
    /Kyle
5.229Left spindle water leakFABSIX::R_GEETue Jul 16 1996 08:2010
	The system failed MQC's for uniformity and head to head removal,
	there was over 1000 ang difference between the two. Found a
	small water leak coming from the left vacuum line filter, the
	water had dripper onto the servo amplifier. We removed, dryer
	and reinstalled the amplifier and connectors. Also tightened the
	filter. We checked the down force and it was in. 

	Had to change the left brush due to high particles.

						B shift
5.230UNIFORMITY / PARTICLESSTRATA::WESTPHALThu Jul 18 1996 20:506
CMP.A1 - DOWN FOR PROBLEMS WITH UNIFORMITY AND PARTICLES.  THE CONSUMABLES
	 WERE SWAPPED OUT, AN O-RING OFF THE LEFT SPINDLE WAS REPLACED, THE 
	 LEFT SCRUB BRUSH ADJUSTED, AND THE TOOL WAS CLEANED UP AND RETURNED TO
	 PRODUCTION.  

5.231PARTICLESSTRATA::WESTPHALSat Jul 20 1996 20:374
CMP.A2 - Down for high particles on the right carrier.  The particles were at 
	 3000.  Changed out the right brush and the particles came in at 1100. 
	 Returned the tool to production.
	  
5.232UNIFORMITY,PARTICLES AND SCRATCHESFABSIX::J_SWENSONThu Jul 25 1996 20:147
    DOWN FOR UNIFORMITY ,PARTICLES AND SCRATCHES.
    REPLACED BOTH CARRIERS, AND PAD.
    CLEANED TOOL AND BROKE IN PAD.
    RETURNED SYSTEM TO PRODUCTION.
    
    
                                         C-SHIFT
5.233robot tubing replaced.....SUBPAC::LANDRYMon Jul 29 1996 19:4310
    The robot was failing to pick up the wafers properly today.
    Found a small gash in the 1/8" tygon tubing, about an inch from the 
    end effector connection.  The tubing was replaced and results were
    good.  
    
    The polish pad was changed out this morning.  Breakin's and MQC's were
    run.  System came in on the first attempt.
    
    						A shift
    
5.234bubblesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Aug 02 1996 05:266


            the pad was replaced due to a large bubble in the 
         middle.
                                         D shift
5.235SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringMon Aug 05 1996 14:486
    
             two break-in wafers broke on the table for
          an unknown cause.  A new pad was installed and
          turned back over to OPS.
                                              A shift
          
5.236Pre-punched films- ok to run prodSTRATA::WOOLDRIDGEPleasure, Spiked With PainThu Aug 08 1996 12:414
    fyi:
    
    Two pre-punched carrier films are on the carriers on A1. They tested
    out nicely at 6% and 8% non-uniformity.  
5.237Left spindle FABSIX::R_GEEMon Aug 12 1996 08:277
	The robot was hitting the bottom of the left carrier when loading 
	and unloading wafers at the left load cup. We adjusted the left
	spindle air cylinder height to match the right, the left carrier 
	now is up 1/2", even with the right. The monthly p.m. also was
	completed. The prepunched carriers were kept wet and placed back
    	on the tool.  
						B shift
5.238RobotFABSIX::R_GEETue Aug 13 1996 08:167
	The robot arm wouldn't home, had to reboot the system 
	to get it to homed.
		
	The robot wouldn't pickup wafers off the left load cup.
	Changed calibration variable #131 from .88 to .92
		
				B shift
5.239Carriers sticking to padFABSIX::R_GEETue Aug 27 1996 07:537
	Carriers were stiking to the pad. We found that on both
	carriers the 3 screws that hold the backing plate to the
	carrier plate were loose, causing the carrier not to seat 
	properly into the backing plate, (finger tight is to loose).
	Replaced and tightened the screws, no more problems.
		
				B shift
5.240POOR UNIFORMITYSTRATA::WESTPHALSat Sep 07 1996 20:223
CMP.A1 - Down for poor uniformity.  Replaced the pad.  

5.241A1 DOWN TO PROCESS THIS WEEKSTRATA::WOOLDRIDGEPleasure, Spiked With PainTue Sep 10 1996 16:157
    A1 is down to process for EX1400 pad experiments most likely for the
    rest of the week.  Please leave the machine as is.  If the machine is
    needed in an EMERGENCY for production (very last resort) is ok to pull
    the pad AND CHANGE THE PLATEN TEMP BACK TO 38'. Now it is @ 22'. 
    Please only do this if the tool is REALLY NEEDED.  I do not want to
    change this pad unless absolutely required.  
                                                   -kyle
5.242load chucksSTRATA::FINNTue Sep 10 1996 20:365
	Installed new left and right load station chucks.  
        Reprogrammed robot. Ran 30 test wafers aok in handling program.

						A shift
5.243A1 still down to processSTRATA::WOOLDRIDGEPleasure, Spiked With PainTue Sep 17 1996 16:266
    Still down to process engineering madness.  If any pm work is done
    please be sure to keep the pad & carriers wet.  If ops HAS to have the
    system because they need the capacity, the pad, carriers, conditioning
    disc should be changed and the chiller temp should be brought up to
    match CMP.A2.  
                   Kyle
5.244CONDITIONING ARM TROUBLESTRATA::WESTPHALFri Sep 20 1996 20:097
CMP.A1 - Down for the conditioner spindle falling off.  The expansion pin was 
	 split and worked its way out.  Put a new pin in from stock.  While 
	 testing the conditioner it was noticed that there was excessive play 
	 in the C. arm while oscillating.  J.P. recommends changing the "rod
	 ends" on the C. arm actuator.  The parts have already been received
	 from stock.  They're sitting on remote cart.

5.245Bent conditioner spindle shaftSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Sep 21 1996 06:5318
        
    
          The rod ends on the conditioner actuator were replaced 
      and the "play" now is identical to A2.
     
          When conditioning the pad, the conditioner assembly is 
      jerking and jumping as the arm sweeps on the pad.  Upon 
      disassembly of the conditioner shaft and bellows, the shaft, 
      when told to rotate is bent slightly.  When the conditioner 
      is on the pad with force, the conditioner is not on the same 
      plane as the table and appears to be causing the skipping 
      and jumping.  We couldn't find the parts so Strasbaugh will 
      have to be called in the a.m. or as late as Monday for the 
      shaft assembly.  It is possible that when the spindle section 
      fell off on days, the bearings inside on top might have got a 
      good jolt as well.  An entire assembly is recommended...
    
                                             D shift
5.246Conditioner trouble continuedLUDWIG::WESTPHALSat Sep 21 1996 20:4210
CMP.A1 - Down for problems w/ the conditioner.  Tool never came up from
	 yesterday when the conditioners's end effector fell off.  After 
	 putting the arm back together there was excessive play in the C.arm
	 while Oscillating.  Previous shift suspected that the shaft was bent.
	 The shaft did'nt appear bent to me so I cleaned everything and put new
         split pins in and reassembled the arm.  Cycled the conditioner about
	 30-40 times in all w/ no problems.  Put the old brushes and carriers
	 back on the tool w/ a new pad.  Conditioned and broke in the pad.  The 
	 tool sets needing only MQCs.

5.247robot missing right loadLUDWIG::RIDLONEliminate the obviousSun Sep 22 1996 05:466
    Robot not picking wafers off of right load chuck. Everything looked
    good operationally except that it didn't appear that the robot was
    going down far enough before it alarmed. Recalibrated #132 "robot up/dn
    right load pos'n" and robot worked fine after that. 
    
    					b-shift
5.248pad/brushes replacedSTRATA::RIDLONEliminate the obviousSun Sep 22 1996 07:334
    Failed particles twice. Replaced pad and brushes. Over to production
    for mqc's.
    
    					B-shift
5.249particles persistSUBPAC::LANDRYSun Sep 22 1996 20:0514
    High particles persist.  Here's what transpired today.
    
    *  Changed the pad again.  Tried a different lot # from the stockroom.
    *  Cleaned all surfaces thoroughly.
    *  Flushed the spindles (carriers off hubs at the time) for 10 minutes.
    *  Flushed the slurry lines for 10 minutes.
    *  Verified brushes making contact in clean station.
    *  Confirmed (using the handling recipe) that the particles are coming
       from the polish area.  The handling run added just a fewe particles.
    
    At shift change, the particles are in the 2500 - 3000 range.
    
                                             A shift
    
5.250Fuel for the fireLUDWIG::RIDLONEliminate the obviousSun Sep 22 1996 23:488
    Tried something different. Made a TEST recipe that only goes up
    to 75lbs DF with vacuum ON in all 3 steps and step 1 is 20 sec,
    step 2 is 30 sec, and step 3 is 10 sec. The right wafer removed only
    455 angstroms and the left 136 agstroms. Particles on the left were 
    0 large and 565 total and 0 scratch count, the right were 0 large,
    1575 total, and .67 scratch count.
    
    						B-shift
5.251new carriers/no particlesLUDWIG::RIDLONEliminate the obviousMon Sep 23 1996 05:306
    As a last resort we decided to build two new carriers with new
    film from stock. First MQC passed with the left at 30-359-0 and
    the right at 56-556-0. Both removal rate and std dev were in spec 
    as well. 
    
    					B-shift
5.252problems found during the PMSUBPAC::LANDRYMon Sep 23 1996 18:4227
    First off, good job by B shift last night.  Glad I didn't have to
    return to the particles issue.  The tool had some idle time today,
    so the weekly pm was performed.  Whilst performing said pm, a couple
    of issues arised:
    
    *  The chiller temp was sitting at 65 degrees C.  Checked the Neslab
       set point.  It was jumping all over the place (+ to -).  Opened up
       the lid to check the electronics.  All connectors/wiring was good.
       Next, tried adjusting the set point.  Seems the pot for the temp
       adjust was erratic.  The contact point must have be oxidized as
       repeated rotation of the pot finally settled everything out.  Temp
       soon returned to normal.
    
    *  Whilst performing the slurry cal, the hose_to_SS tubing connection
       was spewing everywhere.  Cut back a section on the clear tubing and
       retried.  Still spewing.  Found that the SS tubing was clogged solid
       at the discharge end.  Unclogged the tubes to solve spewing problem.
       Cal'd slurry pumps, no adjustment to the variables was necessary.
    
       Cleaned up all dried slurry in and around the polish area.  Also,
       a major amount of dried slurry on the carriers/hubs was cleaned.
    
    *  The condition arm "dog-dish" was overflowing.  Found the drain hose
       to have "arc'd" above the height of the dish.  Repositioned the
       hose.
    
    						A shift
5.253particles 2, techs 0SUBPAC::LANDRYWed Sep 25 1996 20:3910
    The system is currently down for particles.
    
    Earlier in the week, we experienced the same problem.  When the
    particles showed up today, we changed out the pad and rebuilt two
    carriers.  Breakin's were running fine, when a wafer slipped out of the
    right carrier, crash.  Cleaned up the mess, installed a new pad and
    again, built a set of carriers.  Breakins, MQC's were run.  Particles
    are still high.  Thats where we sit.  Good luck!
    
                                               Doc  /  Jim S
5.254ParticlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Sep 26 1996 07:4818
 
         Tool down all night for high particles.  Tried the following:
  
         changed pad
         cleaned tool
         changed brushes
         checked brush heights
         changed and wire brushed conditioner disk
         checked recipes and programs
         checked slurry flow
         checked surfscan recipe
         checked particles with B2, B2 below a 1000, A1 ~4500
         
         latest move: changed pad with one from new lot...  awaiting
         MQC results.....
                                    D shift
         
5.255cmp A-tool particle updateSTRATA::WOOLDRIDGEPleasure, Spiked With PainThu Sep 26 1996 16:5778
     Strasbaugh particle problem update: 9/26
-Kyle

Today:

o  Repeated no touch recipe (eliminating IC pad) and particle counts were good.

o  Particle checks on both machines were in the 4000-6000 range

o  Because of low pad count on A1, pad was conditioned for an extended 
   time.  New MQC was run- no change.

o  Changed conditioning disk on A1- counts went up? - put old disk back on

o  Pad was scrubbed down with a brush and rinsed repeatedly on A1. No 
   change.

o  Noticed that the MQC wafers after polish looked like the polish pad?  
   This is the first time I have seen this.  The wafers show the pad 
   perforation marks perfectly.

o  Looked at pad expiration dates of the IC lots we have in the 
   drawer- one expired Aug. 27th and the other expires Oct 1st.
   (Rodel claims pads are good 6 months after expiration date)

o  Decided to change pad on A1 and install a pad with expiration date 
   of Jan 97'.  Conditioned & broke-in the pad with NEWPAD recipe.
   Ran MQC's and the counts were 4288 & 3323.  Interestingly, no pad 
   perforation marks were seen on these wafers however.
   Followed this by running a modified MQRINSE recipe: Defects went 
   down to 200.

o  Ran a modified MQC recipe with extended rinse at the end of the 
   recipe on CMP.A2.  DEFECTS WERE REDUCED TO 267 on one head and 245 on the 
   head. (no pre-measurements were taken so in fact the defects were 
   actually lower) Specifically changed was:
                                   Step 3: 20 seconds polish time
                                           40 rpm spindle speed
                                           40 rpm table speed
                                   
                                   Step 4:
                                           60 seconds polish time
                                           20 rpm spindle speed
                                           60 rpm table speed
                                  
The difference here is simply the addition of a step 4 rinse time with 
different spindle/table speeds & changing the final rinse step from 
10 seconds to 60 seconds. The recipe name is MQRINSE.  

o  John P. has cut a speedfam pad down to try on A2.  There is also 
   another pad from the batch now on A1 on top on the pad cabinet. 
   (good through Jan '97) if we want to look at that.

o  Speedfam tools defectivity OK?


     Musings:
              
- defects are hard to remove, but an extended rinse brings defects 
into a very low level.

-Pads may be partially the problem- but this is not fully 
understood since the new pad on A1 did not qual.

-It appears for what ever reason, the slurry is NOT rinsing off the 
wafers. This could be related to the pad/polishing OR to something 
going on with the slurry/delivery system.

-It appears we can run the process currently, but the process recipes 
will need to be modified to reflect steps 3 & 4 in the MQRINSE recipe.
A TDCO/edco will need to be put in place tonight (pss) in order to 
reflect this change until the process/hardware fix is fully 
understood.  This is simply an extension & addition to the rinse step.

If needed, I can be reached at home @ 464-5721

/Kyle
5.256It's just to many of them.FABSIX::L_SHKOLNIKOVFri Sep 27 1996 07:2615
We had some problems tonight with a particles. So we run a modified MQC
recipe MQRINSE on CMP.A2 with good results  30-250-0 & 46-483-0.
In order to verify the slurry system we installed 1 micron filters, one per
pump on CMP.A1 and tool came up with following results 75-827-0 & 88-1367-0.
CMP.A1 passed MQC's and released to production.

Well.. Now we changed MQRINSE on A2 back to old MQC recipe and defects
went up to 6000, with micron filters installed tool passed particles
with flying colors and results: 43-256-0 & 63-1289-0. Running full MQC's
right now.


                                                     D-shift
        
5.257cond param wrongSTRATA::RIDLONEliminate the obliviousSat Sep 28 1996 21:3110
    Checking on tools at start of shift and noticed that A1 was
    conditioning with DI water. Checked A2's condition parameters and it
    was still set to slurry 1. Checked notes for any changes that might
    have been made and found none. Called Kyle at home and he was
    surprised. Changed the condition parameters back to slurry 1 with
    500mls/min and conditioned the pad 10 times. FYI, there were 200 wafers
    on the pad. 
    
    
    					B-shift
5.258crash recoverySTRATA::RIDLONEliminate the obliviousTue Oct 01 1996 05:247
    Came in tonight with the machine down from a crash and no carriers
    available. Built two carriers however they didn't have enough wafers
    with enough oxide on them to breakin the machine. They were able to
    scrounge some up and are in the process of breaking it in now.
    
    
    						B-shift
5.259Right carrier problemFABSIX::R_GEETue Oct 01 1996 20:426
	The right carrier kept falling off the right hub. 

	Checked the inner hub and carrier for leaks. Replaced the
	hub o-rings and ran verification wafers, all set.

				A shift
5.260scratches and particlesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Oct 03 1996 07:2014
       Tool failed for nasty rug burns.  MQC's were re-run and
    the problem was still there.   It looks like the wafers were
    dragged off of the load pedestals.  We checked the robot
    heights on dummy wafers and it was fine.  checked for particles
    and there were ~5 areas on the pad that had pieces of either
    film or silicone.  After cleaning the entire tool, MQC's were
    run again and failed.  The pad was replaced and the tool once
    again failed... the scratches are gone but now its failing for 
    particles, 1900 and 1600.  Cleaned and scrubbed entire tool and
    we gave it back to them for one more shot.


                                          D shift
5.261follow up FABSIX::O_WHITFORDThu Oct 03 1996 08:234
    Tool passed MQC's with less than 1300 and 1200 respectively. No
    scratches.
    
    
5.262low polish rate update- slurry statusLUDWIG::WOOLDRIDGEPleasure, Spiked With PainThu Oct 03 1996 17:4619
Low CMP polish rate- UPDATE:

     A new batch of slurry (lot #2h-16-96-1) was manually pumped to 
CMP.A2 due to the problems we are currently having with low rates.
The results were:
                 head 1-1300
                 head 2-1115
Directly after this, the slurry from the day tank was hooked back up 
and the results were as follows:
                head 1-1132
                head 2-1054
Based on the above, I have asked Ashland Chemical to dump the current 
day tank and replenish the slurry with the new batch.  We should have 
new slurry by 6pm or 7pm tonight.  The change over should be transparent
to operations.  Unfortunately, there is only enough slurry of this 
batch to tide us over through the weekend.  Then a new batch will have 
to come on line.
                   - Kyle   
5.263crashSTRATA::RIDLONEliminate the obliviousSun Oct 06 1996 01:577
    Tool had crashed just before passdown. Did a visual and could find no
    evidence for a reason. Pad was flat, it was product, slurry in lines,
    and no way to tell if it was the carriers or not. Changed everything
    except for the conditioner disk. Pad cond x10. Gave back for
    bkin/mqc's.
    
    					B-shift
5.264carrier screws looseSTRATA::RIDLONEliminate the obliviousSun Oct 06 1996 05:548
    Left removal was low and std dev was high at 28. Right was in spec.
    Swapped the carriers and then the left stuck to the table. Took apart
    and found the three carrier plate screws only hand tight, not good.
    Rebuilt and reinstalled. Took apart the right one as well and
    everything looked fine there. Back to them to finish up mqc's.
    
    
    						B-shift
5.265df calLUDWIG::RIDLONEliminate the obliviousMon Oct 07 1996 03:3923
    A shift changed the right carrier and tool passed mqc's. B shift
    ops did send aheads and the left carrier had a removal rate in the
    400's w/ the right being fine. They redid mqc's and the left removal
    was 920, the right 1620. Swapped the carriers. Left remained at
    920 and right remained at 1620. Down force cal was performed. Both
    needed adjustments, however the right was the farthest out.
    
    				  OLD DATA         NEW DATA
                                  -------          -------
    	Left spindle force error = 1.07              1.00
     
	Rght spindle force error = 1.07               .91
    
    	Left dead weight variable= 162.18            160.75
    
    	Rght dead weight variable= 162.62            165.74
    
    
    A new pad was installed and conditioned. The tool is ready for MQC's,
    however, as I understand it, ops is having a hard time getting wafers.
    
    
    						B-shift
5.266Spindle Downforce Cal. FABSIX::R_GEEMon Oct 07 1996 20:369
	The system broke two wafers at the beginning of the shift,
	also the head to head removal was high. 
	We performed the weekly p.m., calibrated the Fairchild 
	regulators and spindle downforce calibration. Process elected
	not to MQC the system because there are very few MQC wafers
	available. There are new carriers and a pad on the system,
	the pad needs breakins run and full MQC's.

				A shift
5.267oh what funSTRATA::RIDLONEliminate the obliviousTue Oct 08 1996 06:478
    Failed mqc's twice for high particles on the right carrier.
    Changed the carrier and on then the mqc wafer from the right carrier
    was left on the pad. Disassembled the carrier and replaced the carrier
    plate screws due to the orings being chewed up. Returned system to
    production.
    
    
    						B-shift
5.268exSTRATA::RIDLONEliminate the obliviousTue Oct 08 1996 08:3310
    Particles on both carriers jumped to over 4000. Checked history 
    in notes and found same problem a few days ago. They ran MQRINSE 
    with good results. We ran MQRINSE with excellent results.
    Particles on left dropped down to 86 total, Right was 69 total!!!
    A side note, the left carrier was sticking to the pad during 
    MQRINSE. It stuck once running the MQC recipe and was taken apart. The
    plate screw were replaced but everything else looked good.
    
    
    						B-shift
5.269software installedSUBPAC::LANDRYTue Oct 08 1996 20:3412
      New software was also installed today.  The current Rev is now 
      version 4.0.  After the software was installed, we came across
      a couple of problems.  We needed to change the state of some of 
      the calibration variables to correct.  There currently exists
      two problems which will need to be addressed.  The Send tank dump
      valves are not responding correctly all the time and the "dunk"
      performance of the send cassettes is off.  We are currently working
      the issue with Strasbaugh.  These problems do NOT necessitate that
      the system remain down.  A warning has been logged in RTC.

      						A shift

5.270Send cassette water dunk featuer now operationalSUBPAC::TOOKERThu Oct 10 1996 16:487
Today the old revision of the Strasbaugh software (4.16.3) was reloaded onto
cmp.a1 & cmp.a2.  This software version will allow us to use the send cassette
water dunk feature.  Strasbaugh is implementing a fix to software version
4.20 that will give all the new features of 4.19 plus the send cassette water
dunk feature.  All warnings have been removed from cmp.a1 & cmp.a2.

Jim T.
5.271Carriers,brushes,particles etc.FABSIX::L_SHKOLNIKOVFri Oct 11 1996 08:1415
	The system has been down all night due to high particles
        We checked the slurry by using micron filters and slurry was fine.
        We replaced the pad and brushes and still failed twice for 
        particles, 2000+. While changing the brushes I found a carrier
        retaining ring screw under the clean station roller and a white
        colored debris all over the left clean station. Found four more 
        screws in left clean station. All 6 rollers of the left clean
        station are damaged by a lose screw and needs to be replaced.
        We do not have them in stock. Replacing the left carrier right
        now and try to convince ops run set of mqc's one more time.

                                                                   
                                                        Leon.            
                                                                         
5.272particles.FABSIX::J_SWENSONFri Oct 11 1996 20:353
             Had high particles on left carrier, swithed carriers problems
    followed replaced carrier cleaned tool still high particles. Eng
    increased di rinse 25 seconds along with a edco particles passed.
5.273WAFER CRASHLUDWIG::WESTPHALSat Oct 12 1996 20:585
CMP.A1 - Down for breaking wafers.  Barry filled out the crash sheet and
	 assisted in cleaning up tool and changing consumables.  The tool 
	 presently sets waiting for MQCs.

5.274particlesLUDWIG::RIDLONEliminate the obliviousSun Oct 13 1996 03:0910
    Failed mqc's for particles X2. No pattern. Had them run mqrinse and
    particles came down considerably but they failed for scratch count. Had
    some spots on the edge with high concentrations. Scrubbed the laod cups
    and checked the pad for any defects around the perimeter or inside
    edge. Found some questionable black marks that I tried to scrub off but
    the brush I was using was full of wafer bits. Its been thrown away. 
    The pad was replaced. Waiting on mqc's.
    
    
    					B-shift
5.275recipe problemsLUDWIG::RIDLONEliminate the obliviousSun Oct 13 1996 05:267
    Failed mqc's again. A2 passed. Compared MQC recipes and they WERE NOT
    THE SAME. Rinse step 3 was 25 seconds as specified by EDCO but the
    spindle and table speeds differed. A2's were both 40 and A1's were 60
    and 20 respectively. Another set of mqc's will be ran.
    
    
    						B-shift
5.276passedLUDWIG::RIDLONEliminate the obliviousSun Oct 13 1996 21:572
    The mqc's passed after the mqc recipe was changed to match A2's.
    The left had a total of ~1000 and the right ~500.
5.277a multitude of stuffLUDWIG::RIDLONEliminate the obliviousMon Oct 14 1996 08:1215
    Tool failed mqc's twice for poor uniformity. The numbers were close so
    another set was done and they passed. The first two dummies of a pmd
    lot crashed, however, and the system went back down. After the tool was
    cleaned up and returned to ops we got an "Axis 1 servo amp fault".
    The tech manual suggests powering down the tool to reset (first option
    of opening and closing the doors didn't work) which corrected this 
    problem. AFter that the robot wouldn't pick up the wafer out of the
    left load bucket. The air fittings on the robot were loose plus the
    wafers had fine silicon chips on them, possibly coming from the new
    carriers. EVerything was cleaned off and the spindles were flushed. 
    No evidence of wafer chips could be seen on the pad so the breakin was
    continued. MQC's wont be done until A-shift.
    
    				
    						B-shift
5.278fyiSTRATA::RIDLONEliminate the obliviousTue Oct 15 1996 02:592
    Just an FYI; the air/vacuum line going to the robot is starting to wear
    from rubbing on the flex track. 
5.279tdco on lineSTRATA::WOOLDRIDGEPleasure, Spiked With PainTue Oct 15 1996 16:514
    There is a TDCO on line which covers an added step for rinse time to
    help with particle problems.  Assuming this works out ok I'll spec it.
    
    /Kyle
5.280WAFER CRASHLUDWIG::WESTPHALWed Oct 16 1996 20:215
CMP.A1 - down for losing a wafer.  Cleaned up the process area and changed the
	 pad and carriers.  Left the same conditioner on but cleaned it.  
	 production filled out the crash sheet.  Tool turned back to VERIFY-U.

5.281particlesSTRATA::RIDLONEliminate the obliviousTue Oct 22 1996 06:2411
    AFter Ashlands work was completed tool failed particles X2 again.
    Particle map showed build up around edge. Installed point of use
    filters on both lines but tool crashed during breakins (fresh 50's).
    Slurry was checked after filters were installed and was good. Slurry
    flow was checked after it crashed (before halting) and there seemed to
    be plenty of flow. Possible air bubbles coming from filter caused
    crash. We purged them with slurry until air bubbles stopped or so we
    thought. New pad and carriers installed. Pad has been conditioned and
    awaits some fresh minds.
    
    					B-shiftoast
5.282pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Oct 25 1996 00:464
    
           Replaced pad due to bubbles....
    
                              D shift
5.283Bubble under padSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Oct 26 1996 04:486

           Tool went to unscheduled down for a bubble.  Pad replaced and 
       came back up.  MQC's passed fine...

                                          D shift
5.284Hub alignment performed on left spindleFABSIX::J_SWENSONWed Oct 30 1996 19:006
    performed hub alignment, run out .0003 gap .0056 - .0086, cleaned tool
    replaced pad and suba, put in verify and returned system back to
    production for mqc's.
    
    
                             C-Shift
5.285Slurry delivery and bubblesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 01 1996 05:5913

              The tool was given to us around 10:30pm for the
          slurry change over to barrel feed.  The tool was 
          converted over and slurry flow was adjusted in software.
  
              The 3-way slurry valves were replaced as well.

              When OPS went to use the tool, several bubbles had
          formed so we had to change the pad.  The tool is turned
          back of to OPS for MQC's.

                                          D shift
5.286Bubbles under padSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 08 1996 06:557


           Tool went to unscheduled DOWN for bubbles.  Pad
       was replaced and tool cleaned up and then was turned
       over to OPS for break-in.
                                          D shift
5.287Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Nov 15 1996 06:207

            Tool failed for deviation and range, replaced pad.



                                          D shift
5.288Wafer Breakage / Left Hub o-rings ShreddedASDG::POIRIERFri Nov 15 1996 19:1717
    
    Wafer fly out on left spindle:
    -----------------------------
    A wafer slid out of the left carrier and broke an oxide monitor
    this afternoon. Two of the four o-rings on the left spindle
    hub were shredded. This o-ring damage problem has been occuring
    on both Strasbaugh tools lately. 
    
    The only apparent reason for this condition is that the carriers
    must be slipping on the hubs. We need to understand why this is
    happening before we throw a carrier during a polish cycle. 
    
    Maybe the vacuum is low during the polish cycle. Possibly a vac leak
    in the rotory unions during spindle rotation.
    
                                                           C Shift.
    
5.289Wafer BreakageASDG::POIRIERSat Nov 16 1996 19:1511
    
    Wafer Breakage at 6:15 p.m.:
    ---------------------------
    The system broke two product wafers tonight. The pad & carriers have
    been removed. The tool still needs some additional cleaning and 
    replacement pad & carriers installed. We'll try to finish the
    work by the end of our shift.  Current status is bgn_repair.
    
    
                                                    C Shift.
    
5.290carriersSTRATA::RIDLONEliminate the obliviousMon Nov 18 1996 20:2616
    New pad and carriers installed. The right carrier wasn't pulling enough
    vacuum when picking up wafers from load cup. Had to replace the check
    valves (both were bad) and the black oring and the plate screws. 
    NOTE: I hope this is common sense but if not...If the oring on the
    plate screw is shredded or crushed than it wont make a good seal when
    tightened (No matter how tight). It would be nice to be able to replace
    these orings but I've checked stock and we don't have the right size.
    So please don't reuse plate screws that have deteriorated orings.
    Same goes for the black plate oring, if its been crimped or crushed in
    spots it wont seal correctly. If you run it between your fingers you
    can easily feel these bad spots. If you do throw it away. 
    
    
    						Mitch
    
    
5.291uniformity problemFABSIX::J_SWENSONSat Nov 23 1996 18:289
    
   -  Tool failed for uniformity
    
   -  Cleaned tool and replaced pad
    
   -  Returned tool to production
    
    
                          C-shift  
5.292Massive wafer breakageSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Nov 30 1996 06:128
           Came into a serious wafer breakage of two monitor
       wafers.  Tool needed an extensive cleaning.  After MQC's
       were run, they failed for scratches.  The crash implanted
       unforseen particles into the brushes.  After the brushes
       were replaced, particles came in.

                                          D shift
5.293wafer crashFABSIX::O_WHITFORDQuticherwhinin&#039;Wed Dec 04 1996 21:345
         Product wafer crashed. Cleaned polish table and table area,
    replaced carriers each with 57 wafers and pad with 114. Rewturned to
    ops to complete verification.
    
                                D Shift 
5.294Wafer crashSTRATA::WESTPHALFri Dec 06 1996 19:295
    CMP.A1 - Wafer crashed during an experimental lot.  The crash appears
    to be caused by a clogged slurry tube.  The same was reamed out, the
    process area cleaned and the consumables changed.  The tool presently
    sets needing breakins and MQCs.
    
5.295Amperage spikes during downforce ramp / Wafer fly outsASDG::POIRIERSat Dec 07 1996 19:4726
    
    Wafer fly outs on left spindle:
    ------------------------------
    The system has been flying out wafers on the left spindle.(no breakage)
    The operator noticed a loud vibration and an amperage spike (> 40 amps) 
    just prior to the flyout. System went down at 7:03 p.m., passing on to 
    B shift.
      
             
    Troubleshooting thus far:
    ------------------------
    1) Checked slurry flow on both pumps. Both at 450 ml/min.
       A little low but not significant.
    
    2) Checked pad for bubbles. Pad O.K.
    
    3) Checked carriers for proper gimbaling.  Carriers O.K.
    
    4) WE JUST OBSERVED CMP.A2 WITH 40 AMP SPIKES DURING DOWNFORCE RAMP!
       The last time we had downforce amperage spikes on BOTH tools,
       it was a problem with the PADS.
    
       Passing troubleshooting onto B Shift.
    
                                                                C Shift. 
    
5.296padFABSIX::M_RIDLONEliminate the obliviousSun Dec 08 1996 06:1313
    Picking up where C left off....we checked the carriers and found
    nothing unusual except for evidence of wafer fly out marks on the left
    carrier; this was replaced. Since all the pads in the room have gone
    well past the manufactures recommended "Use by this date" date, we
    ordered some new ones. Had to dig through the ones in stock to find
    some that hadn't expired. Can someone explain why RODEL puts expiration
    dates on these pads? We've been experiencing alot of crashes lately and
    the pads that were in the room all expired in OCT of 96. Could this be
    our root cause? Anyhow a new left carrier and an unexpired pad were
    installed. 
    
    
    						B-shift
5.297slurry detection installedSTRATA::FINNTue Dec 10 1996 19:306
	Installed slurry detection on CMP.A1. Wired into digital i/o 29.
	Sensor monitors main slurry loop an will alarm during the absence of 
        slurry when slurry pump #1 is demanded. 

						A shift
5.298Low bridge oil alarmSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Dec 11 1996 07:509

         Tool wouldn't start its cycle due to a low bridge oil alarm.
      Found bridge oil level with 1/2" left on the bottom.  Filled
      tank back up with oil.  All set....


 
                                          D shift
5.299POOR UNIFORMITYFABSIX::B_WESTPHALWed Dec 11 1996 19:225
CMP.A1 - Down For poor uniformity.  The pad and conditioner disc were changed 
	 out.  The chiller was found to be at 3 degrees C.  The chiller needed
	 to be topped off.  Tool was cleaned up and returned to production.

5.300Carrier changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 13 1996 05:048

             The right carrier failed for uniformity...  16.5%
         After the carrier was replaced, the tool passed MQC's.


 
                                          D shift
5.301Uniformity issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 13 1996 06:5210

             The right carrier failed for uniformity...  16.5%
         After the carrier was replaced, the tool failed uniformity
         on both heads so the pad was replaced.  MQC's were re-run and
         tool passed MQC's.


 
                                          D shift
5.302Poor UniformityFABSIX::B_WESTPHALSat Dec 14 1996 09:396
    CMP.A1 - Down for poor uniformity.  Due to the short pad life and table
    high currenting I think it must be the pads.  Changed the pad to a 
    lot# different from the previous pad used. Checked and cleaned the
    conditioning disc and did an extended conditioning ( 15 cycles ).  All
    so requested that the operator condition the pad 5 times between boats.
    
5.303Broken Wafers - Pads SuspectASDG::POIRIERSat Dec 14 1996 19:2514
    
    Broken product wafers:
    ---------------------
    The system broke product wafers this afternoon. A new batch of pads
    were brought down from the stock room. The new pads have a different
    lot number than the pad which the wafers broke on.
    
    The new pads are in the black cabinet. The suspected "bad" pads are on
    top of the black cabinet with a note attached. There are approx 200
    pads in the stockroom with approx 15 different lot #'s. Keep trying
    different lot #'s if the "bad Pad" symtoms persist.
    
                                                              C Shift.
    
5.304carrier bluesFABSIX::M_RIDLONEliminate the obliviousSun Dec 15 1996 04:538
    Right carrier lagging behind the left on lift off, also vacuum was a
    little higher than the left. Tore apart the carrier (the right one, for
    those with dislexia) and sure enough we found one of the plate screws
    loose. Replaced the three of them (orings shot) and also replaced the
    black oring which had a cut in it. Reinstalled and tested fine.
    
    
    					B-shift 
5.305more carrier bluesFABSIX::M_RIDLONEliminate the obliviousMon Dec 16 1996 03:498
    The left carrier was intermittently showing higher vacuum then the
    right and would alarm for no wafer found on left carrier. Removed and
    replaced the plate screws, the black oring, and check valves. The check
    valve hoses might have been too short causing the lines to crimp but
    thats speculative. Tested with good vacuum (better than the right now
    but only by .1 or .2 inches of Hg).
    
    						B-shift
5.306mqc failFABSIX::M_RIDLONEliminate the obliviousWed Dec 18 1996 06:4211
    Failed mqc's for uniformity. Pad at 293, carriers at 320. Changed both.
    Checked new carriers and found the one with plate A016 (no initials and
    no step height info) to built wrong. This must have been a test because
    even a new guy could do better. Check valves were facing each other,
    black oring was old, no carrier spacer washers installed and the screws
    were probably used before. I was surprised to see that the height was
    set up right. Anyhow it was rebuilt and installed. Over to production
    for breakins.
    
    
    				B-shift
5.307robot, strange acting.FABSIX::L_SHKOLNIKOVThu Dec 19 1996 06:3313
    

            The robot went in to pick up a wafer from right the send
        elevator and the arm went all way down hit the wafer guide and
        bend the wafer itself. No damage has happened at this time.
        I've tried to rehome the tool - no luck. Had to powered tool
        down and reseated all electrical connections on the robot 
        and on the back of the tool. Powered tool back up and after
        ten wafers notouch tool was running a product lot with no problem.
        
    

                                          D shift
5.308Failed MQC's - Monthly PM completedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 20 1996 06:237
        Tool failed MQC's for bad uniformity.  We decided to take
     it for the MONTHLY PM.  The downforce was out slightly and
     was broughtin tighter.  All of the run-outs were in spec.  
     Tool was turned back over for MQC's.
 
                                          D shift
5.309water leakFABSIX::J_SWENSONFri Dec 20 1996 19:315
    sprayer hose severed
    tool powered down
    di shut off.
    
                     C-shift
5.310Broken DI gun/lineSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Dec 20 1996 20:599



          Repaired broken DI spray gun.  Re-routed direction of line.
       Cleaned up DI under floor.

 
                                          D shift
5.311robot failure.FABSIX::L_SHKOLNIKOVSat Dec 21 1996 06:466
                The tool went down for robot problems. It wouldn't operate
        properly, (robot arm wouldn't move, initialize) Decided to remove
        the robot from the tool and check the encoders and wiring for
        rotational and radial axis.
        
                                    Leon
5.312robot/vacuum lineFABSIX::J_SWENSONSat Dec 21 1996 19:1815
    
      - Removed robot
      - Removed broken elbow for vacuum line from robot 
      - reinstalled robot
      - connected robot up, robot homed without issue
      - replaced vacuum line
      - after replaceing vacuum line and assembling flex tracked
        went to hook up robot, did not leave enough length for bottom
        connection flex track will need to be reopened.
      - a new elbow fitting will be needed.
      - di water is turned off
    
    
    
                           C-Shift
5.313robot readyFABSIX::M_RIDLONEliminate the obliviousSun Dec 22 1996 02:138
    Reinstalled air lines and electrical connections. Reinstalled buckets.
    Ran some wafers through to check robot alignments, made some minor
    adjustments but overall everything looked real good. Installed new pad
    and old carriers and gave back to production for breakins.
    
    
    
    					B-shift
5.314Right carrier / Robot FABSIX::R_GEEMon Dec 23 1996 18:5014
	The right carrier kept leaving wafers on the pad. The carrier
	backing plate screws and o-rings were changed but still was
	leaving wafers on the pad. Changed out the right carrier and
	the problem went away.

	The robot kept missing wafers. Cleaned the robot end effector,
	replaced the 1/8" end effector line & leveled the robot top
	arm. After this the robot top arm failed to home, powered
	down the system to reset the computer. Ran 25 verification
	wafers then released back to production, no errors the rest
	of our shift. (and yes they did run it). 

					A shift
					
5.315No Z'ssss for the robotSUBPAC::LANDRYTue Dec 24 1996 13:505
    The "Z" motion on the robot froze up today.  Re-homed the robot (all
    3 axis') and tested.  Ran fine the remainder of the shift.
    
                                        A shift
    
5.316Robot Errors / Robot top Arm Sensor CleanedASDG::POIRIERFri Dec 27 1996 19:3416
    
    Robot errors:
    ------------
    Robot Z axis travel was not moving. This cause robot errors. We have
    not been able to recreate this condition during testing. 
    
    We did observe robot rotation chatter. Strasbaugh was called and 
    instructed us to check inputs 41, 42 & 43. All checked aok.
    The robot top arm home sensor was covered with slurry. It was cleaned
    and all robot functions appear to be ok. A verification run has been
    started and the system status is now in VERIFY. 
    
    The Strasbaugh Field service reps pager is 1-800-506-4960 if you
    require tech support tonight. (Todd Fordyce)
    
                                                           C Shift.
5.317Robot failed... again...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Dec 28 1996 07:1820


            The tool was down at the beginning of the shift
         for VERIFY on the robot.  The tool ran ok and passed
         MQC's.   Around ~4am, the robot crapped out for Z-axis
         positioning errors.  The robot was pulled and cables
         inspected.  Some moisture was found in the connectors
         and three broken wire were broke which went to the 
         rotation encoder?????   The wires were repaired and
         nothing else was found wrong.  The robot and right
         receive elevator are currently being re-installed.  
         We will pass-on to C shift regarding what's left to
         do for the re-install. At worst, the robot and elevator 
         positioning will need to be checked as well as the 
         vacuum cal.
  

 
                                          D shift
5.318Robot Z Axis Failures / Encoder Cable ReplacementASDG::POIRIERSat Dec 28 1996 19:2044
    
    
    
    
    
    
    
    
    
    Robot Encoder Cable Needs Replacement:
    -------------------------------------
    The Robot Z axis has been intermittently (Failing/ running into end
    stops) for the past two days. 50 + wafer verification runs have been
    performed numerous times. Eventually the robot Z axis runs away. 
    Sometimes in the upward direction and sometimes in the downward
    direction. 
    
    Action thus far:
    
    1) A known good DCX control module was swapped with axis 12 (robot Z).
       Axis 12 failed intermittently a short time later. 
    
    2) The encoder cable connections for the Z axis at the Robot were
       verified AOK.  pins 19,20,21,22,8,9 &10.
    
    3) Jim Tooker was called at home this afternoon and agreed that the 
       cause of the intermittent Z axis encoder failures is most likely 
       Robot Encoder Cable.
    
       Jim doesn't think the problem is being caused by the z axis motor
       encoder due to the intermittent behavior of the failures.
       He wants the Robot Encoder Cable removed and will order a
       replacement asap. Call Jim at home if your have any issues
       that arise relating to the cable over the weekend.
    
       The cable has been snaked out thru the electrical cabinet.
       It is still wired into the ribbon interface cards on the 
       electrical cabinet end. 
    
       Inspection of the cable revealed (knots/twisting/bunching)
       of the conductors in numerous places along the cable. 
    
    
                                                      C Shift.
5.319encoder cable fixed but...FABSIX::M_RIDLONEliminate the obliviousSun Dec 29 1996 06:5514
    Ohmed out the encoder cable removed by C-shift. Found two wires that
    were open. One goes to the z axis and the other to the theta axis.
    Cut almost 2 1/2 feet off the cable and reohmed the cable and all wires
    ohmed good. Snaked cable back thru machine, repinned, reconnected all
    cables and tried powering up the tool to test. However the cpu would not
    reboot. Removed the cpu cover but didn't find any discrepancies.
    Checked the power supplies and 5vdc and 24vdc were good the 15vdc was
    not, however this is controlled by CR28 which is controlled by DI 15.
    None of the DI's are lit so this is probably due to the cpu not being
    booted up. If the problem the cpu can be corrected the flex track for
    the robot still needs to be put back together. 
    
    
    						B-shift
5.320Computer / Robot top armFABSIX::R_GEEMon Dec 30 1996 19:418
	Field Service was in today to investigate the boot-up
	problem. Field service feels the computer is toast. A
	new computer will be arriving thursday.
	
	Replaced the robot top arm belt gears and tensioners.

					A shift
					
5.321status of cmp.a1SUBPAC::TOOKERThu Jan 02 1997 18:4323
The new computer arrived today for cmp.a1. To save time, the hard disk 
drive from the old computer was installed in the new computer and the DCX 
motherboards were installed before Strasbaugh field service arrived.  When
field service arrived they installed the computer.  The new computer booted
up just fine, but they could not get any video on either the front or remote
monitors and there was a keyboard error.  Strasbaugh found that the main 
"black box" was not working and the keyboard cable adapter had a broken pin.
They feel that when the black box went out, it took the computer with it. 
A new "black box" and adapter have been ordered.

Currently, we have made the tool operational by using the keyboard and monitor
from the PC we have in the service aisle.  Now that the tool will boot up, we
can now verify that the fix on the cable performed by D shift on 12/28 is OK.
John Poirier will install carriers and run at least 100 wafers with the
Handling recipe.  This will stress the cable in the flex track.  If all is ok,
he will try to set up the temporary monitor and keyboard in the WIP aisle,
install a pad, and try to qualify the tool.

Please call me home with any questions.

Jim

              
5.322Computer Replaced / video control boxes NGASDG::POIRIERThu Jan 02 1997 19:3116
    
    COMPUTER REPLACED / CURRENT STATUS:
    ----------------------------------
    The Strasbaugh Reps sucessfully replaced the computer today.
    They discovered that the video controller boxes for the monitors
    are not functioning (Both Monitors Blacked out). There are three
    of these boxes in the tool. Replacements have been ordered and 
    will arrive tomorrow.
    
    We have completed 75 wafers of the 100 wafer verification run. Please
    run another 25 on D Shift.
    
    
                                                             C Shift.
    
    
5.323Shuttle doin' the see-saw...SUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Jan 03 1997 05:0014

            The tool finished running the 100 VERIFY wafers
        and all of the consumables were put back on.  After the
        pad was broken-in and MQC's were getting ready to run,
        OPS observed carrier damage.  The white rings were both
        being shaved.  Upon further investigation, it was
        discovered that the front shuttle clamp-down brackets
        were under the shuttle instead of on top.  The shuttle
        was rocking back and forth as a result of it.  Re-positioned
        the brackets and installed new carriers.

 
                                          D shift
5.324Video Controller Box Back Ordered / Temporary Set up in Place.ASDG::POIRIERFri Jan 03 1997 19:3711
    
    Monitor & Keyboard Temporarily Connected:
    ----------------------------------------
    The Dell Monitor & keyboard were placed on the WIP isle clean room 
    rack next to A1. The wires and cables were snaked under the wall and
    connected to the IBM computer in the tool. This temporary set up will 
    remain in place until the back ordered video controller box arrives
    on Monday.
    
                                                       C Shift.
    
5.325Pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Jan 04 1997 04:257


              Pad change because of bubbles....

 
                                          D shift
5.326Front monitor is working againFABSIX::R_GEEMon Jan 06 1997 16:475
	Replaced the bad KVM Extender/Sharer host module (black box).
	The Dell monitor has been removed and the original monitor is
	working again. 
					A shift
					
5.327pad changeSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringWed Jan 08 1997 07:134
    A pad change was needed due to bad uniformity.
 
                                          D shift
5.328pad change - bubblesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringThu Jan 09 1997 06:576

    A pad change was needed due to bad uniformity.
 

                                          D shift
5.329Vacumm leaks at the hubs / o-rings replacedASDG::POIRIERFri Jan 10 1997 18:5822
    
    System leaving wafers on pad:
    ----------------------------
    The tool was leaving wafers on the pad at lift off (right Spindle).
    The two center o-rings on the right hub were replaced. Tested O.K.
    
    
    
    Left Spindle not lifting off pad:
    --------------------------------
    Earlier in the day, the left spindle would take much longer to lift off
    the pad than the right. It was taking it longer to reach the approriate
    vacuum. The center two o-rings on thge left hub were replaced. Tested
    O.K.
    
    FYI: The center two o-rings on both hubs were shredded. We believe this
         was causing the vacuum leaks. We called Strasbaugh a few weeks ago 
         about the o-ring shreding problem. They think the carriers are
         slipping on the hubs. 
    
    
                                                            C Shift. 
5.330tanks filling problemFABSIX::J_SWENSONSat Jan 11 1997 18:1512
    
       Tanks not filling quickly
    _________________________________
    
     Di pressure 37
    
     cda 100 psi
    
     reset tool, tanks filled still slow 
     di line acting as if there is air within the line.
    
                                  C-shift
5.331Problems with the REMOTE/Black BoxSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringTue Jan 14 1997 16:4111



          Since the video repair from the Strasbaugh Rep., it appears
      there is a problem when you have REMOTE unit plugged in and you
      switch the controls back to the front.  You have to have the REMOTE
      unplugged to have the front work.  Jim Tooker was notified today 
      and he is looking into it.

                                 BJ
5.332mqc failedFABSIX::M_RIDLONEliminate the obliviousSun Jan 19 1997 07:446
    Failed mqc's late in the shift. Pulled pad and started cleaning. No
    work for A tools in room so leaving down so A-shift can finish up the
    weekly pm. 
    
    
    						B-shift
5.333replaced clean station brushFABSIX::J_SADINFreedom isn&#039;t free.Mon Jan 20 1997 18:005
    
    
    	replaced left clean station brush. Outer edge was worn.
    
    a-shift
5.334CMP.A1 chillerFABSIX::J_SADINFreedom isn&#039;t free.Sun Jan 26 1997 16:149
    
    
    	The chiller for CMP.A1 was found powered off. Hotline call was
    logged and electrician found breaker tripped in circuit panel.
    Electrician reset breaker and power was restored. Electrician checked
    lines/connections and found nothing wrong (no idea why breaker
    tripped). 
    
    	a-shift
5.335Carrier sticking to padFABSIX::O_WHITFORDLead, follow or step aside.Fri Jan 31 1997 04:067
         Left carrier sticking to pad at lift-off. Swapped left to
    right & vice versa. Problem followed carrier. Went to replace carrier
    and noted both were high on wafer count. (left at 359, right at 358) 
    Only one carrier available. Built up another carrier and replaced both. 
    Turned over to ops for full MQC.
    
                             D shift
5.336Left Spindle Rotory Union Leaking Vacuum.ASDG::POIRIERSat Feb 08 1997 19:2427
    
    Left Spindle Vacuum Failure:
    ---------------------------
    The Left Carrier flew off the spindle while polishing product today.
    Luckily, the product wafers did not get damaged. One of the left Hub 
    o-rings was shredded. A vacuum gage was installed between the carrier
    and the hub on both spindles. Vacuum on the Right was 24 in/hg. We applied
    moderate pressure to the Right Spindle Rotory Union and the vac remained
    at a steady 24 in/hg.
    
    Vacuum on the left spindle rotory union was 19 in/hg. We applied
    extreamly light pressure to the union and the vacuum dropped to 10
    in/hg. We then squirted some SNOOP liquid leak detector to the top
    of the union and the vacuum increased to 24 in/hg. The solution was
    wiped off and the vacuum dropped to 19 in/hg.
    
    It appears that the low vacuum induced by the leaking rotory union 
    is causing the carriers to slip on the hubs. This in turn is 
    shredding the silicone rubber o-rings in the hubs. Once the hub
    o-rings fail, the vacuum drops to very low levels and the carriers
    fly off.
    
    The system is down. Passing on to B Shift for left spindle rotory 
    union replacement.
    
                                                                C Shift.  
                                                                     
5.337Gear box, rotary union, system power problemSUBPAC::LANDRYMon Feb 10 1997 19:5139
    The system is down (scheduled) for an upgrade to the polish table gear
    box.  In fact, it's a new box.  The system is also down for a rotary
    union on the left spindle.  This is due in Tuesday morning.  Also a
    problem right now, is the system is powering itself down.  More on
    this at the end of the note.
    
    Gear Box:  Two East Coast reps and Hallamore riggers performed the swap
    today.  The new one is in.  This action required the removal of the
    polish table as well.  All components associated were re-installed
    and all alignments associated were performed.  The gear box portion is
    complete.  Fixed several damaged sprayers along the way.  Cleaned out
    built up slurry in places unreachable with the table in place.  
    
    Rotary Union:  It has been ordered and should arrive Tuesday AM.  "A"
    shift will install it and perform all the spindle alignments.  Also,
    the monthly portion of the PM will be completed at this time.  After
    all these things are completed, the system will be tested for proper
    operation.  Then, additional breakins and MQC's will be run.
    
    System Power Down:  After the gear box was installed, the system was
    powered up and checks were performed.  A couple of hours into power
    up, the system powered down.  It rebooted fine, but 10 minutes later,
    it powered down again.  The "red" power light on the back of the
    system was not illuminated.  Wiggled some components around, and it was
    back on.  It was first believed to have been contactor #1, as that
    component was being played with when the system powered down.  The
    contactor has been ordered.  We then were going to remove the contactor
    and clean it, when the red light on the back of the system was found
    off again.  This light should be on when the power is making it through
    the first circuit breaker, which it was, but no light.  Traced this
    problem to the 24 VAC relay (which is associated with the interlocks).
    Non in stockroom 6.  Check SR 94 (going through all the relay bins)
    without success.  We cleaned up the old relay's contacts, verified good
    connection of the wiring at the relay plug-in spot, and retested the
    results.  We couldn't get the relay to fail.  The system has remained
    powered up for about 40 - 45 minutes without any problems.  Please
    monitor it occassionally this evening for a power up state.
    
                                           A shift
5.338new rotary union/ table problemFABBIT::B_FINNTue Feb 11 1997 19:3817

	Removed left rotary union per c shift. Installed new rotary 
union which we received today fedx. Installed hub, and continued 
monthly pm (ie spindle alignments) Found left spindle out 10 mils found 
right spindle out 9. Noticed both were low on the same side of the 
table. Did a sweep of the table with via the bridge with dial indicator 
attached to the spindle. Found table to be out left to right by 22 
mils. This is a result of the table being removed yesterday during the 
gear box change. Spoke to Jim T. at home. He wants the rep to fix it 
tomorrow. He also wanted to communicate that the rep will be assigned 
first to either CMP.A1 or A2 depending on productions needs in the morning.
 
We also changed the main contactor as a precautionary measure.
					
							A Shift

5.339CMP.A1 alignments doneFABSIX::J_SADINFreedom isn&#039;t free.Wed Feb 12 1997 19:3210

        Rep was in today to address table alignment issue. He calibrated the
table sweep, as well as calibrated all runouts (table, spindle,
hub, etc). Rep also fixed a conditioner rotation problem by tweaking the
DCX board and then tweaking the BRUSH ROTATION board. Downforce cal was 
completed by us. New pad/conditioner installed and breakins completed with
no furthur issues. Mqcs are in process now and tool is back to production.

        a-shift
5.340Bad unif., monthly completedSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringFri Feb 14 1997 05:168


          Tool failed MQC's for 20%+ uniformity.  Did a 
     pad change and finished outstanding monthly PM item.
     Tool was broke-in and passed MQC's.

                                       D shift
5.341Tool EMO'ing Itself / Table RPM Low ErrorsASDG::POIRIERFri Feb 14 1997 19:2525
    
    Table RPM Low Errors:
    --------------------
    The tool errored 3 times this afternoon for "Table RPM Low" errors.
    The Strasbaugh rep adjusted the Table Motor Controller DIP switches per
    the latest rev Strasbaugh print. We're not sure why the settings
    differed from the print. Maybe it's because the 7.5 to 1 ratio gearbox
    requires different DIP switch settings than the new 15 to1 ratio box.
    The Strasbaugh prints that we have are the latest rev for the 6DS-SP
    polisher.
    
    While running verification wafers, the tool EMO'd itself and stopped
    the run. The last time we encounterd this problem, it was the EMO
    button on the WIP isle side of the tool. We didn't have a replacement
    switch so we disassembled the switch and streched the spring a little.
    The problem did not reoccur for over a year. 
    
    Suggested plan for tonight:
    
    1) Inspect EMO switches and repair.
    
    2) Continue with verification run to ensure Table Controller DIP switch
       change solves the "Table RPM Low" errors.
            
                                                          C Shift.
5.342EMO and stalling issuesSUBPAC::BJUBINVILLECMP-6 Equipment EngineeringSat Feb 15 1997 06:5823
    
         well...  what a night...
    
             Tool EMO'ing before, during and after every two
         wafers being polished.  We did massive trouble-shooting
         with the main breaker box and EMO/interlock circuits.
         We kept having intermitting 24VAC problems.  We by-passed
         the EMO buttons and it kept EMO'ing.  Everything was checked
         for any possible loose wiring.  The #3 in-line 208 line had
         some slop on the terminal block below the main breaker.  We
         noticed some intermitting arching.  We by-passed the 
         interlocks and it worked fine.  While the tool was running
         wafers, we checked all of the interlock wiring which looked
         ok.  We removed the jumper and tried running wafers with 
         everything back to normal.  we ran 24 wafers total from the 
         remote controller and 8 wafers with the front controller.  
         The tool ran fine without any table amperage or stalling
         issues nor any EMO problems.  When all was said and done,
         MQC's are about to be run.
 

                                                D shift
                                   
5.343Left spindle rotary jointFABSIX::R_GEEMon Feb 17 1997 19:338
	The system errored for left spindle not rotating. Found the top
	part of the left spindle rotary joint seized, (the bearing retainer
	ring was ripped out and the bearing were ready to fall out).
	Replaced the rotary joint, re-aligned the hub, replaced the pad &
	conditioner disk, ran verification wafers. 

					A shift
					
5.344Pad changeFABSIX::B_JUBINVILLEFri Feb 21 1997 08:354
    
        Pad was replaced due to failing MQC's for unif.
    
                                  BJ
5.345Faulty Door Interlock Switch / Robot Rotation Homing FailureASDG::POIRIERSat Feb 22 1997 19:3726
    
    Polish Door Interlock Switch:
    ----------------------------
    The system failed to home all axes due to a "Polish Door Open " error
    message on the CRT. The right hand interlock switch was found to be BAD. 
    A replacement switch is sitting on the polish area deck. Please
    replace.
    
    
    
    Robot Rotation Axis Homing Failure:
    ----------------------------------
    The System homed every axis except the Robot Rotation Axis (#11).
    The robot initiates the rotation homing sequence, then continues on
    past the home position until it contacts the end stop. 
    
    The robot home sensor connections were visually inspected at cable
    side and robot side of the miltary style connector. The conductors,
    sockets and pins all look good. 
    
    The robot has been removed from the tool and is on the table in the
    service chase. We're in the process of replacing the robot rotation
    home sensor. It is labled #2.
     
    
                                                               C Shift.
5.346switches and cables and bears, oh my.FABSIX::M_RIDLONEliminate the obliviousSun Feb 23 1997 02:4017
    Continuing on from where C-shift left off. We ohmed out the right hand
    interlock switch and found nothing wrong. The contactor switched when
    the lever arm was pressed. It was wired back up and tested fine. 
    The robot rotation home sensor was replaced and tested and the robot
    still would not home rotationally. Dug out the technicians first line
    of defense (DVM) and started checking for signals. Relay CR24, which is
    energized by the robots theta home sensor, was being energized but it
    was being deenergized. The voltage was only dropping to 9 volts (from
    23) instead of to zero. Pulled the connector off of 13r1 junction block
    and found that pin 10 had shorted to 23 ( or vica versa depending on
    the theory you follow ). Removed the connector housing and inside was
    corroded and soaked. Called Jim and he just happened to have an extra
    cable in his cabinet (oh joy). We're replacing now and will test and
    adjust what ever needs it. If it works you won't here anything more. 
    
    
    						B-shift
5.347CMP.A1 conditioner arm crashFABSIX::J_SADINFreedom isn&#039;t free.Tue Feb 25 1997 20:0513
    
    
    	received a call from a-shift operator that the conditioner arm had
    crashed into the slurry distribution tube. Hence, the slurry tube was
    bent and the conditioner arm would not respond at all. Straightened the
    slurry tube and then found that one of the brushes had fallen out of the
    conditioner motor (the plastic brush holder was cracked/broken). 
    Epoxied/taped the brush holder in place. Reinstalled the motor and
    recalibrated the home sensor. Ran a couple of test conditioning runs
    and all seems to be ok. Back over to production.
    
    	a-shift
     
5.348park itFABSIX::M_RIDLONEliminate the obliviousWed Feb 26 1997 01:3610
    Pad conditioner would not park. We ended up removing the motor/piston
    assembly from the conditioner. Then removed the motor from the piston
    assembly. And then we retaped the brush. Once everything was
    reinstalled we were able to calibrate the park position sensor with
    ease. Later in the shift the tool failed mqc's twice so the pad was
    changed.
    
    
    							b-shift
    
5.349Conditioner sweep motor.FABSIX::L_SHKOLNIKOVFri Feb 28 1997 06:299
       Pad conditioner electric cylinder (arm sweep motor) has been removed 
    from the tool and replaced with a new one. The replacement motor had a
    very short encoder cable which was replaced with the old one. Home
    position sensor - recalibrated. We noticed that conditioner goes too far
    out over the table edge, so outer offset has been changed from 98500 to
    96600. Tested - OK. Returned to production.
    
                                                    D-shift.
    
5.350Dorr Latches Need ReplacementASDG::POIRIERSat Mar 01 1997 19:0411
    
    Door Latches:
    ------------
    The WIP isle side doors need replacement latches.  The top latches on
    both CMP.A1 & CMP.A2 have been replaced with their respective bottom
    latches. (Sensors next to top latch) The doors are working ok for
    now, but we'll be in trouble the next time a top latch fails.
    We can order replacement latches from Strasbaugh or find another
    type that will work.
    
                                                           C Shift.
5.351A dark and stormy nightFABSIX::O_WHITFORDLead, follow or step aside.Thu Mar 06 1997 23:1813
    Tool went down for pad change.
    
       Replace pad, condition positioning off. Noticed with conditioner
    homed and commanded down, it struck the edge of the bucket. Made
    mechanical adjustment to turnbuckle at conditioner end of arm oscillate
    cyl. Homes good, centers up in bucket good, conditione to edge of pad.
    
       Small leak in polish table area. Table rinse bulkhead fitting loose.
    Tightened, no more leak
    
       Right carrier wafer count very high (348) Replaced right carrier.
    
                                       Dave & Leon  
5.352Carriers Sticking To PadASDG::POIRIERSat Mar 08 1997 17:0435
    
    Carriers Not Lifting Off Pad:
    ----------------------------
    The Carriers would not lift off the pad during an auto cycle. The pad
    had only 40 wafers on it and the conditioning disk was changed during
    the weekly pm on Wednesday the 5th. 
    
    We pulled up workstream and found the lot #'s of the previous 3 pad
    changes on both CMP.A1 & CMP.A2. We made sure we chose a different 
    lot # pad due to the "long breaking periods", "short pad life", "quick
    drop-off in uniformity" and "carrier sticking to pad problems" that
    have plauged us all week. 
    
    After the Pad & Conditioning Disk change, we ran 6 break-in wafers
    using the "newpad" recipe. The carrier to pad STICKING problems 
    ceased to exist.
    
    
    The current pad lot # is 6HR-JP2. Please monitor its performance.
    
    Here is a little pad change history from this week:
    
               CMP.A1                                       CMP.A2
    ------------------------------              --------------------------------
    Pad Lot #             Problem                Pad Lot #             Problem
    ---------             -------                ---------             --------
    6FL-JP4 /40 wafers   Carriers sticking       6FE-JP4 /180 waf.  no prob, PM
    
    6HS-JP1 /198 wafers  Uniformity              6HS-JP1 /318 waf.  Unifor & RR
    
    6HS-JP1 /232 wafers  Uniformity              6PE-JP4 /168 waf.  Uniformi
     
    
    
                                                                   C Shift.
5.353Conditioner Head Rotation FailuresASDG::POIRIERSat Mar 15 1997 10:4838
    
    Conditioner Rotation Failures:
    -----------------------------
    The conditioner has been intermittently failing to rotate during the
    its sweep while running the NEWPAD recipe. (higher rotation speed than
    the other recipes) It has been failing between one and four times per
    sweep. Power to the head rotation motor has been interupted during each
    rotation failure.
    
    The yellow LED, (CL) (current limit) on the rotation motor controller board
    illuminates every time the rotation stops. The CL pot was turned 1/8
    turn clockwise to slightly increase the current to the rotation motor.
    This has stopped the conditioner head rotation failures but is not the
    root cause.
     
    The end cover was removed from the conditioner arm and the belt tension
    checked out o.k. But, a loud grinding noise was observed. The
    mechanical drawing for the conditioner assembly shows one ball bearing
    on the top and one on the bottom of the vertical shaft of the
    conditioner. There are also two sets of two ball bearings for the 
    belt tensioners inside the top arm. Its not clear which bearings are
    bad but one or more of them definitly are. Its most likely the top
    bearing on the vertical shaft assembly. Inspection will require major
    disassembly of the conditioner. 
    
    Strasbaugh was called for pricing and avilability of the six bearings.
    Since this happened late in the shift on Friday, they were instructed
    to reply to Jim Tookers voice mail with the ordering info. Our under-
    standing is that they're not available for parts ordering on Sat & Sun.   
    We'll call Jim T. at home this morning to verify this and he'll have
    follow up on Monday morning. 
    
    The conditioner is still working fine at the higher current setting but
    will eventually fail again as the bearing/s degrade further.
    
    
                                                                 C Shift.
    
5.354POOR UNIFORMITYFABSIX::B_WESTPHALWed Mar 19 1997 18:325
CMP.A1 - Down for failing MQCs for poor uniformity.  The process area was
	 cleaned.  The pad, conditioner and Left carrier were swapped out.



5.355Dropped right carrierFABSIX::O_WHITFORDLead, follow or step aside.Wed Mar 19 1997 22:337
    Right carrier fell off spindle with shuttle in motion going to the 
    "clean" (in) position. This damaged the carrier film and the carrier
    had a wafer count of 311... so replaced the carrier & inner most
    silicone o-ring. (o-ring was moderately frayed) Returned tool to ops in
    VERIFY-U.
    
                                       Dave
5.356Right Carrier Fell Off Hub / New Fairchild Vac RegulatorASDG::POIRIERFri Mar 21 1997 19:4019
    
    Right Carrier Fell Off Hub:
    --------------------------
    The Hub o-rings were inspected and found to be in good condition.
    Strasbaugh was called and informed us of a historical problem with
    the Fairchild vacuum regulators. We replaced the aspirator tubes in the 
    regulators with the new stainless steel type that had previously
    been sent to Jim T.  A 20 wafer verification run was sucessfully
    completed and the tool status was changed to UP. 
    
    The needle valve seal in one of the regulators appears to be slightly
    deformed. Jim has arrainged for a Saturday delivery of 2 new
    regulators. We'll install them tomorrow.
    
    
                                                            C Shift.
    
    
    
5.357Carriers are falling....SUBPAC::BJUBINVILLECMP Equipment EngineeringSat Mar 22 1997 05:568
              Within 20 minutes of each other, each carrier
           had fallen off the tool and into the load cups.
           Tool is down waiting for the Fairchilds as indicated
           by C shift in the note previous to this one.  They
           are due in today.

                                       D shift
5.358Left Spindle Dropping CarrierASDG::POIRIERSat Mar 22 1997 19:3223
    
    Left Carrier Falling Off Hub:
    ----------------------------
    Replaced faichild vacuum regulators. The carrier dropping problem 
    continued. We observed the left carrier dropping during the "spindle
    rinse" step of the auto cycle. (just after unload)
    
    The furon air operated isolation valve for the left spindle was
    emitting rust from its exhaust port. The valve was disassembled
    and the diaphram spring was rusty and brittle. It broke as soon
    as we touched it. We did not have a replacement valve in stock but 
    found the correct spring in one of the old 3 way slurry distribution valves
    that were removed for an upgrade.
    
    We ran numerous spindle flush cycles and the carrier vacuum remained
    above 23 in/hg. B SHift, please run 25 wafers using the HANDLING recipe
    as a verification. The system status has been changed to VERIFY-U.
    
    After the verification run, the tool will need a new pad & Carriers.
    
                                                                   C Shift.
    
    
5.359rotary union replacedSUBPAC::LANDRYMon Mar 24 1997 19:2921
Well, the rotary union on the left spindle went bad again.  Same thing
as the previous union, where the race was chewed up pretty bad.  This 
time, it was bad enough to cause a rotation axis error. Seeing this is
at least the 3rd union to go in the past month/month and a half, we 
thoroughly examined all physical aspects of the assembly, in hope of
pinpointing the cause of this dilemma.  No apparent reason was detected.
Todd Fordyce was contact for advice.  All the things he recommended to 
check were previously checked, with no problems found.

A new union was installed and it appears to be fine.  Hub alignment was
performed and is within proper tolerence.  A new pad is being installed
as I type.  The system will be broken in post pad install.  It will need
to have the carriers and brushes re-installed.

In the meantime, Strasbaugh will check further into the problem for us.
They have a couple of leads to chase.  They will get back to us as soon
as possible.

Also today, the weekly pm was performed.

                                          A shift
5.360Failed 24 hr MQC'SFABSIX::R_POIRIERWed Mar 26 1997 06:238
		A1 failed 24 Hr MQC's. Twice. It got worse the
		second time. Changed pad and conditioner.



						B Shift

5.361long pad breakinsFABSIX::J_SWENSONThu Mar 27 1997 19:3915
    long breakin for pad #6FZ-JP1
    -----------------------------
    
   -replaced pad today mqc's failed.
    
   - replaced pad conditioning disk also
    
   - the pad took 48 wafers to breaking before mqc's would pass
    
   - replaced carriers both had 414 wafers on them
    
              Looks like there is a pad problem on the rise
    
                                C-Shift
    
5.362Pad replacedSUBPAC::BJUBINVILLECMP Equipment EngineeringFri Mar 28 1997 05:218
            The carriers started sticking to the pad after
        the polishing was done.  The pad's adhesive wasn't
        holding down to the SUBA and came off the tool with
        two fingers and no resistance.  The pad was replaced
        and was broken in.  We are awaiting MQC's to be run.

                                       D shift
5.363left carrier stickingFABSIX::J_SADINFreedom isn&#039;t free.Tue Apr 01 1997 20:396
    
    
    	left carrier sticking to pad. Replaced plate-oring and plate
    screws. Ran test wafers ok. 
    
    a-shift
5.364Load Stations AdjustedASDG::POIRIERWed Apr 02 1997 20:289
    
    Load Stations Adjusted:
    ----------------------
    Both load stations were adjusted due to loading failures today.
    The thin flat washers on the load station brackets were replaced
    with a thicker type. The thin washers were distorting into a concave
    shape which made the adjustment move while the bolts were being tightened. 
    
                                                                   C Shift.
5.365rotory unionsLUDWIG::FINNSun Apr 06 1997 20:4818
Continued left rotory union installation from B shift. New rotory union is one 
of three spare recently received. Strasbaugh had felt that these premature 
failures were the result of our using their "rebuilt" rotory unions. This one 
installed today is believed to be new, not rebuilt. 

That said, we still choose to install the right rotory union on the left 
spindle, and install the new rotory union on the right spindle. Any further 
premature failures on the left spindle will be due to causes other than rebuilt
rotory unions .

Had wafer vacuum errors on left carrier. Found to be caused bad/old old 
tubing at hub, replaced.  Checked hub gaps and performed hub alignments.

Returned to manufacturing. System is currently MQC.


						A shift
5.366pad change due to high uniformityFABSIX::A_QUILLENNobody will get away with anythingTue Apr 08 1997 08:315
    Pad was changed due to high uniformity (over 20%). The tool was put to
    end repair without recording pad data. The numbers are as follows:
    #6ED-DA5
    #6GF-JP2
    MQC's passed.
5.366Pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 10 1997 07:324
    
       Pad change due to MQC's  - bad unif. 
    
                                  D shift
5.367Carrier dropFABSIX::O_WHITFORDLead, follow or step aside.Wed Apr 16 1997 00:1516
    Problems with the right hand carrier faling off spindle.
    
         1) Checked vacuum at hub and at spindle  24"Hg
    
         2) Checked vacuum at top of spindle (entry port
            into rotary union)                   30"
    
         3) Re connected vac lines after ensuring the mating
            surface of the ferrule & the fitting were clean
            and free of burrs, dents dings etc.
    
         4) Trimmed each end of the tygon tubing between hub
            & spindle. Ops ran some breaking wafers with no 
            problems, logged into Verify-U for them to run MQC
    
                                Dave
5.368Right carrier still falling off.LUDWIG::STANDINGWed Apr 16 1997 20:498
    CMP.A1  Continued with the problem of the right carrier falling off. 
    Also after about three to four wafers it would not be able to get
    sufficient vacuum on the wafer.  Upon doing a wafer blow off found
    water coming out of the chuck.  Removed chuck ring and found nothing
    wrong.  So I replaced the vacuum line and put tie wraps around both
    ends.  Hooked up a chart recorder to monitor the voltage on sols 16 an
    17 to see if they could be closing at the wrong time.  Could not run
    any more wafers due to the tool not seeing any send cassettes.
5.369Uniformity and vacuum issuesSUBPAC::BJUBINVILLECMP Equipment EngineeringThu Apr 17 1997 01:0316
    
        took over from C shift.  SEND elevators were ok.
     Ran 24 wafer verify for the carrier vacuum issue.  
     The right carrier was pulling poor vacuum at LOAD
     (13 psia) with a wafer and it wouldn't come down.
     Swapped carriers left to right and the vacuum was
     4 psia on both.  the tool ran fine with no carriers
     dropping off or any other vacuum problems.  Turned
     tool back over to OPS for MQC's.  Tool failed MQC's 
     for 16+% uniformity.  Checked both hub alignments 
     and found left hub at .004 and the right at .005.  
     Brought both hubs below .0005.  The pad had a bubble
     so that was replaced as well.  Tool ready for MQC's.
    
    
                                             D shift
5.370WAFER CRASHFABSIX::B_WESTPHALThu Apr 17 1997 20:327
CMP.A1- Down for a wafer crash.  Cleaned the tool up and replaced all
	consumables. Checked the carrier to load bucket alignments everything
	was good.  Noticed that the vacuum filter housing up in the bridge was 
	leaking.  The same was removed from the system.  The pad was
	conditioned and returned to production for MQCs.

5.371dropping right carrier againFABSIX::O_WHITFORDLead, follow or step aside.Fri Apr 18 1997 21:2816
    Blew off right carrier again.
    
    Aske ops tech to try two more wafers and observe the vacuum. Everything
    went fine until spindle flush. Lef spindle flushed at @ 40 psiA and
    right spindle flushed in excess of 52 psiA. Hmmm....
    
    PSS made a slight downtweak on the back pressure variable and the next
    two wafers unloaded, and spindles flushed, both at @ 40 psia and the
    right carrier remained with the hub. Ah-mazing.... :)
    
    No for the REAL test, run it for a couple days and see if the problem
    comes back....
    
                               D SHIFT
    
    
5.372PAD CHANGEFABSIX::B_WESTPHALFri Apr 18 1997 21:444
    
CMP.A2 - Down for low removal rates.  Changed the pad conditioner.  cleaned up 
	the tool and returned it to production.

5.373Failed MQC'sFABSIX::R_POIRIERMon Apr 28 1997 08:147
    		
    			Tool failed 24 hr MQC's. Changed pad
    			and conditioning disk. Passed MQC's.
    
    
    					B Shift
    
5.374New robot cable has been orderedSUBPAC::TOOKERTue Apr 29 1997 16:265
    A new cable has been ordered for the cmp.a1 robot.  This takes some
    time to change out, so it will have to be planned.  For the time
    being, there should be no urgent need to change the cable.
    
    Jim
5.375cable for robot SUBPAC::LANDRYTue Apr 29 1997 19:5719
    This note is a followup to the previous note concerning the robot
    cable (it's been ordered).
    
    We arrived today with the system down for a split cable (robot top
    cable) in the robots flex track.  It seems a couple of links from the
    tracks cover were missing, causing the cable to work it's way out and
    eventually pinch a gash in the sheathing.  Fortunetly, no wires within
    the sheathing were damaged.  After consulting with the powers that be,
    it was determined to seal up the gash, rebuild the flex track,
    re-install all removed hardware and test the system for proper
    operation.  Also, the cable end in the electronics cabinet was checked
    to assure that no moisture had permiated back there.  It was dry.
    
    The test runs worked fine.  The polish pad was changed and the tool
    was returned to manufacturing.
    
                 			A shift
    Jim Tooker has placed an order for new cabling.  Time will be arranged
    to replace the damaged one when the new one arrives.  
5.376Slurry supply change.FABSIX::L_SHKOLNIKOVThu May 01 1997 07:078
      
      We hooked up external slurry barrel of 18% solids, per EE
    request. Both slurry pumps were calibrated at 500ml. Pad,carriers
    and conditioner disk were replaced. The tool is currently running
    MQC's. Keep an eye on the level.


                                        D - shift.
5.37718% Slurry Tests for TonightSTRATA::POLLARDThu May 01 1997 18:0337
  SS25 18% Slurry Tests - Plans For Tonight:

    Today a lot of awesome tests were done with TERRIBLE results as we later
found the Cond'n arm was broken and not conditioning correctly.  So this 
means we're back at Square 1 for test.  So I'm going to need some PSS & Ops
support to get the following MQC testing done tonight.

     For wafers - I'm using the EXRSZ Dummies next to the Terminal By CMP.A1 &
the EMQC Monitors sitting next to the EXRSZ dummies. For the EMQC wafers slots
1-8 have been used & Slots 9-12 have been preread for Tx on Thickness.A1 
(under "Proc Eng").  Particles were Pre-read on Slot 9-12 & were < 20.

****IMPORTANT - For ALL MQC Runs *****
   -Please Save all MQC Monitor Wafers in original slots so I can study any 
    Particle Problems.
   -Please Save all Delta Tx Maps on UV1050
   -Record all particle (Large/Total/Scr Lngth) & Tx data in this Note.

----So here's the Plan:
1.  EE will finish up fixing Cond'n Arm and Install New Pad and do 10 Cond'n
    cycles.
2.  PSS/Ops - Run 10 Breakins as normal on "NewPad" recipe.
3.  Do Std MQC Run & Eval.  ("MQC" recipe).
4.  Do 6 Dummy Wafers on "18NewPad" Recipe.
5.  Do Std MQC Run & Eval.  If Uniformity is <10% then Go to Step 6.
    If Uniformity is > 10% then do 10 more Dummies on "18NewPad" recipe &
    repeat Std MQC then go to #6.
6.  Use recipe "18MQC1" recipe for MQC run with 2 monitors.
7.  Run 10 Dummy wafers on "18MQC1" Recipe.
8.  Use recipe "18MQC1" recipe for MQC run with 2 monitors.
9.  Use recipe "18MQC2" recipe for MQC run with 2 monitors.
10. Run 10 Dummy wafers on "18MQC2" Recipe.
11. Use recipe "18MQC2" recipe for MQC run with 2 monitors.

Thanks
Bryan

5.378Pad Conditioner roll pin Shot.FABSIX::J_SWENSONThu May 01 1997 20:0520
    Conditioning Arm
    ----------------
    
                  - During Eng experiment noticed conditioning arm jumping.
                  
                  - found roll pin for disk assy to be oxidized.
    
                  - Bought new roll pins
    
                  - installed new pin
              
                  - lubed assy
    
                  - conditioned pad for 10 cycles 
    
                  - returned system back to ops
    
    
    
                                                  C-Shift
5.379Results for 18% slurry testsFABSIX::R_THOMPSONFri May 02 1997 06:2344
1.  EE will finish up fixing Cond'n Arm and Install New Pad and do 10 Cond'n
    cycles.
2.  PSS/Ops - Run 10 Breakins as normal on "NewPad" recipe.
3.  Do Std MQC Run & Eval.  ("MQC" recipe).
4.  Do 6 Dummy Wafers on "18NewPad" Recipe.
5.  Do Std MQC Run & Eval.  If Uniformity is <10% then Go to Step 6.
    If Uniformity is > 10% then do 10 more Dummies on "18NewPad" recipe &
    repeat Std MQC then go to #6.
6.  Use recipe "18MQC1" recipe for MQC run with 2 monitors.
7.  Run 10 Dummy wafers on "18MQC1" Recipe.
8.  Use recipe "18MQC1" recipe for MQC run with 2 monitors.
9.  Use recipe "18MQC2" recipe for MQC run with 2 monitors.
10. Run 10 Dummy wafers on "18MQC2" Recipe.
11. Use recipe "18MQC2" recipe for MQC run with 2 monitors.
_______________________________________________________________________________
                         
**  processes completed through step 6, we ran out of the 18% slurry
    continue processing at step 7.

                    ** RESULTS**
           RATE DEV.            PARTICLES
                        STEP 3
        rem.rate    std.       lg.  tot.  sc.ct
left-    1594       20.1       602  1758  7.1
right-   1698       22.4       463  2629  48.9
                        STEP 5
        rem.rate    std.       lg.  tot.  sc.ct
left-    1579       21.0       35   378   1.0
right-   1471       18.1       20   170   16.1
                        STEP 5 (second try)
        rem.rate    std.       lg.  tot.  sc.ct
left-    1208       13.5       27   210   2.6
right-   1288       14.8       22   249   0
                        STEP 6
        rem.rate    std.       lg.  tot.  sc.ct
left-    1295       13.1       86   756   30.3
right-   1183       16.0       145  678   188

If you have any questions please contact me.
        Rick Thompson



5.380Scotty.. we need more slurry!!!SUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 02 1997 06:519
    
           We ran out of 18% slurry so the tests were
       halted.  I couldn't find any 18% slurry anywhere
       in HLO3.  Ashland will be here around 6:30am so
       I'll have Leon track them down to try and get
       us another barrel.  The tool is in WET MODE and
       UP/ENG.

                                     D shift   
5.381Back to business.FABSIX::L_SHKOLNIKOVFri May 02 1997 08:214
    There is a new barrel of 18% slurry connected to the tool.
    The slurry flow has been checked.
    
                                                Leon.
5.382addition to .379FABSIX::R_THOMPSONFri May 02 1997 08:407
This is an addition to note .379
We have recieved more of the 18% slurry and the tests have continued
I am currently running step 7. "Run 10 Dummy wafers on "18MQC1" Recipe."
All previous data was recorded on thickness.A4.(pre and post readings)
Pre readings for step 8&9 have been read on thickness.A5.
    Rick

5.383Scratches, bad carrierLUDWIG::STANDINGFri May 02 1997 20:127
    CMP.A1   Removed barrel of slurry and reconnected tool to bulk supply.
             Later tool went down for scratches on wafer, operator ran
    wafers through brushes and without brushes and determined that
    scratches were coming from brushes.  Brushes were replaced.
             Later the wafers were falling off right carrier, you can hear
    a clicking sound coming from the carrier.  Production was running lots
    and would not release the tool for a carrier change.
5.384Failed mqc'sFABSIX::R_POIRIERMon May 05 1997 07:037
    
    		Failed MQC's. Replaced Pad and conditioner.
    		Passed MQC's.
    
    
    				B Shift
    
5.385Failed MQCsFABSIX::R_POIRIERTue May 06 1997 07:3810

			Failed MQC's. Replaced pad, carriers and
			conditioning disk. Also had a problem
			with the conditioning disk dish. Would not
			drain.  Found the drain line plugged up.
			Un-plugged line.

					B Shift

5.386axis 5 motor control moduleSTRATA::FINNWed May 07 1997 21:095
    
    
Changed Axis 5 motor control module at EE request.

						A shift
5.387Pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 09 1997 06:015
    
    
       Pad replaced due to bad uniformity...
    
                              D shift
5.388pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 16 1997 05:585
    
        Replaced pad because it failed MQC's for bad unif. and
     had a bubble
    
                                            D shift
5.389Bad unif. - Robot issueSUBPAC::BJUBINVILLECMP Equipment EngineeringSat May 17 1997 08:2211
    
          Twice the robot crashed into the right and left send 
       elevators.  Found small tear in 1/8" vacuum line.  Cut back line
       and ran 50 wafer VERIFY.  All set....
    
    
           Tool failed 24 hr. MQC's for bad unif.  Replaced pad...
    
           Tool in VERIFY-U at this time and is ready for break-ins.
    
                                            BJ
5.390cmp.a1 crashFABSIX::J_SADINFreedom isn&#039;t free.Sun May 18 1997 19:5213
    
    
    	Replaced both carriers due to max wafer counts and then the right
    carrier began dropping wafers on the polish table (at lift off). It saw
    enough vacuum that it thought it still had the wafer and the
    conditioner crunched the wafer still on the polish table. Stripping the
    carrier revealed the possibility of a pinched hose, so the carrier was
    swapped with a new one. Cleaned up the tool and it's been running well
    since then. 
    
    
    a-team
    
5.391Pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 22 1997 06:056
    
         Replaced pad due to tool failing MQC's for high unif.
      Also, one incident with the right carrier sticking to pad.
      Swapped the right carrier with the left carrier to monitor
      carrier.
                                          D shift
5.392Wafers falling off right carrier.STRATA::STANDINGSat May 24 1997 20:006
    Tool went down for loosing wafers off the right carrier.  Did a spindle
    blow off and saw water coming out from the outer set of holes on the
    carrier.  Removed and disassembled carrier.  Replaced the silicone
    O-ring and the three O-ring screws.  Checked the check valves they were
    OK.  Reassembled carrier and checked for leaks and there were none. 
    Returned tool to production.
5.393A1 pad change status from C shiftFABSIX::B_JUBINVILLEWed May 28 1997 22:1018
From:	ASDG::POIRIER      "FAB6 CMP EQUIPMENT TECHNICIAN. PAGER # 254 FUNCTION 1. TEL:6610,5804,4774,6609." 28-MAY-1997 19:29:29.78
To:	FABSIX::B_JUBINVILLE,FABSIX::L_SHKOLNIKOV
CC:	POIRIER
Subj:	passdown



   Brian & Leon,

   Here is the latest situation:


   
   CMP.A1 - DOWN/REPAIR: Pad change in progress. We won't have time to finish.



                                                            C Shift. 
5.394Pad change - bad uniformitySUBPAC::BJUBINVILLECMP Equipment EngineeringThu May 29 1997 02:5510



        Took over from C shift.  Cleaned tool and replaced
     pad due to bad uniformity on the 24hr MQC's.  Tool 
     passed MQC's after pad change and tool was turned over
     to OPS.

                                       D shift
5.395Pad changeSUBPAC::BJUBINVILLECMP Equipment EngineeringFri May 30 1997 08:146
    
     
        Tool failed 24 hr. MQC's for uniformity.  Pad was replaced.
     The MQC's were run and we are waiting for the results.
    
                                          D shift
5.396Important work for CMP.A tools Sun Nit/Mon MornSTRATA::POLLARDFri May 30 1997 17:4929
**** IMPORTANT Info for Sunday night B shift on CMP.A tools! *******

Please DO the following if a normal PAD change is required on a CMP.A tool
on Sunday Night/Monday morning AFTER 2am:
  (this is in preparation to convert to the new 4" conditioning recipes which
   will be turned on by TDCO Monday during the day)

1. Install the 4" Conditioning Disk in place of 6"
2. Change the following variables:  Var#126- 95,000     Var#71- 0.59
3. Install New Pad as normal.
4. Use Recipe "CRFSTBKN" to do 10 Conditioning Sweeps
   (set # Sweeps back to 2 when finished).
5. Use Recipe "CRFSTBKN" to do 10 Break-ins.
6. On CMP.A1 use Recipe "MQCBP" to do MQC.
   On CMP.A2 use Recipe "MQCBP3" to do MQC.
7. If MQC's pass leave up for Bryan to put TDCO on-line at 8am.
   If fails go to #8.
8. If MQC's failed then evaluate whether the removal rate is Center Fast or
   Center Slow:  
   - If it is Center Fast Repeat #4 and #5 and #6  - if it passes
     system will wait for Bryan to put on TDCO at 8am.  If it fails start
     with another Pad.
   - If it is Center Slow do 10 conditioning cycles with Recipe "CRSLWBKN"
     and do 10 Breakin wafers with the same recipe then repeat #6.  If
     it passes system will wait for Bryan to put on TDCO at 8am.  If it fails
     start with another Pad.

Bryan Pollard

5.397TDCO on-line for CMP.A toolsFABSIX::B_POLLARDMon Jun 02 1997 16:59107
Please read the TDCO which went on-line today for New Conditioning cycles
and a New Pad Breakin Procedure.
Thanks,
Bryan

DISK$DCS2:[UDOC.QUAL]BK20-00000-0006.T   REV L    11-JUN-1996   PAGE  10 OF  13

| 6.1  Temporary D.C.O. Form

  Originator__Bryan Pollard____      TEMPORARY       TDCO # _________________
  Tel Ext__4036_Date__5/29/97__       DOCUMENT       Copy # _____   of ______
                                    CHANGE ORDER     Date Of Issue___________
_______________________________________________________________________________
 PROBLEM:                                             |1ST ISSUE APPROVAL STAMP
Conditioning part of all recipes have been updated for|EXPIRATION DATE
4" disk.  This change will extend the pad life from   |________
the typical 250-300 wafers up to 500 wafers.          |
Also changes have been made to break-in recipes.      |
______________________________________________________|_________________________
DESCRIPTION OF CHANGE:                                |2ND ISSUE APPROVAL STAMP
The conditioning sweep parameters (which are not      |EXPIRATION DATE 
typically seen by operators) have been updated for all|________
production recipes - MQC, PMD, ILD, STI.  The sweep   |
has also been changed from 1 to 2 & from Post         |
to Pre Condition.                                     |
                                                      |
The limit for Pad Life is extended to 500 wafers      |________________________
(with 24 hour MQC's in Spec - Section 4.1.11)         |3RD ISSUE APPROVAL STAMP
                                                      |EXPIRATION DATE 
Also changes have been to the Pad Breakin procedures -|________
these apply to section 3.1.1 and to various parts of  |
4.1 of the spec and are described on the second page. |
                                                      |
______________________________________________________|_______________________
DURATION OF TDCO OR NUMBER OF LOTS TO BE PROCESSED: 30 days
______________________________________________________________________________
DOCUMENT # / TITLE AFFECTED                  
HF6O-0069  /  STRAUSBAUGH 6DS-SP (CMP.A1, CMP.A2) Chemical Mechanical 
Planarization Operations Specification.
_______________________________________________________________________________
1ST ISSUE APPROVALS:| 2ND ISSUE APPROVAL | 3RD ISSUE APPROVAL | EARLY REMOVAL
                    |                    |                    | AUTHORIZATION
SIGNATURE/BADGE/DATE|SIGNATURE/BADGE/DATE|SIGNATURE/BADGE/DATE|SIGN./BADGE/DATE
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
                    |                    |                    |
                    |                    |                    |
____________________|____________________|____________________|_______________
IF LOT SPECIFIC: Upon completion of the lots, the originator must remove, sign, 
date this TDCO form and return it Document Control (HLO1-1/P5).

____________________              ____________              ___________________
    ORIGINATOR                        DATE                  DATE LOTS COMPLETED
                               ---------------
                               |d|i|g|i|t|a|l|
                               ---------------

DISK$DCS2:[UDOC.QUAL]BK20-00000-0006.T   REV L    11-JUN-1996   PAGE  10 OF  13

| 6.1  Temporary D.C.O. Form   
  
  Originator__Bryan Pollard____      TEMPORARY       TDCO # _________________
  Tel Ext__4036_Date__5/29/97__       DOCUMENT       Copy # _____   of ______
  Page 2                              CHANGE ORDER     Date Of Issue___________
_______________________________________________________________________________

The Break-In recipes and procedures for breaking in a New Pad have been altered-
the new procedure will be as follows (applies to section 4.1.6- 4.1.14):
  1. Run 10 Conditioning cycles with the recipe "CRFSTBKN" (stands for
     CenterFast Break-in - most new pads start with a Center Fast polish).
  2. Run 10 - 50K oxide wafers thru the recipe "CRFSTBKN".
  3. Do monitor run with recipe "MQC"
  4. If monitor fails uniformity then do the following:
     - If monitor is Center Fast - run 10 more conditioning cycles with 
       the recipe "CRFSTBKN" then do 10 more 50K oxide dummies with the same 
       recipe and repeat MQC.  If second MQC fails change the pad and start 
       over.  
     - If monitor is Center Slow - run 10 conditioning cycles with the recipe
       "CRSLWBKN" then do 10 more 50K oxide wafer with the same recipe and 
       repeat MQC. If second MQC fails change the pad and start over.
     - For either case, record in CMP Notesfile how many breakins are required 
       to get the uniformity in spec.

The Break-In recipes and procedures for breaking in a USED Pad have also been 
altered- the new procedure will be as follows (applies to section 4.1.6 - 
4.1.14):
   1. All breakins which previously were done with the "USEDPAD" recipe will
      be done with the "CRSLWBKN" recipe except if an MQC has failed
      due to being Center Fast then use the CRFSTBKN recipe.
   2. If a 24 hour MQC fails and is Center Slow then do 20 conditioning
      cycles with the "CRSLWBKN" recipe and repeat MQC - if it passes system
      is up if it fails then the system will require a pad change.

5.3984" conditioner installedFABSIX::J_SADINFreedom isn&#039;t free.Mon Jun 02 1997 20:029
    
    
    	Set up tool for 4" conditioner. Had to move outer bushing on pad
    conditioning arm assembly in order to make the correct outside
    position with the 4" conditioner. Variables 126 and 71 were changed to
    support the 4" conditioner.
    
    
    	a-shift
5.399Please Save the Pads!STRATA::POLLARDWed Jun 04 1997 10:426
NOTE - During the time of the 4" Disk TDCO - Please SAVE a full-length 
       cross-section of all pads on the A tools.  Label them with the tool,
       # of wafers, Date, and "Hold for Bryan Pollard".

Thanks
Bryan Pollard